### v1

# Machine Unlearning: A Comprehensive Survey

Abstract.
As the right to be forgotten has been legislated worldwide, many studies attempt to design unlearning mechanisms to protect users’ privacy when they want to leave machine learning service platforms. Specifically, machine unlearning is to make a trained model to remove the contribution of an erased subset of the training dataset. This survey aims to systematically classify a wide range of machine unlearning and discuss their differences, connections and open problems. We categorize current unlearning methods into four scenarios: centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in unlearning. Since centralized unlearning is the primary domain, we use two parts to introduce: firstly, we classify centralized unlearning into exact unlearning and approximate unlearning; secondly, we offer a detailed introduction to the techniques of these methods. Besides the centralized unlearning, we notice some studies about distributed and irregular data unlearning and introduce federated unlearning and graph unlearning as the two representative directions. After introducing unlearning methods, we review studies about unlearning verification. Moreover, we consider the privacy and security issues essential in machine unlearning and organize the latest related literature. Finally, we discuss the challenges of various unlearning scenarios and address the potential research directions.
machine unlearning
ccs: Computing methodolgies Machine learningccs: Computing methodolgies Aritificial intelligenceccs: Security and privacyccs: Theory of computation
1. Introduction
Over the past decade, enormously increased data and fast hardware improvement have driven machine learning developments quickly. Machine learning (ML) algorithms and artificial intelligence (AI) are embedded into day-to-day applications and wearable devices (Mahdavinejad et al., 2018). It further collects more and more users’ information, including users’ private data, such as driving trajectory, medical and online shopping records. On the one hand, such an enormous amount of data helps to improve ML and AI development further. However, on the other hand, it threatens users’ privacy and demands the enormous requirement of data management to ensure information security and privacy in machine learning.
Machine unlearning is recently proposed to protect users’ privacy when they want their data forgotten by a trained machine learning model. It draws increasing research attention because Right to be forgotten was legislated recently in many countries, such as the General Data Protection Regulation (GDPR) (Mantelero, 2013) in the European Union, PIPEDA privacy legislation in Canada (of the Privacy Commissioner of Canada, 2018), and the California Consumer Privacy Act (de la Torre, 2018) in the United States. The legislation requires that companies must take reasonable measures to achieve the erasure of personal data concerning. For a machine learning application, it means that individual users have the right to ask companies to delete their private data once collected from them. The deletion is not only erasing their data from a database. It needs to remove the contribution of the erased subset of whole training data from the already trained model, which process was first conceptualized as machine unlearning (Cao and Yang, 2015). Specifically, if a user (Alice) wants to exercise her right to be forgotten (Mantelero, 2013) (e.g., quitting a ML application), then the trained model of such application must ”unlearn” her data. Such a process includes two steps: first, a subset of the whole dataset previously used for training a ML model is requested to be erased; second, the ML model provider removes the contribution of the erased data from the trained model. A naive data-erasing method is retraining a new model from scratch (Bourtoule et al., 2021). However, the computation and storage costs of retraining are expensive, especially in complex learning tasks.
Many researchers tried to find efficient and effective methods to realize unlearning rather than naive retraining, and there are several common challenges, which are summarized as follows. Stochasticity of training: A huge amount of randomness exists in the training process in machine learning, especially in complicated models’ training periods such as CNNs (Gu et al., 2018) and DNNs (Sze et al., 2017). This randomness makes the training result non-deterministic (Bourtoule et al., 2021) and raises difficulties for machine unlearning to estimate the contribution of the specific erased data sample. Incrementality of training: The training process in machine learning is an incremental procedure. In other words, the model update on the former data point will influence the contribution of the data points fed into the model after this data. Determining a way to effectively remove the influence of the to-be-erased data samples from the trained model is challenging for machine unlearning (Koh and Liang, 2017). Catastrophe of unlearning: Nguyen et al. (Nguyen et al., 2020) indicated that an unlearned model usually performs worse than the model retrained from scratch. Especially when an unlearning method tries to remove a huge amount of data samples, the degradation could be exponential. They referred to such sharp degradation as catastrophic unlearning (Nguyen et al., 2020). Although several studies have explored ways to mitigate unlearning degradation by bounding the loss function or restricting unlearning threshold, how to eliminate catastrophic unlearning is still an open problem. Recently, many studies put efforts into solving these three main challenges and proposed many novel mechanisms which promote the progress of machine unlearning.
[…]
ARC
Machine UnlearningCentralized UnlearningUnlearning CategoriesExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2021; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021) Unlearning TechniquesSplit Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022)Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2021)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Distributed and Irregular Data UnlearningFederated Unlearning: (Wang et al., 2022; Liu et al., 2021, 2022a; Gupta et al., 2021; Wu et al., 2022b)Graph Unlearning: (Chien et al., 2022; Chen et al., 2021a; Cong and Mahdavi, 2022)Unlearning Verification: (Hu et al., 2022; Thudi et al., 2022; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a) Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021b; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021b; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
Figure 1. Our taxonomy for machine unlearning. The introduction order will also follow this figure. We classify current unlearning literature into four main scenarios: centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning.
The goal of this survey is to classify and systematize machine unlearning according to the reseaching problems and goals in unlearning processes, and to review their differences and connections, advantages and disadvantages. We first catalogue unlearning methods into four main parts, centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning, as shown in Fig. 1. Since there are many studies targeted at centralized unlearning, we use two sections to introduce related work in this scenario. We briefly classify the related unlearning methods into two main kinds, Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022; Nguyen et al., 2020), where exact unlearning extends the idea of naive unlearning to design efficient retraining scheme with less computation cost. Approximate unlearning aims to unlearn a posterior that is approximate to the model retrained on the remaining dataset, only using the original learned model and the erased data. We then detailedly introduce the popular implementing techniques of these two unlearning kinds. In addition to these centralized unlearning methods, we introduce the distributed and irregular data unlearning scenarios, including federated unlearning (Wang et al., 2022; Liu et al., 2021), and graph unlearning (Chien et al., 2022; Chen et al., 2021a). After implementing unlearning, verifying the unlearning effect is equally important and draws lots of research attention. Therefore, we review the studies about unlearning verification subsequently. Moreover, we consider the privacy and security issues essential in machine unlearning, and we organize and review the related unlearning literature about privacy threats, defences, and unlearning applications in the survey.
There is no comprehensive survey about machine unlearning because it is a pretty new research domain. Only several reviews on the specific unlearning scenarios were published on Arxiv. An introduction to machine unlearning, about the exact and approximate unlearning problems and how they are defined and solved by recently proposed methods, see (Mercuri et al., 2022). About certifiable machine unlearning for linear models, the algorithm introduction and experiments analysis can be found in (Mahadevan and Mathioudakis, 2021). Nguyen et al. (Nguyen et al., 2022b) summarized the general unlearning framework and added the unlearning verification part to it. They paid most of their attention to introducing problem formulations and technical definitions. None of them noticed and reviewed the privacy and security issues in machine unlearning. Moreover, they have not comprehensively discussed the differences and connections between different unlearning scenarios.
In our treatment, we review not only centralized machine unlearning but also other scenarios, including distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning. We discuss these unlearning scenarios in the literature and compare the difference and connections from each other. For all four unlearning scenarios, we list open questions that, if solved, would increase our understanding of machine unlearning or enhance our ability to design a suitable unlearning mechanism for different applications.
The main contributions of the paper are as follows.
  • We systematically catalogue machine unlearning studies according to their scenarios, which can be briefly classified into, centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning.
  • For the centralized unlearning scenario, which draws the most unlearning research attention, we further divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
  • We further explore the privacy threats that machine unlearning brings and discuss the applications of unlearning in defending against traditional security and privacy issues.
  • We discuss different challenges between various machine unlearning scenarios, from centralized unlearning to federated unlearning, graph unlearning, and privacy and security issues in unlearning. We list the related open questions of each scenario and present potential research directions to solve them.
The rest of the survey is organized as follows. The background of machine unlearning and the related technical tools are summarized in Section 2. After introducing the background, we will introduce the survey following the taxonomy order in Fig. 1. In Section 3, we present the primary categories of centralized unlearning: exact unlearning and approximate unlearning. In Section 4, techniques of centralized unlearning in different categories introduced in Section 3 are discussed in-depth. We collect and introduce the distributed and irregular data unlearning scenarios that we noticed, which are federated unlearning and graph unlearning, in Section 5. Section 6 introduces the unlearning evaluation and verification methods. The privacy and security issues in machine unlearning are divided into two sections. Section 7 introduces the privacy threats accompanied by machine unlearning. In Section 8, we discuss machine unlearning applications, which are mainly applied in dealing with security issues. In Section 9, we discuss the challenges of current unlearning methods in a general way. After that, we list the open gaps that new studies can target. At last, we summarise the survey in Section 10.
2. Background
2.1. Machine Unlearning
There are many urgent demands that motivate the hot machine unlearning research. The first is the privacy-preserving demand. Since the right to be forgotten has been legislated worldwide, machine unlearning is often formulated as a privacy-preserving problem where users can ask for their data erasure from the ML models. Besides privacy reasons, many other reasons promote the development of machine unlearning. For example, one common reason is usability. During the long time usage period of a recommendation application, users’ information will sometimes change, such as moving changes in their location. The application will produce unsuitable recommendations if it cannot completely delete the incorrect data from the model. Another reason is security. Deep learning models are vulnerable to adversarial attacks (Yuan et al., 2019) such as data poisoning (Tian et al., 2022). Hence, how to detect and remove adversarial poisoned data is essential to ensure models’ security. Once the attack type is detected, the model must erase these adversarial data’s influence using the unlearning mechanism.
Figure 2. Machine unlearning process.
To better understand how unlearning mechanisms work, we first introduce the unlearning problem and process following the machine unlearning framework demonstrated in Fig. 2. Let $\mathcal{Z}$ denote as a space of data items, then the particular (full) training dataset is $D\in\mathcal{Z}$. A learning process can be demonstrated as Step 0 in Fig. 2, i.e., training a model $M$ using an algorithm $\mathcal{A}$ on the training dataset $D$, denoted as $M=\mathcal{A}(D)$, where model $M$ is in a hypothesis space $\mathcal{H}$.
The unlearning process begins at an unlearning request when a user wants to remove his data $D_{e}$ from the trained model, Step 1 in Fig. 2. The requested unlearning data $D_{e}$ can be data samples, classes or graph nodes. Then, in Step 2, the server removes the contribution of $D_{e}$ using a machine unlearning algorithm $\mathcal{U}$. The unlearned model can be expressed as $M_{D\backslash D_{e}}=\mathcal{U}(D,D_{e},\mathcal{A}(D))$. In dealing with an unlearning request, we hope that the unlearned model $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ is the same as the retrained model $\mathcal{A}(D\backslash D_{e})$ (i.e., $\mathcal{U}(D,D_{e},\mathcal{A}(D))\simeq\mathcal{A}(D\backslash D_{e})$). Most unlearning studies ended at this step. However, we do need an effective evaluation metric to assess the unlearning effect as Step 3 in Fig.2. Therefore, we add the verification as the last step in the unlearning process, and we will introduce the related literature later.
2.2. Unlearning Evaluation Metrics
To compare the two models before and after unlearning, we need to define an evaluating metric $d(.)$ between $\mathcal{A}(D\backslash D_{e})$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ from the model parameter space $\mathcal{H}$. To this end, we briefly introduce several common evaluation distances metrics:
$L_{2}$-Norm. Wu et al. (Wu et al., 2020) proposed to utilise the Euclidean norm to evaluate the parameters of $\mathcal{A}(D\backslash D_{e})$ and that of unlearning $\mathcal{U}(D,D_{e},\mathcal{A}(D))$. Suppose $\theta$ are the model parameters learned by algorithm $\mathcal{A}(.)$. The $l_{2}$-norm or distance can be denoted as $||\theta_{\mathcal{A}(D\backslash D_{e})}-\theta_{\mathcal{U}(D,D_{e},\mathcal{A}(D))}||$, where $\theta_{\mathcal{A}(D\backslash D_{e})}$ is the model parameters of retraining from the scratch and $\theta_{\mathcal{U}(D,D_{e},\mathcal{A}(D))}$ is the modal parameters of unlearning algorithms $\mathcal{U}(.)$.
Kullback–Leibler divergence (KLD). In Bayes-based or Markov chain Monte Carlo-based unlearning methods, researchers used the Kullback–Leibler divergence or Jensen-Shannon divergence to optimize the approximate posterior since they can measure the distance of two probability distributions. Therefore, in recent unlearning studies, researchers also used the KLD between the retrained and unlearned models to estimate the unlearning effect (Nguyen et al., 2020). For discrete probability distritbutions of two model parameters $\mathcal{A}(D\backslash D_{e})$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ on the same probability space $\mathcal{H}$, the KLD from $\mathcal{A}(D\backslash D_{e})$ to $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ is denoted as $\text{KL}[\Pr(\mathcal{A}(D\backslash D_{e}))||\Pr(\mathcal{U}(D,D_{e},\mathcal{A}(D)))]$.
Parameter Leakage. Since membership inference attacks can infer if a sample is used in the training dataset of a model, recently, some works have leveraged the property to verify if the specific data is removed by unlearning mechanisms. Some studies (Hu et al., 2022) even proposed to backdoor the erased samples and then attack the unlearned model to check if the unlearned model can be attacked. If the unlearned model is backdoored, it proves that the unlearning algorithm cannot unlearn samples effectively. Conversely, if the backdoor trigger cannot attack the unlearned model, it proves that the unlearning algorithm is effective. Similar methods were also used in (Thudi et al., 2022; Ma et al., 2022) to evaluate the unlearning effect.
2.3. Majority of Tools Used in Unlearning
2.3.1. Differential Privacy (DP)
Differential privacy has become a popular standard for privacy protection in the statistic. In a DP model, a trusted analyzer collects users’ raw data and then executes a private mechanism to achieve differential privacy. The privacy requirement ensures the indistinguishability for any two outputs of neighbouring datasets, where neighbouring datasets mean the dataset only differs by replacing one user’s data, denoted as $X\backsimeq X^{\prime}$. A ($\epsilon,\delta$)-differential privacy algorithm $\mathcal{M}:\mathbb{X}^{n}\to\mathbb{Z}$ means that for every pair of neighboring datasets $X\backsimeq X^{\prime}\in\mathbb{X}^{n}$ and every subset $S\subset\mathbb{Z}$ has that $\mathcal{M}(X)\in S$ and $\mathcal{M}(X^{\prime})\in S$ are $\epsilon$-indistinguishable and $\delta$-approximate. The smaller the $\epsilon$ is, the more the privacy protection level increases. When the value of $\epsilon$ is 0, it implies that the outputted probability distribution of algorithm $\mathcal{M}$ cannot reflect any useful information about the datasets. A general DP mechanism based on adding Laplace noise was presented and theoretically analyzed in (Dwork et al., 2006).
2.3.2. Bayesian Variational Inference (VI)
In machine learning, the bayesian variational inference is used to approximate difficult-to-compute probability densities via optimization. In this part, we revisit the VI framework (Blei et al., 2017) that learns approximate posterior model parameters $\theta$ using Bayesian Theory (Bernardo and Smith, 2009). Given a prior belief $p(\theta)$ of an unknown model and a complete data trainset $D$, an approximate posterior belief $q(\theta|D)\sim p(\theta|D)$ can be optimized by minimizing the Kullback–Leibler divergence (KLD) (Joyce, 2011), $\text{KL}[q(\theta|D)||p(\theta|D)]$. The KLD is a type of statistical distance to measure how one probability distribution $q(\theta|D)$ differs from another probability distribution $p(\theta|D)$. However, it is intractable to compute the KLD exactly or minimize the KLD directly. Instead, the evidence lower bound (ELBO) (Blei et al., 2017) was proposed to be maximized, which is equivalent to minimize KLD between the two probability distributions. ELBO follows directly from $log(p(D))$ subtracting $\text{KL}[q(\theta|D)||p(\theta|D)]$, where the log marginal likelihood $log(p(D))$ is independent of $q(\theta|D)$. Since $\text{KL}[q(\theta|D)||p(D|\theta)]\geq 0$, the ELBO is a lower bound of $log(p(D))$.
In general training situations, ELBO is maximized using stochastic gradient ascent (SGA). The primary process is approximating the expectation $\mathbb{E}_{q(\theta|D)}[log(p(D|\theta))+log(p(\theta)/q(\theta|D))]$ with stochastic sampling in each iteration of SGA. We can use a simple distribution (e.g., the exponential family)to approximate computational ease posterior belief $q(\theta|D)$. When dealing with complex distribution whose density cannot be evaluated, we can maximize the ELBO with adversarial training by alternating between estimating the log-density ratio $log(p(\theta))/q(\theta|D)$ and maximizing the ELBO (Yu et al., 2019).
2.3.3. Privacy Leakage
Figure 3. Privacy leakage: a privacy reconstruction process
Privacy leakage occurs in both unlearning verification and privacy threats in unlearning two parts. In unlearning verification, researchers tried to use privacy leakage attacks to verify if the specific data is unlearned. In the privacy and security issues in unlearning, researchers tried to design effective inference attacks tailored to machine unlearning. The basic attack of privacy leakage in a machine learning setting is membership inference, which determines if a data sample was used in the ML model updating process or not. When a sample is fully known to the adversary, learning which model was trained based on it will leak information about the model. A generic membership inference process was introduced in (Shokri et al., 2017). Shokri et al. first trained the shadow models to simulate the target machine-learning model. Then, they observed and stored the different outputs of the shadow models based on different inputs, in or not, in the trainset. They used these stored outputs as samples to train the membership inference attack model.
Model inversion (Fredrikson et al., 2014), or privacy reconstruction (Salem et al., 2020) is another privacy threat in general machine learning. The goal of model inversion is to infer some missing attributes of input features based on the interaction with a trained machine learning model. Salem et al. (Salem et al., 2020) proposed a reconstruction attack target recovering specific data samples used in the model updating process by the different outputs of the model before updating and after updating. Latter, inferring the private information of updating data in conventional machine learning is transferred to infer the privacy of the erased samples in machine unlearning. In reconstruction attacks, the adversary first collects the different outputs using his probing data $D_{probe}$, including the original outputs $\hat{Y}_{\mathcal{M}}$ before unlearning, and the outputs $\hat{Y}_{\mathcal{M}_{-D_{e}}}$ after unlearning. Then he constructs the attack model based on the posterior difference $\delta=\hat{Y}_{\mathcal{M}_{-D_{e}}}-\hat{Y}_{\mathcal{M}}$. The attack model contains an encoder and decoder, which has a similar structure as VAEs (Kingma and Welling, 2014), and the main process is shown in Fig. 3.
3. Categories of Centralized Machine Unlearning
In this section, we classify existing centralized unlearning methods by their inherent mechanism and designed purposes. Detailed techniques will be introduced in the next section. From the former introduction, we know that naive retraining is the most effective manner to realize machine unlearning. However, it is inefficient because it needs to store the whole training dataset and retrain a new model from scratch. It consumes a huge amount of storage and computation costs, especially in deep learning scenarios. Therefore, researchers tried to design effective and efficient unlearning mechanisms. However, there were difficult challenges when we tried other unlearning methods. Three representative challenges are the stochasticity of training, the incrementality of training, and the catastrophe of unlearning. Exact unlearning was proposed to reduce the computation cost by splitting training sub-models based on pre-divided data sub-sets. It also divides the stochasticity and the incrementality into sub-models. It can unlearn an exact model by ensembling the consisting submodels, but it still needs to store all the split data subsets for fast retraining. Another approach is approximate unlearning. It can reduce both storage and computation costs because it unlearns only based on the erased data. However, it is difficult for approximate unlearning methods to control the accuracy degradation, due to the challenges in estimating stochasticity and incrementality during the training process. Most of them bounded their estimation to avoid dramatic accuracy reduction.
3.1. Exact Unlearning
Exact unlearning could also be called fast retraining, whose basic idea is derived from naive retraining from scratch. Following the background introduction of unlearning, we know the learning model $\mathcal{A}$ and unlearning model $\mathcal{U}$ based on the trainset $D$ and erased dataset $D_{e}\subseteq D$. If $\mathcal{U}(.)$ is implemented as naive retraining, the equality between $\mathcal{A}(D\backslash D_{e})\in\mathcal{H}$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))\in\mathcal{H}$ is absolutely guaranteed. However, naive retraining involves high computation and storage costs, especially for deep learning models and complex datasets (Thudi et al., 2022). Unlike naive retraining needs to retrain based on the whole remaining dataset, exact unlearning tries to retrain a sub-model based on the subset of remaining data to reduce computation cost. A general operation of exact unlearning is that they first divide the training dataset into several sub-sets. Then, they transform the learning process by ensembling the sub-models trained based on each sub-set as the final model. So that when unlearning request comes, they only need to retrain the trained sub-model based on the sub-set containing the erased data. They then ensemble the retrained sub-model and other sub-models as the unlearned model.
As we introduced above, we know that the primary goal of exact unlearning is to reduce the computation cost of retraining a new model by transforming the original learning algorithms into an ensembling form. It divides the stochasticity and incrementality into several sub-models to reduce their influence. However, to some extent, they sacrificed the storage cost because they needed to store the whole training dataset in a divided form. In (Cao and Yang, 2015), Cao and Yang transformed the traditional ML algorithms into a summation form. When an unlearning requirement comes, they only need to update several summations, ensuring the process runs faster than retraining from scratch. Bourtoule et al. proposed SISA (Bourtoule et al., 2021), which is short for Sharded, Isolated, Sliced, and Aggregated training. SISA divided the full data into shards and trained models separately in each shard. For unlearning, they simply need to retrain the shard that includes the erased data. Wu et al. (Wu et al., 2022a) proposed a framework that precisely models the influence of each training data point on the model concerning various performance criteria and removes the influence of the data that is required to delete. Golatkar et al. (Golatkar et al., 2021) proposed an unlearning method on deep networks, splitting the trained model into two parts. The core part based on the data will not delete, and the unlearning part on the erased data with parameters bound. These methods are efficient in computation, but they sacrifice the storage space to store the intermediate training parameters of different slices and the related training sub-sets.
Besides the huge storage cost, another big problem in exact unlearning is that it only suits the situation, in which the unlearning request only needs to remove a few samples with a low frequency. Suppose an unlearning request needs to remove many data samples (usually, they are not in the same previous divided sub-set). In that case, exact unlearning must retrains all these related sub-models or even all the sub-models in the worst situation. At the moment, exact unlearning is no longer computation efficient, and it still needs to store the whole training dataset or intermediate parameters.
3.2. Approximate Unlearning
Figure 4. Model changes when adding a new point or removing a point. (a) A normally trained classifying model classifies classes 1 and 2. (b) When a new point appears, the model is trained based on it, and the classifying line is pushed to classify it. (c) When we need to remove an influential point, we should recover the contribution of this point to the model. (d) When we remove a Non-influential point, the model may not need to change a lot.
Unlike exact unlearning could only reduce the retraining computation cost, approximate unlearning tries to directly unlearn based on the learned model and the erased data sample, which saves the computation and storage costs together. The goal of approximate unlearning is to unlearn a posterior that approximates the model trained on the remaining dataset, based on only the erased sub-dataset while not the remaining dataset (used by exact unlearning).
As we introduced before, the unlearning purpose is expecting that the unlearned model $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ to equal to the retrained model $\mathcal{A}(D\backslash D_{e})$. Since exact unlearning is implemented by retraining from the remaining dataset or sub-sets, they can almost guarantee equality before and after unlearning. However, since approximate unlearning tries to directly delete the contribution of the erased data from the trained model, how to exactly estimate the contribution of the erased data (the contribution contains the stochasticity and the incrementality ) and remove it becomes the core problem.
The text description of the changes between two different distribution spaces before and after removing the specific data is not intuitive. Fig. 4 shows illustrated changes when adding a new point or removing a point in a classifying model. When an influential point appears, it usually pushes the line to move forward than the original classifying line to identify it, as shown in Fig. 4 (b). When this influential point is requested to remove, the unlearning mechanism must recover the model to the original one that has not been trained by this specific point, as shown in Fig. 4 (c). However, when only unlearning a non-influential point, which may have almost non-influence on the model, the unlearned model may not change compared to the original trained model in this situation, as shown in Fig. 4 (d).
Many methods were proposed to implement approximate unlearning efficiently and effectively. The popular solutions are certified-removal and Bayes-based mechanisms, which will be introduced in technical detail in Section 4. Here, we briefly summarize current studies about approximate unlearning. A certified-removal mechanism (Guo et al., 2020) was an approximate unlearning method similar to differential privacy. Guo et al. (Guo et al., 2020) defined the $\epsilon-$approximate unlearning, which makes sure that the model after and before unlearning must be $\epsilon$-indistinguishable as the definition in DP. The difference between $\epsilon-$approximate unlearning and $\epsilon$-DP is that the mechanism $\mathcal{A}$ on differential privacy is needed never to memorize the data in the first place, which is impossible in machine unlearning. The machine learning model does not learn anything from the training dataset if $\mathcal{A}$ is differentially private (Bourtoule et al., 2021). Sekhari et al. (Sekhari et al., 2021) introduced a similar solution to (Guo et al., 2020), which removes the contribution of the erased data from the gradients during the unlearning process. To mitigate the unlearning accuracy reduction, they bounded the model parameters by their differential unlearning definition. Neel et al. (Neel et al., 2021) proposed unlearning methods allowing arbitrary sequences of updates, including addition and deletion, and defined the corresponding weak and strong unlearning algorithms.
Different from indistinguishable unlearning, other methods used the Bayes theorem to unlearn an approximate posterior directly. In (Nguyen et al., 2020), Nguyen et al. unlearned an approximate posterior based on the erased data via the Bayesian inference. Similarly, Fu et al. (Fu et al., 2022) proposed a machine unlearning algorithm for Markov chain Monte Carlo (MCMC) and further defined a knowledge removal estimator the evaluate the knowledge erasing. Nguyen et al. (Nguyen et al., 2022a) also explained the effectiveness of the approximate machine unlearning algorithm from the perspective of Markov Chain Monte Carlo. However, since those methods are approximately unlearning the contribution of all the erasing data, including the inputs and labels, it is easy to decrease the model accuracy to some extent after unlearning.
4. Techniques of Centralized Unlearning
Table 1. Summary of popular unlearning techniques
Unlearning Literature Taxonomy Requests Type Techniques Realization Method Year
SISA (Bourtoule et al., 2021) Exact unlearning Samples Split learning Data and Model Partition 2021
GraphEraser (Chen et al., 2021a) Exact unlearning Graph nodes Split learning Data Partition 2021
Amnesiac unl. (Graves et al., 2021) Exact unlearning Samples Split learning Partially retraining 2021
RecEraser (Chen et al., 2022) Exact unlearning Samples Split learning Balanced Data Partition 2022
ARCANE (Yan et al., 2022) Exact unlearning Samples Split learning Parition by Class 2022
HedgeCut (Schelter et al., 2021) Exact unlearning Samples Split learning Tree ensemble learning 2021
DeltaGrad (Wu et al., 2020) Exact unlearning Samples Split learning L-BFGS (Byrd et al., 1994) 2020
L-CODEC (Mehta et al., 2022) Approximate unlearning Samples Certified Data Removal Markov Blanket selection 2022
PUMA (Wu et al., 2022a) Approximate unlearning Samples Certified Data Removal SME 2022
Certified Removal (Guo et al., 2020) Approximate unlearning Samples Certified Data Removal LP 2019
$(\epsilon,\delta)$-unl. (Neel et al., 2021) Approximate unlearning Samples Certified Data Removal Perturbed gradient descent 2021
(Ginart et al., 2019) Approximate unlearning Samples Certified Data Removal Quantized and DC-k-means 2019
Graph unl. (Chien et al., 2022) Approximate unlearning Graph nodes Certified Data Removal Certified removal 2022
EUBO, rKL (Nguyen et al., 2020) Approximate unlearning Samples Bayesian-based Unlearning VBI 2020
MCU (Nguyen et al., 2022a) Approximate unlearning Samples Bayesian-based Unlearning Monte cario-based 2022
BIF (Fu et al., 2022) Approximate unlearning Samples Bayesian-based Unlearning MCMC 2022
  • Alogrihtm abbreviations
  • VBI: Variational Bayesian Inference, SME: Store medial estimation, MCU: Monte Cario-based machine unlearning, FIM: Fisher Information Matrix, MCMC: Markov chain Monte Carlo, BIF: Bayesian inference forgetting, LP: Loss perturbation (Chaudhuri et al., 2011), TF-IDF: Term Frequency Inverse Document Frequency, EUBO: Evidence upper bound, rKL: reverse Kullback–Leibler
The naive retraining method is the most effective way to realize machine unlearning. However, it is expensive in both computation and storage. Exact unlearning extends the idea of retraining and tries to reduce the computation cost of unlearning. Approximate unlearning was proposed to find a way to reduce computation and storage consumption together. It is usually realized by unlearning based on the erased data only. In this section, we will introduce the popular techniques used in existing unlearning methods, in both exact unlearning and approximate unlearning. The dominant studies are summarized in Table 1, where the primary technique used in exact unlearning is split learning, and two primary techniques used in approximate unlearning are certified data removal and bayesian-based unlearning, respectively. The introduction order of these techniques will mainly follow the order of the unlearning techniques in Fig. 1.
4.1. Split Unlearning
Since most of the exact unlearning methods tried to split the full dataset into multi sub-sets and divide the model training process, we call this kind of unlearning technique split unlearning. The main process of split unlearning is illustrated in Fig. 5 (b). By contrast, we show the process of the naive retraining method in Fig. 5 (a). From Fig. 5 (a), we can see that when we unlearn the erased data by naive unlearning, we need two steps to realize it: first, remove the data from the original training dataset; second, retrain a new model based on the remaining dataset. Since it needs to store the whole training dataset and retrain from the remaining dataset, it often entails significant computational and stored overhead. Researchers saw the huge computation cost of naive retraining. They proposed many exact unlearning methods, and we summary as split unlearning because they were all based on data or model split techniques. As shown in Fig. 5 (b), the split unlearning technique can be summarized into four steps. Different from naive unlearning trained based on the whole remaining trainset, the first phase of split unlearning is dividing the original full trainset into multiple disjoint shards. All the constituent models are trained based on each split data slice. Then, in the second phase, when the unlearning request comes, they only need to remove the requested data from the split slice and retrain this slice’s constituent model in the third phase. In the last phase, the split unlearning aggregates the retrained and other constituent models together as a new unlearned model.
The first split unlearning is proposed by Cao and Yang (Cao and Yang, 2015). They split the original learning algorithms into a summation form. In a regular machine learning form, the model directly learns from the training dataset. However, in the summation form, they first train a small number of constituent models, which learn from several parts of the full trainset and then aggregate these intermediate models as the final learning model. So that when unlearning, they only need to retrain the constituent model that contains the information of erased data. It can efficiently speed up retraining time and reduce computation costs. In (Cao and Yang, 2015), the authors indicated that support vector machines, naive Bayes classifiers, k-means clustering, and many machine learning algorithms could be implemented in a summation form to reduce the retraining cost. The statistical query (SQ) learning (Kearns, 1998) guarantees the summation form. Although algorithms in Probably Approximately Correct (PAC) setting can be converted to the SQ learning setting, a huge amount of complex models, such as DNNs, cannot be converted to SQ learning efficiently.
Figure 5. (a) Naive unlearning. Only two steps: remove the erased data from the whole dataset and retrain the model based on the remaining dataset. (b) Split unlearning. It contains four steps: 1. split the original dataset into $n$ shards, 2. remove the erased data from the corresponding shard, 3. retrain the sub-model of this shard, 4. ensemble all sub-models as the final model.
Then, Bourtoule et al. (Bourtoule et al., 2021) and Yan et al. (Yan et al., 2022) proposed advantaged methods unlearn samples suitable on deep neural networks. The primary idea of (Bourtoule et al., 2021; Yan et al., 2022) is also similar to the process shown in Fig. 5 (b). In (Bourtoule et al., 2021), Bourtoule et al. named their unlearning method the SISA training approach. SISA was implemented on deep neural networks, including two Conv. layers followed by two FC layers, Wide ResNet-1-1 and ResNet-50 models on multiple datasets, including MNIST, Purchase, SVHN, CIFAR100, Imagenet, and Mini-Imagenet. Unlike the original split unlearning dividing the dataset and transforming learning algorithms to summation form, Yan et al. proposed ARCANE (Yan et al., 2022), which transforms conventional ML into ensembling multiple one-class classification tasks. When many unlearning requests come, it can reduce retraining costs, which is not considered by previous work.
Chen et al. extended exact unlearning methods to recommendation tasks and proposed RecEraser (Chen et al., 2022), which has similar architecture as split unlearning in Fig. 5 (b). RecEraser is tailored to recommendation systems, which can efficiently implement unlearning. Specifically, they designed three data partition schemes to split recommendation data into balanced pieces and propose an adaptive aggregation method based on an attention mechanism. They conducted the experimental evaluation on three real-world datasets, YELP2018, Movielens-1m, and Movielens-10m. These three datasets are widely used to assess the effectiveness and efficiency of recommendation models.
Schelter et al. proposed HedgeCut (Schelter et al., 2021), which implemented machine unlearning on tree-based ML models in a split unlearning similar form. Tree-based learning algorithms are trained by recursively partitioning the training data, locally optimizing a criterion such as Gini gain (Wehenkel and Pavella, 1991). HedgeCut focuses on implementing fast retraining for these methods. Furthermore, they evaluated their method on five publicly available datasets on both accuracy and running time.
Another method that is similar to split unlearning is Amnesiac Unlearning (Graves et al., 2021). The intuitive idea of Amnesiac Unlearning is to store the parameters of training batches and then subtract them when unlearning requests appear. In particular, it first trains the learning model by adding the total gradients $\sum_{e=1}^{E}\sum_{b=1}^{B}\nabla_{\theta_{e,b}}$ to the initial model parameters $\theta_{initial}$, where $E$ is the training epochs, and $B$ is the data batches. During the training process, they kept a list of $SB$, which records the batches containing the private data. The list could be in the form of an index of batches for each training example, an index of batches for each training category or any other information expected. When the unlearning request comes, a model using Amnesiac unlearning needs only to remove the updates from each batch $sb\in SB$ from the learned model $\theta_{M}$. As Graves et al. mentioned, using Amnesiac unlearning effectively and efficiently removes traces of the erased data that could be inferred via state-of-the-art privacy inference attacks and does not degrade the accuracy of the model in any other way.
4.2. Certified Data Removal
Certified data removal unlearning methods usually define their unlearning algorithms as $\epsilon$-indistinguishable unlearning, which is similar to the differential privacy definition. Most of them use the Hessian matrix (Bishop, 1992) to evaluate the contribution of erased data samples for unlearning subtraction. After estimating the contribution of the erased data samples, they unlearn by subtracting these contributions with an updating bound from the unlearning model.
In (Guo et al., 2020), Guo et al. proposed a certified data removal mechanism, which assumes removing the last training sample, $(x_{n},y_{n})$. Specifically, they defined a removal mechanism that approximately minimizes $\mathcal{L}(\theta;D^{\prime})$ with $D^{\prime}=D\backslash(x_{n},y_{n})$. The loss gradient at sample $(x_{n},y_{n})$ can be denoted as $\Delta=\lambda\theta^{*}+\nabla l((\theta^{*})^{T}x_{n},y_{n})$ and the Hessian of $L(\cdot;D^{\prime})$ at $\theta^{*}$ by $H_{\theta^{*}}=\nabla^{2}L(\theta^{*};D^{\prime})$. Then they applied a one-step Newton update to the gradient impact of the removed point $(x_{n},y_{n})$ on the model $\theta$. Under their observation, they found that directly removing the Hessian contribution from the direction of the gradient will leak the privacy of the training data that was expected to be removed. They used the loss perturbation technique (Chaudhuri et al., 2011) to hide this information. It used a random linear term to perturb the empirical risk and ensures that the outputs of their method $\mathcal{U}(D,D_{e},\mathcal{A})$ is $\epsilon$-indistinguishable between the retrained model $\mathcal{A}(D\backslash D_{e})$. In(Mehta et al., 2022) and (Thudi et al., 2022), they designed unlearning algorithms following the certified data removal definition in (Guo et al., 2020). Mehta et al. unlearned via their proposed efficient Hessians, L-FOCI (Mehta et al., 2022). Thudi et al. (Thudi et al., 2022) used membership inference as a verification error to adjust the unlearning process on stochastic gradient descent (SGD) optimization.
Another similar unlearning method is PUMA. In (Wu et al., 2022a), Wu et al. proposed a new data removal method through gradient re-weighting called PUMA, which also used the Hessian Vector Product (HVP) term. They first estimated and recorded individual contributions of $(x_{i},y_{i})$, where the estimation is limited to less than a dot product between the pre-cached HVP term and individual gradient. When the unlearning request comes, they subtract the estimation of the erased samples to update the model parameters.
Ginart et al. (Ginart et al., 2019) extended certified data removal to k-means clustering algorithms. They formulated the unlearning problem of efficiently removing personal data information from trained clustering models and proposed two different efficient deletions for k-means clustering, quantized k-mean and divide-and-conquer k-means. In their work, both algorithms have theoretical guarantees and strong empirical results.
DeltaGrad (Wu et al., 2020), an algorithm for fast retraining of SGD-based models, was proposed by Wu et al. to small changes of data, based on the idea of ”differentiating the optimization path” concerning the data and Quasi-Newton methods. They theoretically proved that their algorithm could approximate the right optimization path rapidly for the strongly convex objective. DeltaGrad starts with a ”burn-in” period of first iterations, where it computes the full gradients exactly. After that, it only calculates the complete gradients for every first iteration. For other iterating rounds, it operates the L-BGFS algorithm (Byrd et al., 1994) to compute Quasi-Hessians approximating the true Hessians, keeping a set of updates at some prior iterations.
Sekhari et al. (Sekhari et al., 2021) further given a strict separation between $\epsilon$-indistinguishable unlearning and differential privacy. Different from (Guo et al., 2020), in order to utilise tools of differential privacy (DP) for ML, the most straightforward manner is to ignore the special set of erasure demands $D_{e}$ and create an unlearning mechanism $\mathcal{U}$ that only relies on the learned algorithm $\mathcal{A}(D)$. In particular, the unlearning algorithm is of the form $\mathcal{U}(D_{e},\mathcal{A}(D))=\mathcal{U}(\mathcal{A}(D))$ and makes sure the true unlearned model $\mathcal{U}(\mathcal{A}(D))$ is $\epsilon$-indistinguishable to $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$. Notice the difference between (Guo et al., 2020) and (Sekhari et al., 2021). In the definition of (Guo et al., 2020), their $\epsilon$-indistinguishable unlearning is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{A}(D\backslash D_{e})$, but here is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$. Such a pair of algorithms in (Sekhari et al., 2021) would be differential private for $D$, where the neighbouring datasets mean that for two datasets with an edit distance of $m$ samples. The guarantee of DP unlearning is stronger than the model distribution undistinguishable unlearning guarantee in (Guo et al., 2020), and thus it suffices to satisfy it. Based on the definition of (Sekhari et al., 2021), they pointed out that any DP algorithm automatically unlearns any $m$ data samples if it is private for datasets with edit distance $m$. Therefore, they derivate the bound on deletion capacity from the standard performance guarantees for DP learning. Furthermore, they determine that the existing unlearning algorithms can delete up to $\frac{n}{d^{1/4}}$ samples meanwhile still maintaining the performance guarantee w.r.t. the test loss, where $n$ is the size of the original trainset and $d$ is the dimension of trainset inputs.
The above methods are all answering the problem of unlearning. Neel et al. (Neel et al., 2021) extended the unlearning definition to the update, which includes both ’add’, $D\cup\{z\}$, and ’delete’ as $D\backslash\{z\}$, where $z\in\mathcal{Z}$ is a data point. They follow the definition in (Guo et al., 2020) and define the similar $(\epsilon,\delta)$-indistinguishability. Moreover, they extend $(\epsilon,\delta)$-indistinguishability for both ’add’ and ’delete’ updates, which can also be denoted as $(\epsilon,\delta)$-publishing.
4.3. Bayesian-based Unlearning
Different from certified data removal that unlearns samples by subtracting corresponding Hessian matrix estimation from trained models, Bayesian-based unlearning tries to optimize an approximate posterior of the remaining dataset only using the trained model (the posterior of the original full dataset) and the erased data. The exact Bayesian unlearning posterior can be derivated from the Bayesian rule as $p(\theta|D_{r})=p(\theta|D)\ p(D_{e}|D_{r})/p(D_{e}|\theta)$, where $\theta$ is the posterior (i.e., model parameters). $D$ is the original full trainset, and $D_{r}$ is the remaining dataset which removes the erased data $D_{e}$, denoted as $D=D_{e}\cup D_{r}$ and $D_{e}\cap D_{r}=\emptyset$. If the model parameters $\theta$ are discrete-valued, $p(\theta|D_{r})$ can be directly obtained from the Bayesian rule. The use of a conjugate prior furthermore makes unlearning relatively simple.
However, in practice, it is not easy to achieve the exact posterior, not to mention the unlearning posterior. In (Nguyen et al., 2020), Nguyen et al. tried doing likewise at the beginning. They defined the loss function using the KLD between the approximate predictive distribution $q_{u}(y|D_{r})$ and the exact predictive distribution $p(y|D_{r})$ where the observation $y$ (i.e., drawn from a model with parameters $\theta$) is conditionally independent of $D_{r}$ given $\theta$. They bounded this loss function by the KLD between posterior beliefs $q_{u}(\theta|D_{r})$ and $p(\theta|D_{r})$ and further proposed evidence upper bound (EUBO) as the loss function to unlearn the approximate unlearning posterior. To avoid the overestimation of using KL divergence to optimize the posterior, they introduced an adjusted likelihood to control the unlearning extent.
In (Nguyen et al., 2022a), the authors also studied the problem of ” unlearning” particular erased subset samples from a trained model with better efficiency than retraining a new model from scratch. Toward this purpose, Nguyen et al. proposed a Markov chain Monte Cario-based machine unlearning (MCU) method deriving from the Bayesian rule. They experimentally proved that MCU could unlearn the erased subsets of the whole training dataset from a prepared model effectively and efficiently.
Fu et al. (Fu et al., 2022) converted the Markov chain Monte Carlo unlearning problem into an explicit optimization problem. Then they proposed $\epsilon-$Knowledge removal, which was a little similar to certified removal methods, but they defined that the KLD between the unlearned and the retrained model must be less than $\epsilon$. To quantify the explicit $\epsilon-$knowledge removal, they proposed a knowledge removal estimator to evaluate the difference between the original and unlearned distributions. As they indicated, though their algorithm cannot wholly remove the learned knowledge from the already trained distribution, their method can still help the unlearned model approach its local minimum.
5. Distributed and Irregular Data Unlearning
We have introduced popular centralized machine unlearning in the former two sections. Most of them are requested to delete the regularly structured data. Their main purpose is to reduce the computation and storage costs and the accuracy degradation caused by unlearning. Recently, we noticed that some researchers paid attention to realizing unlearning in other scenarios, such as in distributed and irregular data learning scenarios. In this section, we will introduce the literature about federated unlearning and graph unlearning as two representative directions.
5.1. Federated Unlearning
Federated learning (McMahan et al., 2016, 2017) is first proposed to protect the participating users’ privacy during a machine learning training process in distributed situations. All participants will only upload their locally trained model parameters instead of their sensitive local data to the server for the model training. Therefore, in a federated learning scenario, the limited access to the dataset will become a unique challenge to implement unlearning.
Since the local data cannot be uploaded to the FL server side, most federated unlearning methods tried to unlearn a certain client’s influence from the trained model by storing and estimating the contribution of uploaded parameters. In this situation, they can implement federated unlearning without interacting with the client. The two representative methods are (Liu et al., 2021; Wu et al., 2022b). Liu et al. (Liu et al., 2021) proposed FedEraser to sanitize the contribution of a federated client on the global federated learning model. In particular, during the FL global training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates. Based on these operations, they reconstructed the unlearned global model instead of retraining a new model from scratch. However, FedEraser can only unlearn one client’s data, which means it must unlearn all the contributions of this specific client’s data. It is unsuitable for a client who wants to unlearn a small piece of his data. Wu et al. (Wu et al., 2022b) tried to unlearn a client’s influence from the federated global model by removing the historical updates from the global model. They implemented federated unlearning by using knowledge distillation to restore the contribution of clients’ models, which does not need to rely on clients’ participation and any data restriction.
Wang et al. (Wang et al., 2022) explored the problem of how to selectively unlearn a category from a trained CNN classification model in FL. The biggest challenge is that the local data used for FL global model’s training cannot be accessed globally. Therefore, they explored the inner influence of each channel in deep and observed that various channels have distinct impacts on different categories. They proposed a method that does not require accessing the data used for training and retraining from scratch to cleanly scrub the information of particular categories from the global FL model. A method called Term Frequency Inverse Document Frequency (TF-IDF) was introduced to quantize the class discrimination of channels. Then, they pruned the channels with high TF-IDF scores on the erased classifications to unlearn them. Evaluations of CIFAR10 and CIFAR100 demonstrate the unlearning effect of their method.
Different from unlearning a whole client’s influence and unlearning a class, Liu et al. (Liu et al., 2022a) investigated how to unlearn data samples in FL. They first defined a federated unlearning problem and proposed a fast retraining method to remove data points from the federated learning global model. Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods (Xu et al., 2020) and the first-order Taylor approximate technique. They used the practical Fisher Information Matrix (FIM) (Ly et al., 2017) to model the Hessian matrix at a low cost. When unlearning, they subtract the estimation matrix from gradients with a bound.
5.2. Graph Unlearning
We introduce graph unlearning as a representative kind of irregular data unlearning. In (Chien et al., 2022; Chen et al., 2021a; Cong and Mahdavi, 2022), researchers extend regular data machine unlearning to a graph data scenario. Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig. 6. Therefore, Chien et al. (Chien et al., 2022) proposed node feature unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC). Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees. They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
Figure 6. (a) Regular data unlearning, which only unlearns the data sample. (b) Graph unlearning. Since the graph data contains both node features and connective edge information, we may need to unlearn only edges information or both edges and nodes in graph unlearning.
Chen et al. (Chen et al., 2021a) found that applying SISA (Bourtoule et al., 2021) unlearning methods to graph data learning will severely harm the graph-structured information, resulting in model utility degradation. Therefore, they proposed GraphEraser to implement unlearning tailored to graph data. Similar to SISA, they first cut off some connecting edges to split the total graph into some sub-graphs. Then they trained the constituent graph models on these sub-graphs and ensemble them for the final prediction task. To realize graph unlearning efficiently, they proposed two graph partition algorithms and corresponding aggregation methods based on them.
In (Cong and Mahdavi, 2022), Cong and Mahdavi filled in the gap between regularly structured data unlearning and graph data unlearning by studying the unlearning problem on the linear-GNN. To unlearn the knowledge of a specific node, they propose a projection-based unlearning approach PROJECTOR that projects the weight parameters of the pre-trained model onto a subspace that is irrelevant to the deleted node features. PROJECTOR could overcome the challenges caused by node dependency and is guaranteed to unlearn the deleted node features from the pre-trained model.
6. Unlearning Evaluation and Verification
Table 2. Evaluation and Verification Metrics, Datasets
Evaluation Metrics Formula/Description Usage Literature Popular Datasets
Accuracy Accuracy of the unlearned model on erased datasets and remaining datasets To evaluate the predictive accuracy of the unlearned model (Bourtoule et al., 2021; Guo et al., 2020), … MNIST (LeCun et al., 1998) CIFAR SVHN LSUN Photo (Shchur et al., 2018) ImageNet (Deng et al., 2009) Adult Credit info PID (Asuncion and Newman, 2007) PWD (Liu et al., 2010) AG News (Zhang et al., 2015) GTSRB (Stallkamp et al., 2011) Covtype (Blackard and Dean, 1999) HIGGS (Baldi et al., 2014) RCV1 (Glewis et al., 2004) Market-1501 (Zheng et al., 2015) YELP2018 Movielens-1m Movielens-10m Cora (Sen et al., 2008) Citseer (Yang et al., 2016) Pubmed (Fey and Lenssen, 2019) ogbn-arxiv (Hu et al., 2020) Computers (McAuley et al., 2015) CS (Shchur et al., 2018) Physics (Shchur et al., 2018)
Running time The training time of unlearning process To evaluate the unlearning efficiency (Cao and Yang, 2015; Guo et al., 2020), …
L2-norm The parameters differences between the retrained and unlearned models $||\theta_{1}-\theta_{2}||$ To evaluate the indistinguishability between two models (Wu et al., 2020)
KL-Divergence The KLD between the distribution of the unlearned and retrained model: $\text{KL}(A(D_{r})||U(D_{e},A(D)))$ To evaluate the indistinguishability between model parameters (Nguyen et al., 2020)
JS-Divergence The JS divergence between the predictions of retrained and unlearned model: $JS(A,U)=0.5\cdot KL(A||Q)+0.5\cdot KL(U||Q)$ To evaluate the indistinguishability between model outputs (Chundawat et al., 2022)
Membership inference Recall (##\##detected items /#absent#/\#/ #erased items) To verify if the erased data is unlearned by the model (Graves et al., 2021)
Forgetting rate (FR) $FR=\frac{AF-BF}{BT}$, where AF, BF and BT are defined below To measure the rate of samples that are changed from member to non -member after unlearning (Ma et al., 2022)
[…]
EMA Ensembled Membership Auditing: Ensemble multiple membership metrics and utilizes Kolmogorov-Smirno (KS) statistical tools to obtain a final auditing score To verify if a unlearned model mem -orizes a query dataset (Liu and Tsaftaris, 2020)
MIB Membership Inference via Backdooring: Achieve membership inference for the backdoored data by querying a certain number of black-box queries to the model To verify if the model unlearns the backdoored data (Hu et al., 2022)
Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effect is also an important problem in unlearning. In Section 2, we have introduced the L2-norm (Wu et al., 2020), KLD (Nguyen et al., 2020), and privacy leakage as the unlearning verification metrics. The common metrics also include the accuracy to assess the predictive accuracy of the unlearned model and running time to evaluate the unlearning efficiency. We have listed recent evaluation and verification metrics in Table 6, where the popular experimental datasets are furthermore presented. Since accuracy, running time, L2-norm, and KLD are fixed evaluation metrics, we do not discuss them in detail here. In this section, we pay attention to the new evaluation metrics that recent studies tailored to unlearning. Most of them are based on privacy leakage.
In (Liu and Tsaftaris, 2020), Liu and Tsaftaris indicated that the data auditing problem in machine unlearning is becoming popular. They used a calibration dataset and the Kolmogorov-Smirno (KS) (Demidenko, 2004) distance to confirm if an erased dataset is removed. Their challenge was that the method would fail when the calibration dataset was out of high quality or when the query samples were similar to the training dataset. An Ensembled Membership Auditing (EMA) (Huang et al., 2021a) method was proposed to overcome these limitations when auditing data erasure from a trained model. EMA is a 2-step data removal auditing method that ensembles several metrics and statistical tools. In the first step of EMA, the auditor uses various metrics to infer if the model remembers the samples of the query dataset. In the second step, EMA aggregates a final auditing score by ensembling many membership metrics and statistical tools. They verified whether a trained model forgets query data by the final auditing score. They conducted the experiments using standard datasets, including MNIST, SVHN, and the Chest X-ray dataset.
Inspired by backdoor attacks in ML, Hu et al. (Hu et al., 2022) proposed Membership Inference via Backdooring (MIB). MIB leverage the property of backdoor attacks that backdoor triggers will misadvise the trained model to predict the backdoored sample to other wrong class. The main idea of MIB is that the user proactively adds the trigger to her data when publishing them online so that she can implement the backdoor attacks to determine if a model has been trained based on her data. MIB evaluate the membership inference for the triggered data by calculating the results of a certain number of black-box queries to the targetted model. Then, they observed the results between the targeted model and clean models to infer if the model was trained using the backdoored data.
A similar membership-inference-based method was proposed in (Ma et al., 2022). Ma et al. verified the effectiveness of an unlearning method by their proposed uniform metric called forgetting rate (FR). Their methods were inspired by membership inference attacks and defined the evaluation metrics based on the observation of membership inference. Suppose the dataset $D_{e}$ is the erased dataset; the FR of an unlearning method is denoted as $FR=\frac{AF-BF}{BT}$. In their FR definition, $BF$ and $AF$ mean the number of samples in $D_{e}$ that are predicted to be false by a membership inference attack before and after unlearning. $BT$ is the number of samples in $D_{e}$ predicted to be true by a membership inference attack before machine unlearning. According to this definition, $FR$ presents an instinctive evaluation of how many data points are changed from member to non-member after unlearning. If an unlearning method achieves that $AF>BF$ on the condition that $BT>0$, it means this method is effective. By contrast, the unlearning will be meaningless. They also proposed an unlearning method called Forsaken and evaluated Forsaken’s performance on eight standard benchmark datasets. The main idea of Forsaken is to continuously generate masked gradients, apply them to the model’s neurons, and model them to forget the contribution of the specific samples. To implement this process, they introduced a mask gradient generator. The proposed mask gradient generator could adjust the volume of the mask gradients to avoid unlearning catastrophes based on the observation of the updated model.
Sommer et al. (Sommer et al., 2022) introduced Athena, which leverages the property of backdoors, to verify the effectiveness of unlearning. Athena effectively certifies whether the data is deleted from an unlearning method with high confidence. Thus, it provides a basis for quantitatively inferring unlearning. In their backdoor-based verification scheme, they first backdoor users’ data and then test the backdoor success probability to infer if the data is deleted.
7. Privacy Threats on Machine Unlearning
Although machine unlearning was first proposed to protect users’ privacy, many researchers have noticed that unlearning brings new privacy threats simultaneously. Similar to machine learning scenarios, there are two main kinds of privacy threats in unlearning. They are membership inference attacks and private information reconstruction attacks. In this section, we introduce the two prevalent unlearning privacy threats detailedly.
7.1. Membership Inference Attacks in Unlearning
Chen et al. (Chen et al., 2021b) first pointed out that when a model is unlearned, the difference in the outputs from the model before and after unlearning leaks the privacy of the erased data. Then they proposed the corresponding membership inference attack pipeline in unlearning, which includes three phases: posteriors generation, feature construction, and membership inference.
  1. (1)
    Posteriors Generation. Suppose that the attacker has the ability to access two versions of the ML model, the model $\theta_{\mathcal{A}}$ before unlearning and the unlearned model $\theta_{\mathcal{U}}$. Assume a target sample $e$, the attacker queries $\theta_{\mathcal{A}}$ and $\theta_{\mathcal{U}}$, and obtains the corresponding posteriors, $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, which also called as confidence values in (Shokri et al., 2017).
  2. (2)
    Feature Construction. After achieving the two posteriors $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, the attacker sums them to make the inference feature vector $F$. Common methods exist to construct the feature vector shown in (Chen et al., 2021b).
  3. (3)
    Inference. After the attacker finishes training the attack model based on the created features $F$, he inputs the collected feature to the inference model to predict if the specific sample $e$ is in the erased dataset of unlearning models.
In (Chen et al., 2021b), they assumed that the attacker has admission to two versions of ML models before and after unlearning, but it is sometimes impractical, especially in black-box learning scenarios. Lu et al. (Lu et al., 2022) further proposed a label-only membership inference attack method to infer if a data point is unlearned by the model, eliminating the dependence on accessing posteriors. Their basic idea is that the same noise injection on candidate data points will show different results for the sample in or not in the training dataset. Thus, they made the adversary continuously query the original and unlearned models and add noise to change their outputing labels. By observing the disturbance amplitude, they can determine whether an item is deleted.
Golatkar et al. (Golatkar et al., 2020b) derived an upper bound to confirm the maximizing knowledge that can be extracted from a black-box model. They queried the model with an image and obtained the related output. They then used the entropy of the result probabilities to construct an effective black-box membership inference (Shokri and Shmatikov, 2015) attack in machine unlearning.
7.2. Privacy Reconstruction Attacks in Unlearning
Privacy reconstruction is another popular attacking kind in machine unlearning. In an unlearning scenario, Gao et al. (Gao et al., 2022) proposed the deleted reconstruction attacks to recover the removed data from the outputs of the original and unlearning models. In their work, they formalized erasure inference and erasure reconstruction attacks, in which the attacker aims to either infer which record is deleted or reconstruct the erased sample. In particular, for the deletion inference, they formalize the goal of an erasure inference to distinguish a data example $e$ was in or not in the erased dataset, $e\in D_{e}$ or $e^{\prime}\notin D_{e}$. For the deletion reconstruction, they focused on reconstructing the erased example $e$. In all their reconstruction attacks, the attacker is not given any particular examples, and its goal is to extract knowledge about the features of the erased instance. Specifically, the deletion reconstructions include the deleted instance reconstruction and deleted label reconstruction. As named, the deleted instance reconstruction is to extract all of the features of the erased instance, and the erased label reconstruction is to infer the label of the erased point in the classification problem.
Zanella-Béguelin et al. (Zanella-Béguelin et al., 2020) indicated that the releasing snapshot of overlapped language models would leak the privacy of the training dataset. They verified that the model updates pose a significant risk to the private information added to or removed from the training dataset by a huge amount of experimental results. Zanella-Béguelin et al. found five phenomenons. First, an attacker can extract particular sentences used or not in the training dataset by comparing two models. Second, analysing more model snapshots shows more information about the updated data than considering fewer model snapshots. Third, adding or deleting other non-private data during model updates can not mitigate privacy leakage. Fourth, differential privacy can reduce privacy leakage risks, but it also decreases the accuracy of trained models. Fifth, to mitigate the privacy leakage risks while keeping the model utility, the server can limit the model parameters access or only output a subset of the results.
Many studies furthermore utilised these privacy threats to evaluate the unlearning effect. Huang et al. (Huang et al., 2021a) proposed Ensembled Membership Auditing (EMA) for auditing data erasure. They use the membership inference to evaluate the removing effect of unlearning. Graves et al. (Graves et al., 2021) indicated that if an attacker can infer the sensitive information that was wanted to be erased, then it means that the server has not guarded the user’s rights to be forgotten. Baumhauer et al. (Baumhauer et al., 2022) developed linear filtration to sanitise classification models with logits prediction after class-wide deletion requests. They verified their methods by testing how well it defends against privacy attacks.
Both (Chen et al., 2021b) and (Zanella-Béguelin et al., 2020) pointed out that differential privacy guarantees that a model does not reveal too much information about any particular training data point. A differentially private model can furthermore guarantee the group’s privacy by bounding the impacts of a group of training samples. If using DP to protect the privacy of a group of $|D\backslash D_{e}|$training samples against snapshot attacks on $\theta_{D},\theta_{D\backslash D_{e}}$, it means that $\theta_{D\backslash D_{e}}$ cannot be more useful than $\theta_{D}$.
8. Machine Unlearning Applications
Table 3. Machine Unlearning Application
Literature Application Scenarios Algorithms Realization Methods Evaluation Metric Year
(Huang et al., 2021b) Data unlearnable EMP Add noise ACC 2021
(Liu et al., 2022b) Mitigate backdoor BAERASER Gradient ascent method ASR and ACC 2022
(Wang et al., 2019) Mitigate backdoor - Median Absolute Deviation L1-Norm, ##\##FP, ACC, ASR 2019
(Du et al., 2019) Anomaly detection Unlearning Maintain memory set ##\##FP and ##\##FN 2019
(Cao et al., 2018) Repair pollution KARMA Cluster and unlearn DA 2018
  • ASR:Attack Success Rate, ACC: Accuracy
  • DA:Detection Accuracy, EMP: Error-minimizing Perturbations
  • ##\##FP: number of false positive, ##\##FN: number of false negative
Besides the inherent demands in machine unlearning that draw much research attention, many researchers also find that machine unlearning can be applied in many other scenarios to solve related problems. We list the recent unlearning applications in Table 3. The most popular application of machine unlearning is to mitigate the anomaly, including the backdoor triggers and pollution, from a trained model.
Wang et al. (Wang et al., 2019) applied machine unlearning in detecting and mitigating backdoor attacks in DNNs. They designed two approaches to eliminate backdoor-related neurons/weights from the infected model and repair the infected model to be robust against malicious images. Then, Liu et al. (Liu et al., 2022b) proposed that a backdoor model learned the poisoned decision boundary. Data points with triggers are all classified into the target class. They reversed the backdoor injection process to defend against it in machine unlearning, which is simple but effective. Their method contains two primary steps: first, they use a max-entropy staircase approximator to complete trigger reconstruction; second, they remove the added backdoor triggers using unlearning. They named these two key steps of BAERASER as trigger pattern recovery and trigger pattern unlearning. Via a dynamic penalty mechanism, they mitigated the sharp accuracy degradation of gradient-ascent-based machine unlearning methods.
Repairing pollution is another successful unlearning application. Cao et al. (Cao et al., 2018) proposed KARMA to search different subsets of training datasets and return the subset that with the highest misclassifications. KARMA includes three parts. First, KARMA searches for potential causality that leads to the wrong classification of a ML model. It clusters the misclassified data points into different parts and extracts the centres of clusters. Then KARMA prioritizes the search for similar data samples in the training datasets based on these extracted centres. Second, KARMA grows the causality found in the first stage by finding more training samples and forming a cluster. Third, KARMA determines if a causality cluster is polluted and calculates how many samples the cluster contains.
Machine unlearning was also applied to lifelong anomaly detection problems. Du et al. (Du et al., 2019) proposed new approaches to solve the challenges of exploding loss and sharp accuracy degradation caused by unlearning. In their presentation, they explain their techniques, emphasizing handling false negatives. When meeting such case $x_{t}$, their goal is, therefore, to modify the model to reduce the probability of $p(x_{t})$, or to increase the loss $\mathcal{L}(x_{t})$.
Huang et al. (Huang et al., 2021b) presented a method that can make training examples unlearnable by injecting a type of error-minimizing noise. This kind of noise is intentionally generated to reduce the error of the training examples close to zero, which can mislead the model into considering there is ”nothing” to learn from these examples. They first tried the error-maximizing (Goodfellow et al., 2014) noise but found that this noise cannot prevent DNN learning when applied in a sample-wise manner to the training data points. Therefore, they then begin to study the opposite direction of error-maximizing noise. In particular, they proposed the error-minimizing noise to stop the model from being punished by the loss function during traditional ML model training. Therefore, it can mislead the ML model to consider that there is ”nothing” to learn.
9. Lessons Learnt and Discussions
In the research domain of machine unlearning, researchers mainly face three difficult challenges, the stochasticity of training, incrementality of training, and catastrophe of unlearning, when they try to design efficient and effective unlearning mechanisms replacing naive retraining. They tried many mechanisms to mitigate the influence of these challenges. For example, in exact unlearning, they designed split algorithms that divide the final model into several sub-models and avoid the stochasticity and incrementality of one sub-model to influence other sub-models. In approximate unlearning, they bounded the removed estimation to aviod the accuracy degradation. However, it is still not easy to analyze them clearly and solve them totally. Moreover, as the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met new challenges. In this section, we will discuss the differences between various unlearning scenarios and their corresponding challenges. Based on these challenges, we list some potential research directions on which we can focus.
In the common centralized unlearning scenario, retraining from scratch can achieve the best unlearning effect, but it is expensive in both computation and storage. Existing methods try to design new unlearning mechanisms to reduce the cost from the two aspects. Although they proposed many methods, they just mitigated the influence of the main challenges and still have not solved them. Existing unlearning methods were at the beginning step that tried to implement effective unlearning. They proposed mechanisms, but most of them cannot guarantee the final results, so they also bound the unlearned item (Guo et al., 2020) or set an unlearning threshold (Nguyen et al., 2020). Although they also introduced many metrics to verify unlearning effect, the unlearning and verification processes are split. It means that the unlearning result is uncertain during the unlearning period before they finish the verification. Therefore, we find and list some open questions in centralized unlearning.
  1. (1)
    Optimizing the original three challenges. The first two challenges are hard to solve, so recent researches focus on reducing the utility degradation after unlearning. Even though many studies exist, there is still a long way to eliminate the unlearning degradation.
  2. (2)
    How to unlearn with a certain or exact goal? In other words, can we unlearn as learning? We know the purpose, to what extent we have unlearned and when we can finish or stop unlearning process.
Machine unlearning in distributed scenarios has many differences to which in centralized scenarios. We take federated unlearning as a representative example of distributed unlearning. The first difference is that federated unlearning can only be implemented locally on the client’s side if they want to unlearn some specific samples because clients’ do not upload their data to the FL server in a federated scenario. To avoid interacting with clients during unlearning, researchers (Liu et al., 2021; Wu et al., 2022b) proposed to unlearn the contribution of a whole client while not some samples of the client. The second difference is that when unlearning requests come during the FL training process, the FL server must first execute the unlearning process and broadcast the unlearned model for later updating to avoid other clients wasting computation on the before-unlearned model. The third difference is that federated unlearning is more vulnerable to catastrophic degradation than centralized unlearning because if the FL server broadcasts the catastrophic unlearned model and other clients update based on the unlearned model, it will vanish the efforts of other clients that trained before. After introducing these differences, we can see that the challenges in federated unlearning are more complex than in centralized unlearning, and we conclude the following open problems in federated unlearning.
  1. (1)
    Federated unlearning cannot use the fast retraining methods because data is out of reach for the server. Therefore, federated unlearning can only be implemented using approximate unlearning methods. However, as we know, approximate unlearning is easy to cause catastrophic unlearning, and federated learning is more vulnerable to degradation, so how to control the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
  2. (2)
    Privacy in federated learning is more sensitive than in other situations because FL was designed to protect participants’ privacy at the beginning. May unlearning bring other privacy demands in federated learning, such as secret or unnoticeable unlearning?
Besides exploring machine unlearning based on regularly structured data, researchers tried to implement unlearning in graph data. Exact unlearning may be suitable for centralized graph unlearning if graph data is sparse. However, the challenges of approximate unlearning may be more difficult than structured-data-based unlearning because, in graph unlearning, the relationship and influence between data samples are more complex than structured data (Chen et al., 2021a; Chien et al., 2022). In particular, graph data includes not only the node feature value but also the connecting edge information. Therefore, in graph unlearning, the estimation of the contribution of a node will be more difficult than in regular data unlearning. The original problems in regular data unlearning will be more challenging in graph unlearning. Besides these problems, graph unlearning also face unique problems that is related to edge structure information. In graph unlearning, how to unlearn some edges or a sub-graph is a big question.
Another important part is the privacy and security issues in machine unlearning. Although machine unlearning was proposed to protect users’ privacy when they want to leave, it brings new threats in that adversaries have a chance to infer the information about the removed data. Literature (Chen et al., 2021b; Zanella-Béguelin et al., 2020) have pointed out that updates of unlearning will leak privacy information, and they proposed corresponding attacks to infer this private information. However, most recent unlearning privacy leakage attacks are similar to those in a learning situation. An attack that is tailored to unlearning mechanisms is expected. A similar situation exists in unlearning applications. Although researchers proposed to unlearn a backdoor trigger (Liu et al., 2022b) or pollution (Cao et al., 2018), they only used a few unlearning techniques and paid more effort to detect those anomalies. One important reason is that unlearning mechanism is not mature enough. We are at the beginning of machine unlearning, and there are still many problems in unlearning itself that need to be solved. Finding machine unlearning applying situation and tailoring unlearning techniques to this situation is the direction of unlearning application.
10. Summary
The survey aims to offer a comprehensive and up-to-date overview of machine unlearning techniques in a systematic way. We organize the main challenges, research advancements, corresponding techniques, and privacy and security issues in machine unlearning. Additionally, we presented a detailed and unified classification of machine unlearning. We first illustrate the complete unlearning framework, including the learning, request, unlearning, and verification. Then, we catalogue recent studies briefly into exact unlearning and approximate unlearning, and we introduce the technical details of these unlearning methods. Moreover, we noticed some new unlearning scenarios, such as federated unlearning and graph unlearning, which are also introduced in the survey. Besides, we consider privacy and security issues in machine unlearning as an important consisting part. We collect and summarize the related literature about unlearning privacy threats and applications. Ultimately, the survey provides clear summaries and comparisons between various unlearning scenarios and corresponding methods, giving a comprehensive view of current work and listing the challenges and open problems of different scenarios of machine unlearning.
We hope our survey can help classify future unlearning studies, achieve a more in-depth understanding of unlearning methods, and address complex challenges. We believe the open problems listed in Section 9 will still be challenging in the following years, and we will try to optimize some of them. Last but not least, we expect this survey can help researchers in the study of machine unlearning, regardless of the unlearning strategies or unlearning privacy and security threats or applications of unlearning.

### v2

@@
 Abstract.
-As the right to be forgotten has been legislated worldwide, many studies attempt to design unlearning mechanisms to protect users’ privacy when they want to leave machine learning service platforms.
-Specifically, machine unlearning is to make a trained model to remove the contribution of an erased subset of the training dataset.
-This survey aims to systematically classify a wide range of machine unlearning and discuss their differences, connections and open problems.
-We categorize current unlearning methods into four scenarios: centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in unlearning.
-Since centralized unlearning is the primary domain, we use two parts to introduce: firstly, we classify centralized unlearning into exact unlearning and approximate unlearning; secondly, we offer a detailed introduction to the techniques of these methods.
-Besides the centralized unlearning, we notice some studies about distributed and irregular data unlearning and introduce federated unlearning and graph unlearning as the two representative directions.
-After introducing unlearning methods, we review studies about unlearning verification.
-Moreover, we consider the privacy and security issues essential in machine unlearning and organize the latest related literature.
-Finally, we discuss the challenges of various unlearning scenarios and address the potential research directions.
-machine unlearning
+As the right to be forgotten has been legislated worldwide, numerous studies have sought to design unlearning methods to protect users’ privacy when they want to remove their data from machine learning service platforms.
+This has given rise to the concept of “machine unlearning” — a field dedicated to removing the influence of specified samples from trained models.
+This survey aims to systematically classify a wide range of machine unlearning studies, discussing their differences, connections, and open problems.
+We categorize current unlearning methods into four key areas: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in unlearning.
+Given that centralized unlearning is the primary focus, we further divide its discussion into two parts: first, at a higher level, we classify centralized unlearning into exact unlearning and approximate unlearning; second, we provide a detailed introduction to the techniques used in these methods.
+Next, we introduce federated unlearning, another emerging area that has garnered significant attention.
+Following the discussion on unlearning methods, we review studies on unlearning verification and audit, which assess the effectiveness of existing unlearning algorithms.
+Additionally, we consider the privacy and security issues essential in machine unlearning and compile the related literature.
+Finally, we discuss the challenges of various unlearning scenarios and highlight potential research directions.
+Federated Unlearning; Machine Unlearning; Unlearning Verification
 ccs: Computing methodolgies Machine learningccs: Computing methodolgies Aritificial intelligenceccs: Security and privacyccs: Theory of computation
@@
 Machine learning (ML) algorithms and artificial intelligence (AI) are embedded into day-to-day applications and wearable devices (Mahdavinejad et al., 2018).
-It further collects more and more users’ information, including users’ private data, such as driving trajectory, medical and online shopping records.
-On the one hand, such an enormous amount of data helps to improve ML and AI development further.
-However, on the other hand, it threatens users’ privacy and demands the enormous requirement of data management to ensure information security and privacy in machine learning.
-Machine unlearning is recently proposed to protect users’ privacy when they want their data forgotten by a trained machine learning model.
-It draws increasing research attention because Right to be forgotten was legislated recently in many countries, such as the General Data Protection Regulation (GDPR) (Mantelero, 2013) in the European Union, PIPEDA privacy legislation in Canada (of the Privacy Commissioner of Canada, 2018), and the California Consumer Privacy Act (de la Torre, 2018) in the United States.
-The legislation requires that companies must take reasonable measures to achieve the erasure of personal data concerning.
-For a machine learning application, it means that individual users have the right to ask companies to delete their private data once collected from them.
-The deletion is not only erasing their data from a database.
-It needs to remove the contribution of the erased subset of whole training data from the already trained model, which process was first conceptualized as machine unlearning (Cao and Yang, 2015).
-Specifically, if a user (Alice) wants to exercise her right to be forgotten (Mantelero, 2013) (e.g., quitting a ML application), then the trained model of such application must ”unlearn” her data.
-Such a process includes two steps: first, a subset of the whole dataset previously used for training a ML model is requested to be erased; second, the ML model provider removes the contribution of the erased data from the trained model.
-A naive data-erasing method is retraining a new model from scratch (Bourtoule et al., 2021).
+It continuously collects increasing amounts of user information, including private data such as driving trajectories, medical records, and online shopping histories (Kirkpatrick et al., 2017; Rolnick et al., 2019).
+On the one hand, such an enormous amount of data helps to further advance ML and AI development.
+On the other hand, however, it poses a threat to users’ privacy and creates a significant need for robust data management to ensure information security and privacy in ML (Tian et al., 2022).
+Machine unlearning has drawn growing research attention as the recent legislation of the ”Right to be Forgotten” in many countries.
+Notable instances include the GDPR in the European Union (Mantelero, 2013), the PIPEDA privacy legislation in Canada (of the Privacy Commissioner of Canada, ), and the California Consumer Privacy Act in the United States (de la Torre, 2018).
+According to these laws, companies must take reasonable measures to guarantee that personal data is deleted upon request.
+It indicates that individual users have the right to request companies to remove their private data, which was previously collected for ML model training.
+The deletion is not only erasing their data from a database, but it also needs to delete the influence of the specified samples from trained models.
+The process of data removal from models was first conceptualized as machine unlearning (Cao and Yang, 2015).
+Specifically, suppose a user (Alice) wants to exercise her right (Mantelero, 2013) when quitting a ML application, then the trained model of such application must ”unlearn” her data.
+Such a process includes two steps: first, a subset of the dataset previously used for ML model training is requested to be deleted; second, the ML model provider erases the contribution of these data from the trained models.
+A naive data-erasing method is retraining a new ML model from scratch (Bourtoule et al., 2021).
 However, the computation and storage costs of retraining are expensive, especially in complex learning tasks.
-Many researchers tried to find efficient and effective methods to realize unlearning rather than naive retraining, and there are several common challenges, which are summarized as follows.
-Stochasticity of training: A huge amount of randomness exists in the training process in machine learning, especially in complicated models’ training periods such as CNNs (Gu et al., 2018) and DNNs (Sze et al., 2017).
-This randomness makes the training result non-deterministic (Bourtoule et al., 2021) and raises difficulties for machine unlearning to estimate the contribution of the specific erased data sample.
-Incrementality of training: The training process in machine learning is an incremental procedure.
-In other words, the model update on the former data point will influence the contribution of the data points fed into the model after this data.
-Determining a way to effectively remove the influence of the to-be-erased data samples from the trained model is challenging for machine unlearning (Koh and Liang, 2017).
-Catastrophe of unlearning: Nguyen et al.
-(Nguyen et al., 2020) indicated that an unlearned model usually performs worse than the model retrained from scratch.
-Especially when an unlearning method tries to remove a huge amount of data samples, the degradation could be exponential.
+Many researchers have tried to find efficient and effective methods to implement unlearning rather than naive retraining, and there are several common challenges, which are summarized as follows.
+(1) Stochasticity of training: A huge amount of randomness exists in the training process in machine learning, especially in complicated models’ training periods such as CNNs (Gu et al., 2018) and DNNs (Sze et al., 2017).
+This randomness makes the training results non-deterministic (Bourtoule et al., 2021) and raises challenges for machine unlearning to estimate the impact of the typical erased samples.
+(2) Incrementality of training: The training process in machine learning is incremental, meaning that the model update from one data point influences the contribution of subsequent data points fed into the model.
+Deciding a way to effectively remove the contributions of the to-be-erased samples from the trained model is challenging for machine unlearning (Koh and Liang, 2017).
+(3) Catastrophe of unlearning: Nguyen et al.
+(Nguyen et al., 2020) indicated that an unlearned model typically has worse model utility than the model retrained from scratch.
+The degradation would be exponential, especially when a method tries to delete a huge amount of data samples.
 They referred to such sharp degradation as catastrophic unlearning (Nguyen et al., 2020).
-Although several studies have explored ways to mitigate unlearning degradation by bounding the loss function or restricting unlearning threshold, how to eliminate catastrophic unlearning is still an open problem.
-Recently, many studies put efforts into solving these three main challenges and proposed many novel mechanisms which promote the progress of machine unlearning.
+Although several studies mitigate model utility degradation by bounding the loss function or restricting the unlearning update threshold, eliminating catastrophic unlearning remains an open problem.
+Recently, many studies have put efforts into solving these three main challenges and proposed many novel mechanisms that promote the progress of machine unlearning.
 […]
 ARC
-Machine UnlearningCentralized UnlearningUnlearning CategoriesExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2021; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021) Unlearning TechniquesSplit Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022)Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2021)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Distributed and Irregular Data UnlearningFederated Unlearning: (Wang et al., 2022; Liu et al., 2021, 2022a; Gupta et al., 2021; Wu et al., 2022b)Graph Unlearning: (Chien et al., 2022; Chen et al., 2021a; Cong and Mahdavi, 2022)Unlearning Verification: (Hu et al., 2022; Thudi et al., 2022; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a) Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021b; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021b; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
+Machine UnlearningCentralized UnlearningUnlearning CategoriesExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a; Schelter et al., 2023; Koch and Soll, 2023; Hu et al., 2024a)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021; Tarun et al., 2023; Wang et al., 2023a; Chundawat et al., 2023; Warnecke et al., 2024; Liu et al., 2023a; Lin et al., 2023; Cha et al., 2024) Unlearning TechniquesSplit Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a)Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Graph Unlearning: (Chien et al., 2022a; Chen et al., 2022b; Li et al., 2024; Cong and Mahdavi, 2022; Wang et al., 2023b; Wu et al., 2023a; Zhang et al., 2024; Wu et al., 2023b)Federated (Distributed) Unlearning: (Wang et al., 2022a, 2024; Fraboni et al., 2024; Wang et al., 2023c; Liu et al., 2021, 2022a; Zhang et al., 2023a; Gupta et al., 2021; Wu et al., 2022b)Unlearning Verification: (Hu et al., 2022; Guo et al., 2023; Thudi et al., 2022a; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a; Thudi et al., 2022b) Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b; Zhang et al., 2023b)Security Threats in Unlearning : (Marchant et al., 2022; Zhao et al., 2024; Di et al., 2022; Hu et al., ; Liu et al., 2024)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Wei et al., 2023; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
 Figure 1.
@@
 The introduction order will also follow this figure.
-We classify current unlearning literature into four main scenarios: centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning.
-The goal of this survey is to classify and systematize machine unlearning according to the reseaching problems and goals in unlearning processes, and to review their differences and connections, advantages and disadvantages.
-We first catalogue unlearning methods into four main parts, centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning, as shown in Fig.
+We classify the current unlearning literature into four main scenarios: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in machine unlearning.
+This survey aims to classify and systematize machine unlearning methods based on the research problems and objectives in the unlearning process and to review their differences and connections, as well as their advantages and disadvantages.
+We first categorize unlearning approaches into four main areas: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in machine unlearning, as shown in Fig.
 1.
-Since there are many studies targeted at centralized unlearning, we use two sections to introduce related work in this scenario.
-We briefly classify the related unlearning methods into two main kinds, Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022; Nguyen et al., 2020), where exact unlearning extends the idea of naive unlearning to design efficient retraining scheme with less computation cost.
-Approximate unlearning aims to unlearn a posterior that is approximate to the model retrained on the remaining dataset, only using the original learned model and the erased data.
-We then detailedly introduce the popular implementing techniques of these two unlearning kinds.
-In addition to these centralized unlearning methods, we introduce the distributed and irregular data unlearning scenarios, including federated unlearning (Wang et al., 2022; Liu et al., 2021), and graph unlearning (Chien et al., 2022; Chen et al., 2021a).
-After implementing unlearning, verifying the unlearning effect is equally important and draws lots of research attention.
-Therefore, we review the studies about unlearning verification subsequently.
-Moreover, we consider the privacy and security issues essential in machine unlearning, and we organize and review the related unlearning literature about privacy threats, defences, and unlearning applications in the survey.
-There is no comprehensive survey about machine unlearning because it is a pretty new research domain.
-Only several reviews on the specific unlearning scenarios were published on Arxiv.
-An introduction to machine unlearning, about the exact and approximate unlearning problems and how they are defined and solved by recently proposed methods, see (Mercuri et al., 2022).
-About certifiable machine unlearning for linear models, the algorithm introduction and experiments analysis can be found in (Mahadevan and Mathioudakis, 2021).
-Nguyen et al.
-(Nguyen et al., 2022b) summarized the general unlearning framework and added the unlearning verification part to it.
-They paid most of their attention to introducing problem formulations and technical definitions.
-None of them noticed and reviewed the privacy and security issues in machine unlearning.
-Moreover, they have not comprehensively discussed the differences and connections between different unlearning scenarios.
-In our treatment, we review not only centralized machine unlearning but also other scenarios, including distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning.
-We discuss these unlearning scenarios in the literature and compare the difference and connections from each other.
-For all four unlearning scenarios, we list open questions that, if solved, would increase our understanding of machine unlearning or enhance our ability to design a suitable unlearning mechanism for different applications.
-The main contributions of the paper are as follows.
-  • We systematically catalogue machine unlearning studies according to their scenarios, which can be briefly classified into, centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning.
-  • For the centralized unlearning scenario, which draws the most unlearning research attention, we further divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
-  • We further explore the privacy threats that machine unlearning brings and discuss the applications of unlearning in defending against traditional security and privacy issues.
-  • We discuss different challenges between various machine unlearning scenarios, from centralized unlearning to federated unlearning, graph unlearning, and privacy and security issues in unlearning.
+Since many studies focus on centralized unlearning, we dedicate two sections to introducing corresponding studies in this area.
+We classify these unlearning methods into: Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022a; Nguyen et al., 2020), where exact unlearning extends the idea of naive unlearning to design efficient retraining scheme with less computation cost.
+Approximate unlearning tries to unlearn a model approximate to the model retrained using the remaining dataset.
+We also provide a detailed introduction to the popular techniques used to implement these two types of unlearning.
+In addition to centralized unlearning methods, we introduce federated (distributed) unlearning studies (Wang et al., 2022a; Liu et al., 2021), which are gaining increasing importance.
+After discussing unlearning implementation methods, verifying the unlearning effect is equally crucial and has garnered significant research attention (Thudi et al., 2022b).
+Therefore, we subsequently review related research on unlearning verification.
+Finally, we consider privacy and security issues in machine unlearning essential, organizing and reviewing the related publications on privacy and security threats, defenses, and unlearning applications.
+There are only a few surveys on machine unlearning because it is a relatively new research domain, and most of them are preprints on Arxiv, focusing on certain unlearning scenarios.
+For an introduction to machine unlearning, including discussions on exact and approximate unlearning problems and their solutions through recently proposed methods, see (Mercuri et al., 2022).
+For information on provable machine unlearning for linear models, including algorithm introductions and experimental analysis, refer to (Mahadevan and Mathioudakis, 2021).
+For an overview of federated unlearning, see (Liu et al., 2023b; Wang et al., 2023d), and for graph unlearning, see (Said et al., 2023).
+While Nguyen et al.
+(Nguyen et al., 2022b) summarized the general unlearning framework and added the unlearning verification part to it, they focused primarily on introducing problem formulations and technical definitions.
+However, none of these surveys reviewed the privacy and security issues in machine unlearning, nor did they comprehensively discuss the differences and connections between various unlearning scenarios.
+Compared with existing surveys on machine unlearning, we performed a systematic literature review (SLR) following the guidelines presented by (Keele et al., 2007) and as exemplified in (Kubrak et al., 2022).
+The contributions of this paper include:
+  • We systematically catalog machine unlearning studies according to their scenarios, which can be briefly classified into centralized unlearning, distributed unlearning, unlearning verification, and privacy and security issues in machine unlearning.
+  • For the centralized unlearning scenario, which draws the most attention, we divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
+  • We further explore the privacy and security threats that target machine unlearning and discuss the applications of unlearning in defending against traditional security and privacy issues.
+  • We discuss different challenges between various machine unlearning areas, from centralized unlearning to federated unlearning, unlearning verification, and privacy and security issues in unlearning.
 We list the related open questions of each scenario and present potential research directions to solve them.
-The rest of the survey is organized as follows.
-The background of machine unlearning and the related technical tools are summarized in Section 2.
-After introducing the background, we will introduce the survey following the taxonomy order in Fig.
+The survey’s remaining sections are arranged as follows.
+We show a systematic literature review model of the survey in Section 2.
+The basic knowledge of machine unlearning and related technical tools are summarized in Section 3.
+After introducing the background, we introduce the survey following the taxonomy order in Fig.
 1.
-In Section 3, we present the primary categories of centralized unlearning: exact unlearning and approximate unlearning.
-In Section 4, techniques of centralized unlearning in different categories introduced in Section 3 are discussed in-depth.
-We collect and introduce the distributed and irregular data unlearning scenarios that we noticed, which are federated unlearning and graph unlearning, in Section 5.
+The primary content of centralized unlearning is presented in Section 4, which includes an introduction from two main unlearning categories and corresponding in-depth techniques.
+We collect and introduce the federated unlearning scenarios in Section 5.
 Section 6 introduces the unlearning evaluation and verification methods.
 The privacy and security issues in machine unlearning are divided into two sections.
-Section 7 introduces the privacy threats accompanied by machine unlearning.
-In Section 8, we discuss machine unlearning applications, which are mainly applied in dealing with security issues.
+Section 7 introduces the privacy and security threats accompanied by machine unlearning.
+Section 8 discusses machine unlearning applications, which are mainly applied in dealing with security issues.
 In Section 9, we discuss the challenges of current unlearning methods in a general way.
-After that, we list the open gaps that new studies can target.
-At last, we summarise the survey in Section 10.
+After that, we enumerate the unaddressed issues for further research.
+At last, in Section 10, we provide a summary of the survey.
+2. The SLR Method of the Survey
+We performed a systematic literature review (SLR) of the extant research on machine unlearning.
+The main goal was to compile a large body of research on machine unlearning and categorize them from various perspectives to facilitate analysis.
+Therefore, we employ the SLR method because it is especially appropriate for finding pertinent literature on a certain research topic (Keele et al., 2007).
+The main process of the SLR method for the survey is presented as follows.
+We first designed a search string according to the review protocol (Keele et al., 2007), identified appropriate digital databases, and defined the data extraction strategy.
+We focused on the keywords “unlearning” and “machine unlearning” in the search string, formulated as “unlearning OR machine unlearning”.
+We used this search string in IEEE Xplore, ACM Digital Library, Scopus, and the Web of Science to find relevant papers.
+Additionally, we conducted a search on Arxiv to identify further relevant literature.
+Consequently, a total of 972 papers were retrieved (the search was conducted on 7 July 2024).
+We then limited the publication years to those after 2020 and ensured that the keyword “unlearning” appeared in the title and abstract.
+Moreover, after filtering out duplicates and papers with fewer than six pages, 261 papers remained.
+Referring to the Google Scholar top publications and the China Computer Federation (CCF) recommendations lists, we focused on reviewing 103 papers from top venues.
+After including 33 references for related techniques, we reviewed a total of 136 references.
+During the systematic literature review process, the data extraction strategy is defined as follows.
+Besides the basic information of papers (title, author, publication venue, year), we first categorize papers into the four main classes as Fig.
+1 according to the research problems.
+Then, we extract the techniques utilized, the evaluating metrics, and the datasets employed in these papers, which will all be introduced in the following context.
 3. Background
 3.1. Machine Unlearning
-There are many urgent demands that motivate the hot machine unlearning research.
-The first is the privacy-preserving demand.
-Since the right to be forgotten has been legislated worldwide, machine unlearning is often formulated as a privacy-preserving problem where users can ask for their data erasure from the ML models.
-Besides privacy reasons, many other reasons promote the development of machine unlearning.
-For example, one common reason is usability.
-During the long time usage period of a recommendation application, users’ information will sometimes change, such as moving changes in their location.
-The application will produce unsuitable recommendations if it cannot completely delete the incorrect data from the model.
-Another reason is security.
-Deep learning models are vulnerable to adversarial attacks (Yuan et al., 2019) such as data poisoning (Tian et al., 2022).
-Hence, how to detect and remove adversarial poisoned data is essential to ensure models’ security.
-Once the attack type is detected, the model must erase these adversarial data’s influence using the unlearning mechanism.
-Figure 2. Machine unlearning process.
+There are several urgent demands driving machine unlearning research.
+The foremost is the demand for privacy preservation.
+With the “right to be forgotten” being legislated globally (Mantelero, 2013; of the Privacy Commissioner of Canada, ), machine unlearning ensures that users can request the erasure of their data from trained ML models.
+In addition to privacy concerns, other factors are promoting the development of machine unlearning.
+One significant factor is model utility.
+In the real world, vast amounts of data are generated daily, necessitating prompt updates to model services, as outdated data can negatively impact model performance (Wang et al., 2022b; Rolnick et al., 2019; Lopez-Paz and Ranzato, 2017).
+An effective unlearning mechanism is crucial for mitigating the adverse effects of outdated and incorrect data on model utility.
+Another critical factor is security.
+Adversarial attacks and data poisoning (Tian et al., 2022) can easily compromise deep learning models.
+Therefore, detecting and removing adversarial and poisoned data is essential to ensure model security.
+Once an attack type is identified, the model must erase the influence of these adversarial data using the unlearning mechanism.
 To better understand how unlearning mechanisms work, we first introduce the unlearning problem and process following the machine unlearning framework demonstrated in Fig.
-2.
-Let $\mathcal{Z}$ denote as a space of data items, then the particular (full) training dataset is $D\in\mathcal{Z}$.
+2, and we summarize the notations in Table 1.
+Let $\mathcal{Z}$ denote as a space of data items; the particular (full) training dataset is $D\in\mathcal{Z}$.
 A learning process can be demonstrated as Step 0 in Fig.
 2, i.e., training a model $M$ using an algorithm $\mathcal{A}$ on the training dataset $D$, denoted as $M=\mathcal{A}(D)$, where model $M$ is in a hypothesis space $\mathcal{H}$.
-The unlearning process begins at an unlearning request when a user wants to remove his data $D_{e}$ from the trained model, Step 1 in Fig.
+The unlearning process begins at an unlearning request when a user wants to erase his specified data $D_{e}$ from the trained model, Step 1 in Fig.
 2.
-The requested unlearning data $D_{e}$ can be data samples, classes or graph nodes.
+The requested unlearning data $D_{e}$ can be data samples (Bourtoule et al., 2021), classes (Warnecke et al., 2023), or graph nodes (Chen et al., 2022b).
 Then, in Step 2, the server removes the contribution of $D_{e}$ using a machine unlearning algorithm $\mathcal{U}$.
-The unlearned model can be expressed as $M_{D\backslash D_{e}}=\mathcal{U}(D,D_{e},\mathcal{A}(D))$.
-In dealing with an unlearning request, we hope that the unlearned model $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ is the same as the retrained model $\mathcal{A}(D\backslash D_{e})$ (i.e., $\mathcal{U}(D,D_{e},\mathcal{A}(D))\simeq\mathcal{A}(D\backslash D_{e})$).
+The unlearned model can be described as $M_{D\setminus D_{e}}=\mathcal{U}(M,D,D_{e})$.
+When dealing with an unlearning request, we hope that the unlearned model $\mathcal{U}(M,D,D_{e})$ is the same as the retrained model $\mathcal{A}(D\setminus D_{e})$ (i.e., $\mathcal{U}(M,D,D_{e})\simeq\mathcal{A}(D\setminus D_{e})$).
 Most unlearning studies ended at this step.
-However, we do need an effective evaluation metric to assess the unlearning effect as Step 3 in Fig.2.
-Therefore, we add the verification as the last step in the unlearning process, and we will introduce the related literature later.
-2.2. Unlearning Evaluation Metrics
-To compare the two models before and after unlearning, we need to define an evaluating metric $d(.)$ between $\mathcal{A}(D\backslash D_{e})$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ from the model parameter space $\mathcal{H}$.
-To this end, we briefly introduce several common evaluation distances metrics:
+However, we do need an effective evaluation metric to assess the unlearning effect, as in Step 3 in Fig.
+2.
+Therefore, we add the verification step as the last step in the unlearning process, and we will introduce the related literature later.
+Figure 2. Machine Unlearning Process
+Table 1. Notations in Machine Unlearning
+Symbols Description Symbols Description
+$\mathcal{Z}$ data items space $\mathcal{H}$ model parameters space
+$D=(X,Y)$ the training dataset with inputs $X$ and labels $Y$ $\mathcal{A}(\cdot)$ the ML algorithm
+$D_{e}=(X_{e},Y_{e})$ the unlearned (erased) dataset $\mathcal{U}(\cdot)$ the unlearning algorithm
+$D_{r}=D\setminus D_{e}$ the remaining dataset $M$ the model with parameters $\theta$
+$D_{\text{probe}}$ the probing data used in attacking $M_{D\backslash D_{e}}$ the unlearned model
+3.2. Machine Unlearning Evaluation Metrics
+To compare the two models before and after unlearning, we need to define an evaluating metric $d(\cdot)$ between $\mathcal{A}(D\setminus D_{e})$ and $\mathcal{U}(M,D,D_{e})$.
+To this end, we briefly introduce several common evaluation distance metrics:
 $L_{2}$-Norm.
-Wu et al.
-(Wu et al., 2020) proposed to utilise the Euclidean norm to evaluate the parameters of $\mathcal{A}(D\backslash D_{e})$ and that of unlearning $\mathcal{U}(D,D_{e},\mathcal{A}(D))$.
-Suppose $\theta$ are the model parameters learned by algorithm $\mathcal{A}(.)$.
-The $l_{2}$-norm or distance can be denoted as $||\theta_{\mathcal{A}(D\backslash D_{e})}-\theta_{\mathcal{U}(D,D_{e},\mathcal{A}(D))}||$, where $\theta_{\mathcal{A}(D\backslash D_{e})}$ is the model parameters of retraining from the scratch and $\theta_{\mathcal{U}(D,D_{e},\mathcal{A}(D))}$ is the modal parameters of unlearning algorithms $\mathcal{U}(.)$.
+In (Wu et al., 2020), the authors propose utilizing the Euclidean distance to evaluate the parameters of the retrained model and the unlearned model.
+Let $\theta$ represent the model parameters learned by the algorithm $\mathcal{A}(\cdot)$.
+The $L_{2}$-norm measures the distance between $\theta_{\mathcal{A}(D\setminus D_{e})}$ and $\theta_{\mathcal{U}(M,D,D_{e})}$, where $\theta_{\mathcal{A}(D\setminus D_{e})}$ are the model parameters retrained from scratch, and $\theta_{\mathcal{U}(M,D,D_{e})}$ are the model parameters resulting from the unlearning algorithm $\mathcal{U}(\cdot)$.
 Kullback–Leibler divergence (KLD).
-In Bayes-based or Markov chain Monte Carlo-based unlearning methods, researchers used the Kullback–Leibler divergence or Jensen-Shannon divergence to optimize the approximate posterior since they can measure the distance of two probability distributions.
-Therefore, in recent unlearning studies, researchers also used the KLD between the retrained and unlearned models to estimate the unlearning effect (Nguyen et al., 2020).
-For discrete probability distritbutions of two model parameters $\mathcal{A}(D\backslash D_{e})$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ on the same probability space $\mathcal{H}$, the KLD from $\mathcal{A}(D\backslash D_{e})$ to $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ is denoted as $\text{KL}[\Pr(\mathcal{A}(D\backslash D_{e}))||\Pr(\mathcal{U}(D,D_{e},\mathcal{A}(D)))]$.
-Parameter Leakage.
-Since membership inference attacks can infer if a sample is used in the training dataset of a model, recently, some works have leveraged the property to verify if the specific data is removed by unlearning mechanisms.
-Some studies (Hu et al., 2022) even proposed to backdoor the erased samples and then attack the unlearned model to check if the unlearned model can be attacked.
-If the unlearned model is backdoored, it proves that the unlearning algorithm cannot unlearn samples effectively.
+KLD is commonly used to measure the divergence between two probability distributions, often assessing the distance between retrained and unlearned models.
+In Bayes-based or Markov chain Monte Carlo-based unlearning methods (Nguyen et al., 2020), researchers utilize KLD (Joyce, 2011) to optimize approximate models, employing it to measure the distance between two probability distributions.
+Recent unlearning studies have also used KLD to estimate the unlearning effect by comparing the distributions of retrained and unlearned models (Nguyen et al., 2020).
+Evaluation Metric based on Privacy Leakage.
+Since membership inference attacks (Chen et al., 2021) can decide whether a sample was utilized for training a model, recently, some works have leveraged this property to verify if unlearning mechanisms remove the specific data.
+Some studies (Hu et al., 2022) even proposed to backdoor the unlearning samples for initial model training and then attack the unlearned model.
+If the unlearned model is still backdoored, this proves that the unlearning algorithm cannot unlearn samples effectively.
 Conversely, if the backdoor trigger cannot attack the unlearned model, it proves that the unlearning algorithm is effective.
-Similar methods were also used in (Thudi et al., 2022; Ma et al., 2022) to evaluate the unlearning effect.
+Similar methods were also used in (Thudi et al., 2022a; Ma et al., 2022) to evaluate the unlearning effect.
 3.3. Majority of Tools Used in Unlearning
-2.3.1. Differential Privacy (DP)
-Differential privacy has become a popular standard for privacy protection in the statistic.
-In a DP model, a trusted analyzer collects users’ raw data and then executes a private mechanism to achieve differential privacy.
-The privacy requirement ensures the indistinguishability for any two outputs of neighbouring datasets, where neighbouring datasets mean the dataset only differs by replacing one user’s data, denoted as $X\backsimeq X^{\prime}$.
-A ($\epsilon,\delta$)-differential privacy algorithm $\mathcal{M}:\mathbb{X}^{n}\to\mathbb{Z}$ means that for every pair of neighboring datasets $X\backsimeq X^{\prime}\in\mathbb{X}^{n}$ and every subset $S\subset\mathbb{Z}$ has that $\mathcal{M}(X)\in S$ and $\mathcal{M}(X^{\prime})\in S$ are $\epsilon$-indistinguishable and $\delta$-approximate.
-The smaller the $\epsilon$ is, the more the privacy protection level increases.
-When the value of $\epsilon$ is 0, it implies that the outputted probability distribution of algorithm $\mathcal{M}$ cannot reflect any useful information about the datasets.
+Differential Privacy (DP).
+Differential privacy is a popular benchmark for privacy protection in the Statistic (Dwork et al., 2006).
+In a DP model, a trusted analyzer collects users’ raw data and then executes a private method to guarantee differential privacy.
+The DP protection ensures the indistinguishability for any two outputs of neighboring datasets, where neighboring datasets mean the dataset only differs by replacing one user’s data, denoted as $X\backsimeq X^{\prime}$.
+A ($\epsilon,\delta$)-differential privacy algorithm $\mathcal{M}:\mathbb{X}^{n}\to\mathbb{Z}$ means that for every neighboring dataset pair $X\backsimeq X^{\prime}\in\mathbb{X}^{n}$ and every subset $S\subset\mathbb{Z}$ has that $\mathcal{M}(X)\in S$ and $\mathcal{M}(X^{\prime})\in S$ are $\epsilon$-indistinguishable and $\delta$-approximate.
+The degree of privacy protection rises with decreasing $\epsilon$.
+When $\epsilon=0$, it implies that the outputted probability distribution of mechanism $\mathcal{M}$ cannot represent any meaningful information.
 A general DP mechanism based on adding Laplace noise was presented and theoretically analyzed in (Dwork et al., 2006).
-2.3.2. Bayesian Variational Inference (VI)
-In machine learning, the bayesian variational inference is used to approximate difficult-to-compute probability densities via optimization.
-In this part, we revisit the VI framework (Blei et al., 2017) that learns approximate posterior model parameters $\theta$ using Bayesian Theory (Bernardo and Smith, 2009).
-Given a prior belief $p(\theta)$ of an unknown model and a complete data trainset $D$, an approximate posterior belief $q(\theta|D)\sim p(\theta|D)$ can be optimized by minimizing the Kullback–Leibler divergence (KLD) (Joyce, 2011), $\text{KL}[q(\theta|D)||p(\theta|D)]$.
-The KLD is a type of statistical distance to measure how one probability distribution $q(\theta|D)$ differs from another probability distribution $p(\theta|D)$.
+Bayesian Variational Inference.
+In machine learning, the Bayesian variational inference is used to approximate difficult-to-compute probability densities via optimization (Kingma and Welling, 2014; Blei et al., 2017).
+We revisit the variational inference framework that learns approximate posterior model parameters $\theta$ using Bayesian Theory in this part.
+Suppose a prior belief $p(\theta)$ of an unidentified model and a complete data trainset $D$, an approximate posterior belief $q(\theta|D)\sim p(\theta|D)$ can be optimized by minimizing the KLD (Joyce, 2011), $\text{KL}[q(\theta|D)||p(\theta|D)]$.
+KLD measures how one probability distribution $q(\theta|D)$ differs from another probability distribution $p(\theta|D)$.
 However, it is intractable to compute the KLD exactly or minimize the KLD directly.
-Instead, the evidence lower bound (ELBO) (Blei et al., 2017) was proposed to be maximized, which is equivalent to minimize KLD between the two probability distributions.
-ELBO follows directly from $log(p(D))$ subtracting $\text{KL}[q(\theta|D)||p(\theta|D)]$, where the log marginal likelihood $log(p(D))$ is independent of $q(\theta|D)$.
-Since $\text{KL}[q(\theta|D)||p(D|\theta)]\geq 0$, the ELBO is a lower bound of $log(p(D))$.
-In general training situations, ELBO is maximized using stochastic gradient ascent (SGA).
-The primary process is approximating the expectation $\mathbb{E}_{q(\theta|D)}[log(p(D|\theta))+log(p(\theta)/q(\theta|D))]$ with stochastic sampling in each iteration of SGA.
+Instead, the evidence lower bound (ELBO) (Kingma and Welling, 2014) was proposed to be maximized, which is equivalent to minimize KLD between the two probability distributions.
+ELBO follows directly from $\log(p(D))$ subtracting $\text{KL}[q(\theta|D)||p(\theta|D)]$, where $\log(p(D))$ is independent of $q(\theta|D)$.
+The ELBO is a lower bound of $\log(p(D))$ as $\text{KL}[q(\theta|D)||p(D|\theta)]\geq 0$.
+In general training situations, ELBO is maximized using stochastic gradient ascent (SGA) (Kingma and Welling, 2014).
+The primary process is approximating the expectation $\mathbb{E}_{q(\theta|D)}[\log(p(D|\theta))+\log(p(\theta)/q(\theta|D))]$ with stochastic sampling in each iteration of SGA.
 We can use a simple distribution (e.g., the exponential family)to approximate computational ease posterior belief $q(\theta|D)$.
-When dealing with complex distribution whose density cannot be evaluated, we can maximize the ELBO with adversarial training by alternating between estimating the log-density ratio $log(p(\theta))/q(\theta|D)$ and maximizing the ELBO (Yu et al., 2019).
-2.3.3. Privacy Leakage
 Figure 3. Privacy Leakage: a Privacy Reconstruction Process
-Privacy leakage occurs in both unlearning verification and privacy threats in unlearning two parts.
-In unlearning verification, researchers tried to use privacy leakage attacks to verify if the specific data is unlearned.
-In the privacy and security issues in unlearning, researchers tried to design effective inference attacks tailored to machine unlearning.
-The basic attack of privacy leakage in a machine learning setting is membership inference, which determines if a data sample was used in the ML model updating process or not.
-When a sample is fully known to the adversary, learning which model was trained based on it will leak information about the model.
+Privacy Leakage Attacks.
+Privacy leakage occurs in both unlearning verification and privacy threats in two parts of unlearning.
+In unlearning verification, researchers tried to use privacy leakage attacks to verify whether the specific data is unlearned.
+Regarding the privacy and security issues in unlearning, researchers have tried to design effective inference attacks tailored to machine unlearning.
+The basic attack of privacy leakage in a machine learning setting is membership inference, which determines if a sample was employed in the model updating process or not.
+When an attacker fully knows a sample, knowing which model was trained on it will leak information about the model.
 A generic membership inference process was introduced in (Shokri et al., 2017).
 Shokri et al.
-first trained the shadow models to simulate the target machine-learning model.
+first trained the shadow models to approach the target ML models.
 Then, they observed and stored the different outputs of the shadow models based on different inputs, in or not, in the trainset.
@@
 Model inversion (Fredrikson et al., 2014), or privacy reconstruction (Salem et al., 2020) is another privacy threat in general machine learning.
-The goal of model inversion is to infer some missing attributes of input features based on the interaction with a trained machine learning model.
+Model inversion aims to infer some lacking attributes of input features based on the interaction with the trained ML model.
 Salem et al.
-(Salem et al., 2020) proposed a reconstruction attack target recovering specific data samples used in the model updating process by the different outputs of the model before updating and after updating.
-Latter, inferring the private information of updating data in conventional machine learning is transferred to infer the privacy of the erased samples in machine unlearning.
-In reconstruction attacks, the adversary first collects the different outputs using his probing data $D_{probe}$, including the original outputs $\hat{Y}_{\mathcal{M}}$ before unlearning, and the outputs $\hat{Y}_{\mathcal{M}_{-D_{e}}}$ after unlearning.
-Then he constructs the attack model based on the posterior difference $\delta=\hat{Y}_{\mathcal{M}_{-D_{e}}}-\hat{Y}_{\mathcal{M}}$.
+(Salem et al., 2020) proposed a reconstruction attack target recovering specific data samples used in the model updating by different model outputs before and after updating.
+Later, inferring the private information of updating data in conventional machine learning is transferred to inferring the privacy of the erased samples in machine unlearning.
+In reconstruction attacks, the adversary first collects the different outputs using his probing data $D_{\text{probe}}$, including the original outputs $\hat{Y}_{M}$ before unlearning, and the outputs $\hat{Y}_{M_{-D_{e}}}$ after unlearning.
+Then, he constructs the attack model based on the posterior difference $\delta=\hat{Y}_{M_{-D_{e}}}-\hat{Y}_{M}$.
 The attack model contains an encoder and decoder, which has a similar structure as VAEs (Kingma and Welling, 2014), and the main process is shown in Fig.
 3.
-3. Categories of Centralized Machine Unlearning
-In this section, we classify existing centralized unlearning methods by their inherent mechanism and designed purposes.
-Detailed techniques will be introduced in the next section.
+4. Centralized Machine Unlearning
+In this section, we classify existing centralized unlearning methods by their inherent mechanism and designed purposes and present the corresponding detailed techniques.
+4.1. Unlearning Solution Categories
 From the former introduction, we know that naive retraining is the most effective manner to realize machine unlearning.
-However, it is inefficient because it needs to store the whole training dataset and retrain a new model from scratch.
-It consumes a huge amount of storage and computation costs, especially in deep learning scenarios.
+However, it is inefficient because it requires storing the entire original dataset and retraining the model from scratch, which consumes significant storage and computational resources, especially in deep learning scenarios.
 Therefore, researchers tried to design effective and efficient unlearning mechanisms.
-However, there were difficult challenges when we tried other unlearning methods.
-Three representative challenges are the stochasticity of training, the incrementality of training, and the catastrophe of unlearning.
-Exact unlearning was proposed to reduce the computation cost by splitting training sub-models based on pre-divided data sub-sets.
+Exact unlearning was proposed to reduce the computation cost by splitting training sub-models based on pre-divided data sub-sets (Bourtoule et al., 2021).
 It also divides the stochasticity and the incrementality into sub-models.
 It can unlearn an exact model by ensembling the consisting submodels, but it still needs to store all the split data subsets for fast retraining.
-Another approach is approximate unlearning.
+Another approach is approximate unlearning (Sekhari et al., 2021).
 It can reduce both storage and computation costs because it unlearns only based on the erased data.
-However, it is difficult for approximate unlearning methods to control the accuracy degradation, due to the challenges in estimating stochasticity and incrementality during the training process.
+However, it is difficult for approximate unlearning methods to control the accuracy degradation due to the challenges in estimating stochasticity and incrementality during the training process.
 Most of them bounded their estimation to avoid dramatic accuracy reduction.
-3.1. Exact Unlearning
+4.1.1. Exact Unlearning
 Exact unlearning could also be called fast retraining, whose basic idea is derived from naive retraining from scratch.
-Following the background introduction of unlearning, we know the learning model $\mathcal{A}$ and unlearning model $\mathcal{U}$ based on the trainset $D$ and erased dataset $D_{e}\subseteq D$.
-If $\mathcal{U}(.)$ is implemented as naive retraining, the equality between $\mathcal{A}(D\backslash D_{e})\in\mathcal{H}$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))\in\mathcal{H}$ is absolutely guaranteed.
-However, naive retraining involves high computation and storage costs, especially for deep learning models and complex datasets (Thudi et al., 2022).
-Unlike naive retraining needs to retrain based on the whole remaining dataset, exact unlearning tries to retrain a sub-model based on the subset of remaining data to reduce computation cost.
-A general operation of exact unlearning is that they first divide the training dataset into several sub-sets.
-Then, they transform the learning process by ensembling the sub-models trained based on each sub-set as the final model.
-So that when unlearning request comes, they only need to retrain the trained sub-model based on the sub-set containing the erased data.
+Following the background introduction of unlearning, we know the learning and unlearning algorithm, $\mathcal{A}(D)$ and $\mathcal{U}$, based on the trainset $D$ and erased dataset $D_{e}\subseteq D$, respectively.
+If $\mathcal{U}(\cdot)$ is implemented as naive retraining, the equality between $\mathcal{A}(D\setminus D_{e})\in\mathcal{H}$ and $\mathcal{U}(M,D,D_{e})\in\mathcal{H}$ is absolutely guaranteed.
+However, naive retraining involves high computation and storage costs, especially for deep learning models and complex datasets (Thudi et al., 2022a).
+Unlike naive retraining, which relies on the whole remaining dataset, exact unlearning tries to retrain a sub-model only using a subset of the remaining dataset to reduce calculation cost.
+A general operation of exact unlearning is that they first divide the dataset into several small sub-sets.
+Then, they transform the learning process by ensembling the sub-models trained with each sub-set as the final model (Bourtoule et al., 2021; Cao and Yang, 2015).
+So that when an unlearning request comes, they are just required to retrain the sub-model corresponding to the sub-set containing the erased data.
 They then ensemble the retrained sub-model and other sub-models as the unlearned model.
-As we introduced above, we know that the primary goal of exact unlearning is to reduce the computation cost of retraining a new model by transforming the original learning algorithms into an ensembling form.
+As we introduced above, exact unlearning aims to mitigate the computation cost when retraining a new model by transforming the original learning algorithms into an ensembling form.
 It divides the stochasticity and incrementality into several sub-models to reduce their influence.
@@
 In (Cao and Yang, 2015), Cao and Yang transformed the traditional ML algorithms into a summation form.
-When an unlearning requirement comes, they only need to update several summations, ensuring the process runs faster than retraining from scratch.
-Bourtoule et al.
-proposed SISA (Bourtoule et al., 2021), which is short for Sharded, Isolated, Sliced, and Aggregated training.
-SISA divided the full data into shards and trained models separately in each shard.
+They are only required to update several summations when an unlearning requirement comes, ensuring the method runs faster than retraining from scratch.
+SISA (Bourtoule et al., 2021) is a representative exact unlearning algorithm, which splits the full training dataset into shards and trained models separately in each shard.
 For unlearning, they simply need to retrain the shard that includes the erased data.
-Wu et al.
-(Wu et al., 2022a) proposed a framework that precisely models the influence of each training data point on the model concerning various performance criteria and removes the influence of the data that is required to delete.
+Study (Wu et al., 2022a) proposed a framework that precisely models the impact of individual training sample on the model concerning various performance criteria and removes the impact of samples that are required to be removed.
 Golatkar et al.
 (Golatkar et al., 2021) proposed an unlearning method on deep networks, splitting the trained model into two parts.
-The core part based on the data will not delete, and the unlearning part on the erased data with parameters bound.
+The core part based on the data will not be deleted, and the unlearning part with the erased data will be unlearned with parameters bound.
 These methods are efficient in computation, but they sacrifice the storage space to store the intermediate training parameters of different slices and the related training sub-sets.
-Besides the huge storage cost, another big problem in exact unlearning is that it only suits the situation, in which the unlearning request only needs to remove a few samples with a low frequency.
+Besides the high storage cost, another major issue with exact unlearning is that it is only suitable for scenarios where the unlearning request involves removing a few samples with low frequency.
 Suppose an unlearning request needs to remove many data samples (usually, they are not in the same previous divided sub-set).
-In that case, exact unlearning must retrains all these related sub-models or even all the sub-models in the worst situation.
-At the moment, exact unlearning is no longer computation efficient, and it still needs to store the whole training dataset or intermediate parameters.
-3.2. Approximate Unlearning
+In that case, exact unlearning must retrain all these related sub-models or even all the sub-models in the worst situation.
+At the moment, exact unlearning is no longer computation efficient, and the whole training dataset and intermediate parameters still need to be stored.
+4.1.2. Approximate Unlearning
 Figure 4.
-Model changes when adding a new point or removing a point.
+The model changes when adding a new point or removing a point.
 (a) A normally trained classifying model classifies classes 1 and 2.
 (b) When a new point appears, the model is trained based on it, and the classifying line is pushed to classify it.
-(c) When we need to remove an influential point, we should recover the contribution of this point to the model.
+(c) When we need to remove an influential point, we should recover the contribution of this data point on the model.
 (d) When we remove a Non-influential point, the model may not need to change a lot.
-Unlike exact unlearning could only reduce the retraining computation cost, approximate unlearning tries to directly unlearn based on the learned model and the erased data sample, which saves the computation and storage costs together.
-The goal of approximate unlearning is to unlearn a posterior that approximates the model trained on the remaining dataset, based on only the erased sub-dataset while not the remaining dataset (used by exact unlearning).
-As we introduced before, the unlearning purpose is expecting that the unlearned model $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ to equal to the retrained model $\mathcal{A}(D\backslash D_{e})$.
+Unlike exact unlearning, which only aims to reduce the retraining computation cost, approximate unlearning tries to directly unlearn based on the trained model and the erased data sample, which saves the computation and storage costs together.
+Approximate unlearning studies aim to unlearn a model that approaches the model trained on the remaining dataset, i.e., the unlearned model $\mathcal{U}(M,D,D_{e})$ should match the retrained model $\mathcal{A}(D\setminus D_{e})$.
 Since exact unlearning is implemented by retraining from the remaining dataset or sub-sets, they can almost guarantee equality before and after unlearning.
-However, since approximate unlearning tries to directly delete the contribution of the erased data from the trained model, how to exactly estimate the contribution of the erased data (the contribution contains the stochasticity and the incrementality ) and remove it becomes the core problem.
+However, since approximate unlearning tries to directly delete the influence of the unlearned samples from trained models, the core problem lies in precisely estimating and removing this contribution, which includes both stochasticity and incrementality.
 The text description of the changes between two different distribution spaces before and after removing the specific data is not intuitive.
@@
 4 (b).
-When this influential point is requested to remove, the unlearning mechanism must recover the model to the original one that has not been trained by this specific point, as shown in Fig.
+When this influential point is requested to be removed, the unlearning mechanism must recover the model to the original one that has not been trained by this specific point, as shown in Fig.
 4 (c).
@@
 Many methods were proposed to implement approximate unlearning efficiently and effectively.
-The popular solutions are certified-removal and Bayes-based mechanisms, which will be introduced in technical detail in Section 4.
+The popular solutions are certified-removal (Guo et al., 2020) and Bayes-based mechanisms (Nguyen et al., 2020), which are introduced in technical detail in Section 4.2.
 Here, we briefly summarize current studies about approximate unlearning.
@@
 The machine learning model does not learn anything from the training dataset if $\mathcal{A}$ is differentially private (Bourtoule et al., 2021).
-Sekhari et al.
-(Sekhari et al., 2021) introduced a similar solution to (Guo et al., 2020), which removes the contribution of the erased data from the gradients during the unlearning process.
-To mitigate the unlearning accuracy reduction, they bounded the model parameters by their differential unlearning definition.
-Neel et al.
-(Neel et al., 2021) proposed unlearning methods allowing arbitrary sequences of updates, including addition and deletion, and defined the corresponding weak and strong unlearning algorithms.
-Different from indistinguishable unlearning, other methods used the Bayes theorem to unlearn an approximate posterior directly.
+A similar solution to (Guo et al., 2020) is introduced in (Sekhari et al., 2021), which erases the influence of the specified samples from the gradients during unlearning.
+To mitigate accuracy degradation, model parameter updates are constrained by a differential unlearning definition.
+Different from indistinguishable unlearning, other methods used the Bayes theorem to unlearn an approximate posterior.
 In (Nguyen et al., 2020), Nguyen et al.
-unlearned an approximate posterior based on the erased data via the Bayesian inference.
-Similarly, Fu et al.
-(Fu et al., 2022) proposed a machine unlearning algorithm for Markov chain Monte Carlo (MCMC) and further defined a knowledge removal estimator the evaluate the knowledge erasing.
-Nguyen et al.
-(Nguyen et al., 2022a) also explained the effectiveness of the approximate machine unlearning algorithm from the perspective of Markov Chain Monte Carlo.
-However, since those methods are approximately unlearning the contribution of all the erasing data, including the inputs and labels, it is easy to decrease the model accuracy to some extent after unlearning.
-4. Techniques of Centralized Unlearning
-Table 1. Summary of popular unlearning techniques
+employed Bayesian inference to unlearn an approximate model using the erased data.
+Moreover, Fu et al.
+(Fu et al., 2022) developed a similar unlearning method, which uses Markov chain Monte Carlo (MCMC) methods.
+In (Nguyen et al., 2022a), the authors also explained the effectiveness of the approximate unlearning method from the perspective of MCMC.
+Nevertheless, since those techniques are approximately unlearning the contribution of all the erasing data, including the inputs and labels, they inevitably decrease the model accuracy to some extent after unlearning.
+Main Challenges of Approximate Unlearning.
+As we discussed at the beginning, there are three main challenges in machine unlearning.
+In centralized scenarios, these studies mainly aim at solving the basic machine unlearning problem, which unavoidably faces three challenges: stochasticity of training, incrementality of training, and catastrophe of unlearning.
+The exact unlearning methods extend the retraining idea, which avoids facing these challenges but consumes lots of storage costs.
+The approximate unlearning methods face these challenges directly, and we here list the relevant work about how to estimate the contribution of erased samples to overcome the stochasticity and incrementality of training, and how to prevent unlearning catastrophe.
+  • To overcome the stochasticity and incrementality challenges when estimating the unlearning influence, one popular strategy is based on the first-order and second-order influence function (Basu et al., 2020), which is calculated based on the perturbation theory (Avrachenkov et al., 2013).
+  • The unlearning catastrophe appears commonly in approximate unlearning, and many studies try to propose some methods to solve this problem.
+In certified removal and Bayesian-based methods, they usually set a threshold to limit the unlearning update extent (Guo et al., 2020; Nguyen et al., 2020).
+In (Wang et al., 2023a), Wang et al.
+solves this problem by adding a model utility compensation task during unlearning optimization and finding the optimal balance based on multi-objective training methods.
+4.2. Detailed Techniques of Centralized Unlearning
+Table 2. Popular Unlearning Techniques
 Unlearning Literature Taxonomy Requests Type Techniques Realization Method Year
-SISA (Bourtoule et al., 2021) Exact unlearning Samples Split learning Data and Model Partition 2021
-GraphEraser (Chen et al., 2021a) Exact unlearning Graph nodes Split learning Data Partition 2021
-Amnesiac unl. (Graves et al., 2021) Exact unlearning Samples Split learning Partially retraining 2021
-RecEraser (Chen et al., 2022) Exact unlearning Samples Split learning Balanced Data Partition 2022
-ARCANE (Yan et al., 2022) Exact unlearning Samples Split learning Parition by Class 2022
-HedgeCut (Schelter et al., 2021) Exact unlearning Samples Split learning Tree ensemble learning 2021
-DeltaGrad (Wu et al., 2020) Exact unlearning Samples Split learning L-BFGS (Byrd et al., 1994) 2020
+SISA (Bourtoule et al., 2021) Exact unlearning Samples Split Unlearning Data and Model Partition 2021
+Amnesiac unl. (Graves et al., 2021) Exact unlearning Samples Split Unlearning Partially retraining 2021
+GraphEraser (Chen et al., 2022b) Exact unlearning Graph nodes Split Unlearning Data Partition 2022
+RecEraser (Chen et al., 2022a) Exact unlearning Samples Split Unlearning Balanced Data Partition 2022
+ARCANE (Yan et al., 2022) Exact unlearning Samples Split Unlearning Parition by Class 2022
+HedgeCut (Schelter et al., 2021) Exact unlearning Samples Split Unlearning Tree ensemble learning 2021
+DeltaGrad (Wu et al., 2020) Exact unlearning Samples Split Unlearning L-BFGS (Byrd et al., 1994) 2020
+ERASER (Hu et al., 2024a) Exact unlearning Samples Split Unlearning Inference Serving-Aware 2024
 L-CODEC (Mehta et al., 2022) Approximate unlearning Samples Certified Data Removal Markov Blanket selection 2022
@@
 $(\epsilon,\delta)$-unl. (Neel et al., 2021) Approximate unlearning Samples Certified Data Removal Perturbed gradient descent 2021
-(Ginart et al., 2019) Approximate unlearning Samples Certified Data Removal Quantized and DC-k-means 2019
-Graph unl. (Chien et al., 2022) Approximate unlearning Graph nodes Certified Data Removal Certified removal 2022
-EUBO, rKL (Nguyen et al., 2020) Approximate unlearning Samples Bayesian-based Unlearning VBI 2020
-MCU (Nguyen et al., 2022a) Approximate unlearning Samples Bayesian-based Unlearning Monte cario-based 2022
-BIF (Fu et al., 2022) Approximate unlearning Samples Bayesian-based Unlearning MCMC 2022
+(Warnecke et al., 2023) Approximate unlearning Samples Certified Data Removal Influence Theory 2023
+Graph unl. (Chien et al., 2022a) Approximate unlearning Graph nodes Certified Data Removal Certified removal 2022
+Gif (Wu et al., 2023b) Approximate unlearning Graph nodes Certified Data Removal Influence function 2023
+SUMMIT (Zhang et al., 2024) Approximate unlearning Graph nodes Certified Data Removal Multi-Objective Optimization 2024
+EUBO, rKL (Nguyen et al., 2020) Approximate unlearning Samples Bayesian Unlearning VBI 2020
+MCU (Nguyen et al., 2022a) Approximate unlearning Samples Bayesian Unlearning Monte cario-based 2022
+BIF (Fu et al., 2022) Approximate unlearning Samples Bayesian Unlearning MCMC 2022
   • Alogrihtm abbreviations
-  • VBI: Variational Bayesian Inference, SME: Store medial estimation, MCU: Monte Cario-based machine unlearning, FIM: Fisher Information Matrix, MCMC: Markov chain Monte Carlo, BIF: Bayesian inference forgetting, LP: Loss perturbation (Chaudhuri et al., 2011), TF-IDF: Term Frequency Inverse Document Frequency, EUBO: Evidence upper bound, rKL: reverse Kullback–Leibler
-The naive retraining method is the most effective way to realize machine unlearning.
-However, it is expensive in both computation and storage.
+  • VBI: Variational Bayesian Inference, SME: Store medial estimation, MCU: Monte Cario-based machine unlearning, FIM: Fisher Information Matrix, MCMC: Markov chain Monte Carlo, BIF: Bayesian inference forgetting, LP: Loss perturbation, TF-IDF: Term Frequency Inverse Document Frequency, EUBO: Evidence upper bound, rKL: reverse Kullback–Leibler
+This section presents the popular techniques used in existing unlearning methods, in both exact and approximate unlearning studies.
 Exact unlearning extends the idea of retraining and tries to reduce the computation cost of unlearning.
 Approximate unlearning was proposed to find a way to reduce computation and storage consumption together.
-It is usually realized by unlearning based on the erased data only.
-In this section, we will introduce the popular techniques used in existing unlearning methods, in both exact unlearning and approximate unlearning.
-The dominant studies are summarized in Table 1, where the primary technique used in exact unlearning is split learning, and two primary techniques used in approximate unlearning are certified data removal and bayesian-based unlearning, respectively.
-The introduction order of these techniques will mainly follow the order of the unlearning techniques in Fig.
+The dominant studies are summarized in Table 2, where the primary technique used in exact unlearning is split learning.
+Two primary techniques used in approximate unlearning are certified data removal and Bayesian-based unlearning.
+The introduction of these techniques will primarily follow the sequence of the unlearning techniques presented in Fig.
 1.
-4.1. Split Unlearning
-Since most of the exact unlearning methods tried to split the full dataset into multi sub-sets and divide the model training process, we call this kind of unlearning technique split unlearning.
-The main process of split unlearning is illustrated in Fig.
+4.2.1. Split Unlearning
+Figure 5.
+(a) Naive unlearning.
+There are only two steps: delete the specified samples from the whole dataset and retrain a model based on the remaining dataset.
+(b) Split unlearning.
+It contains four steps: 1.
+split the original dataset into $n$ shards, 2.
+remove the erased data from the corresponding shard, 3.
+retrain the sub-model of this shard, 4.
+ensemble all sub-models as the final model.
+Since most exact unlearning methods attempted to partition the training dataset into multiple subsets and divide the ML model learning process, we call this kind of unlearning technique split unlearning.
+The main procedure of split unlearning is illustrated in Fig.
 5 (b).
-By contrast, we show the process of the naive retraining method in Fig.
-5 (a).
-From Fig.
-5 (a), we can see that when we unlearn the erased data by naive unlearning, we need two steps to realize it: first, remove the data from the original training dataset; second, retrain a new model based on the remaining dataset.
+By contrast, the process of the naive retraining method is shown in Fig.
+5 (a), where we need two steps to realize it: first, delete the samples from the training dataset; second, retrain a new model using the remaining dataset.
 Since it needs to store the whole training dataset and retrain from the remaining dataset, it often entails significant computational and stored overhead.
-Researchers saw the huge computation cost of naive retraining.
-They proposed many exact unlearning methods, and we summary as split unlearning because they were all based on data or model split techniques.
+They proposed many exact unlearning methods to reduce the huge computation cost of naive retraining, and the majority of them are based on split learning techniques, either on data or model.
 As shown in Fig.
 5 (b), the split unlearning technique can be summarized into four steps.
-Different from naive unlearning trained based on the whole remaining trainset, the first phase of split unlearning is dividing the original full trainset into multiple disjoint shards.
+Unlike naive unlearning trained based on the remaining trainset, the first phase of split unlearning is dividing the original full trainset into multiple disjoint shards.
 All the constituent models are trained based on each split data slice.
-Then, in the second phase, when the unlearning request comes, they only need to remove the requested data from the split slice and retrain this slice’s constituent model in the third phase.
+Then, in the second phase, when the unlearning request comes, they only need to erase the requested samples from the split slice and retrain this slice’s constituent model in the third phase.
 In the last phase, the split unlearning aggregates the retrained and other constituent models together as a new unlearned model.
@@
 It can efficiently speed up retraining time and reduce computation costs.
-In (Cao and Yang, 2015), the authors indicated that support vector machines, naive Bayes classifiers, k-means clustering, and many machine learning algorithms could be implemented in a summation form to reduce the retraining cost.
+In (Cao and Yang, 2015), the authors indicated that support vector machines, naive Bayes classifiers, k-means clustering, and many ML algorithms could be implemented in a summation form to reduce the retraining cost.
 The statistical query (SQ) learning (Kearns, 1998) guarantees the summation form.
-Although algorithms in Probably Approximately Correct (PAC) setting can be converted to the SQ learning setting, a huge amount of complex models, such as DNNs, cannot be converted to SQ learning efficiently.
-Figure 5.
-(a) Naive unlearning.
-Only two steps: remove the erased data from the whole dataset and retrain the model based on the remaining dataset.
-(b) Split unlearning.
-It contains four steps: 1.
-split the original dataset into $n$ shards, 2.
-remove the erased data from the corresponding shard, 3.
-retrain the sub-model of this shard, 4.
-ensemble all sub-models as the final model.
+Although algorithms in the Probably Approximately Correct (PAC) setting can transform to the SQ learning setting, many complex models, such as DNNs, cannot be efficiently converted to SQ learning.
 Then, Bourtoule et al.
@@
 named their unlearning method the SISA training approach.
-SISA was implemented on deep neural networks, including two Conv.
-layers followed by two FC layers, Wide ResNet-1-1 and ResNet-50 models on multiple datasets, including MNIST, Purchase, SVHN, CIFAR100, Imagenet, and Mini-Imagenet.
+SISA can be implemented on deep neural networks, training multiple sub-neural networks based on divided sub-datasets.
+When the unlearning request comes, SISA retrains the model of the shard, which contains the information about the erased samples.
+SISA is effective and efficient as it aggregates all sub-models final prediction results rather than aggregates all these models.
 Unlike the original split unlearning dividing the dataset and transforming learning algorithms to summation form, Yan et al.
 proposed ARCANE (Yan et al., 2022), which transforms conventional ML into ensembling multiple one-class classification tasks.
-When many unlearning requests come, it can reduce retraining costs, which is not considered by previous work.
+When many unlearning requests come, it can reduce retraining costs, which was not considered in previous work.
 Chen et al.
-extended exact unlearning methods to recommendation tasks and proposed RecEraser (Chen et al., 2022), which has similar architecture as split unlearning in Fig.
+(Chen et al., 2022a) extended exact unlearning methods to recommendation tasks and proposed RecEraser, which has similar architecture as split unlearning in Fig.
 5 (b).
 RecEraser is tailored to recommendation systems, which can efficiently implement unlearning.
-Specifically, they designed three data partition schemes to split recommendation data into balanced pieces and propose an adaptive aggregation method based on an attention mechanism.
-They conducted the experimental evaluation on three real-world datasets, YELP2018, Movielens-1m, and Movielens-10m.
-These three datasets are widely used to assess the effectiveness and efficiency of recommendation models.
-Schelter et al.
+Specifically, they designed three data division schemes to partition recommendation data into balanced pieces and created an adaptive aggregation algorithm utilizing an attention mechanism.
+They conducted the experiments on representative real-world datasets, which are usually employed to assess the effectiveness and efficiency of recommendation models.
+Besides the above popular ML models, Schelter et al.
 proposed HedgeCut (Schelter et al., 2021), which implemented machine unlearning on tree-based ML models in a split unlearning similar form.
-Tree-based learning algorithms are trained by recursively partitioning the training data, locally optimizing a criterion such as Gini gain (Wehenkel and Pavella, 1991).
+Tree-based learning algorithms are developed by recursively partitioning the training dataset, locally optimizing a metric such as Gini gain (Wehenkel and Pavella, 1991).
 HedgeCut focuses on implementing fast retraining for these methods.
@@
 The intuitive idea of Amnesiac Unlearning is to store the parameters of training batches and then subtract them when unlearning requests appear.
-In particular, it first trains the learning model by adding the total gradients $\sum_{e=1}^{E}\sum_{b=1}^{B}\nabla_{\theta_{e,b}}$ to the initial model parameters $\theta_{initial}$, where $E$ is the training epochs, and $B$ is the data batches.
-During the training process, they kept a list of $SB$, which records the batches containing the private data.
-The list could be in the form of an index of batches for each training example, an index of batches for each training category or any other information expected.
+In particular, it first trains the learning model by adding the total gradients $\sum_{e=1}^{E}\sum_{b=1}^{B}\nabla_{\theta_{e,b}}$ to the initial model parameters $\theta_{\text{initial}}$, where $E$ is the training epochs, and $B$ is the data batches.
+In the model training process, they kept a list called $SB$, which records the batches holding the private data.
+This list could be formed as an index of batches for each training example, an index of batches for each category or any other information expected.
 When the unlearning request comes, a model using Amnesiac unlearning needs only to remove the updates from each batch $sb\in SB$ from the learned model $\theta_{M}$.
 As Graves et al.
-mentioned, using Amnesiac unlearning effectively and efficiently removes traces of the erased data that could be inferred via state-of-the-art privacy inference attacks and does not degrade the accuracy of the model in any other way.
-4.2. Certified Data Removal
-Certified data removal unlearning methods usually define their unlearning algorithms as $\epsilon$-indistinguishable unlearning, which is similar to the differential privacy definition.
+(Graves et al., 2021) stated, using Amnesiac unlearning effectively and efficiently removes the contribution of the erased samples that could be detected through state-of-the-art privacy inference attacks and does not degrade the accuracy of the model in any other way.
+4.2.2. Certified Data Removal
+Figure 6.
+The approximate unlearning by certified data removal.
+The unlearning algorithm $\mathcal{U}$ includes estimating the influence of specified data and removing the estimation from trained models.
+The unlearned model is expected to approach the retrained model.
+Certified data removal unlearning methods usually define their unlearning algorithms as $\epsilon$-indistinguishable unlearning, which is similar to the differential privacy definition (Dwork et al., 2014).
+An example is presented in Figure 6.
 Most of them use the Hessian matrix (Bishop, 1992) to evaluate the contribution of erased data samples for unlearning subtraction.
-After estimating the contribution of the erased data samples, they unlearn by subtracting these contributions with an updating bound from the unlearning model.
+After estimating the impact of the erased data samples, they unlearn by subtracting these impacts with an updating bound from the unlearning model.
 In (Guo et al., 2020), Guo et al.
-proposed a certified data removal mechanism, which assumes removing the last training sample, $(x_{n},y_{n})$.
+proposed a certified data removal method, which assumes removing the last training sample, $(x_{n},y_{n})$.
 Specifically, they defined a removal mechanism that approximately minimizes $\mathcal{L}(\theta;D^{\prime})$ with $D^{\prime}=D\backslash(x_{n},y_{n})$.
-The loss gradient at sample $(x_{n},y_{n})$ can be denoted as $\Delta=\lambda\theta^{*}+\nabla l((\theta^{*})^{T}x_{n},y_{n})$ and the Hessian of $L(\cdot;D^{\prime})$ at $\theta^{*}$ by $H_{\theta^{*}}=\nabla^{2}L(\theta^{*};D^{\prime})$.
-Then they applied a one-step Newton update to the gradient impact of the removed point $(x_{n},y_{n})$ on the model $\theta$.
-Under their observation, they found that directly removing the Hessian contribution from the direction of the gradient will leak the privacy of the training data that was expected to be removed.
+The loss gradient at sample $(x_{n},y_{n})$ can be denoted as $\Delta=\lambda\nabla\ell(\theta^{T}\cdot x_{n},y_{n})$ and the Hessian of $L(\cdot;D^{\prime})$ at $\theta$ by $H_{\theta}=\nabla^{2}L(\theta;D^{\prime})$.
+Then, they applied a one-step Newton update to the model parameters impact of the erased point $(x_{n},y_{n})$ on the model $\theta$.
+Under their observation, they found that directly removing the Hessian contribution from the gradient will reveal the private information of the erased data.
 They used the loss perturbation technique (Chaudhuri et al., 2011) to hide this information.
-It used a random linear term to perturb the empirical risk and ensures that the outputs of their method $\mathcal{U}(D,D_{e},\mathcal{A})$ is $\epsilon$-indistinguishable between the retrained model $\mathcal{A}(D\backslash D_{e})$.
-In(Mehta et al., 2022) and (Thudi et al., 2022), they designed unlearning algorithms following the certified data removal definition in (Guo et al., 2020).
+It used a random linear term to perturb the empirical risk and ensure that the outputs of their method $\mathcal{U}(D,D_{e},\mathcal{A})$ is $\epsilon$-indistinguishable between the retrained model $\mathcal{A}(D\backslash D_{e})$.
+In(Mehta et al., 2022) and (Thudi et al., 2022a), they designed unlearning algorithms following the certified data removal definition in (Guo et al., 2020).
 Mehta et al.
-unlearned via their proposed efficient Hessians, L-FOCI (Mehta et al., 2022).
+(Mehta et al., 2022) unlearned via their proposed efficient Hessians, L-FOCI (Mehta et al., 2022).
 Thudi et al.
-(Thudi et al., 2022) used membership inference as a verification error to adjust the unlearning process on stochastic gradient descent (SGD) optimization.
+(Thudi et al., 2022a) used membership inference as a verification error to adjust the unlearning process on stochastic gradient descent (SGD) optimization.
 Another similar unlearning method is PUMA.
@@
 proposed a new data removal method through gradient re-weighting called PUMA, which also used the Hessian Vector Product (HVP) term.
-They first estimated and recorded individual contributions of $(x_{i},y_{i})$, where the estimation is limited to less than a dot product between the pre-cached HVP term and individual gradient.
-When the unlearning request comes, they subtract the estimation of the erased samples to update the model parameters.
+They first estimated and recorded individual contributions of $(x_{i},y_{i})$, where the estimation is limited to less than one dot product between the pre-cached HVP term and individual gradient.
+When the unlearning request comes, they subtract the estimate of the erased samples to revise the model.
 Ginart et al.
 (Ginart et al., 2019) extended certified data removal to k-means clustering algorithms.
-They formulated the unlearning problem of efficiently removing personal data information from trained clustering models and proposed two different efficient deletions for k-means clustering, quantized k-mean and divide-and-conquer k-means.
+They formulated the unlearning problem of efficiently removing personal data information from trained clustering models.
+They offered two different deletions for k-means clustering, quantized k-mean and divide-and-conquer k-means.
 In their work, both algorithms have theoretical guarantees and strong empirical results.
-DeltaGrad (Wu et al., 2020), an algorithm for fast retraining of SGD-based models, was proposed by Wu et al.
-to small changes of data, based on the idea of ”differentiating the optimization path” concerning the data and Quasi-Newton methods.
+To retrain SGD-based models fast, DeltaGrad was proposed by Wu et al.
+(Wu et al., 2020) to unlearn small changes of data inspired by the idea of ”differentiating the optimization path” concerning the training dataset and Quasi-Newton methods.
 They theoretically proved that their algorithm could approximate the right optimization path rapidly for the strongly convex objective.
-DeltaGrad starts with a ”burn-in” period of first iterations, where it computes the full gradients exactly.
+DeltaGrad starts with a ”burn-in” period of first iterations, where it computes the full gradients precisely.
 After that, it only calculates the complete gradients for every first iteration.
 For other iterating rounds, it operates the L-BGFS algorithm (Byrd et al., 1994) to compute Quasi-Hessians approximating the true Hessians, keeping a set of updates at some prior iterations.
-Sekhari et al.
+For a deeper understanding of certified machine unlearning, Sekhari et al.
 (Sekhari et al., 2021) further given a strict separation between $\epsilon$-indistinguishable unlearning and differential privacy.
-Different from (Guo et al., 2020), in order to utilise tools of differential privacy (DP) for ML, the most straightforward manner is to ignore the special set of erasure demands $D_{e}$ and create an unlearning mechanism $\mathcal{U}$ that only relies on the learned algorithm $\mathcal{A}(D)$.
-In particular, the unlearning algorithm is of the form $\mathcal{U}(D_{e},\mathcal{A}(D))=\mathcal{U}(\mathcal{A}(D))$ and makes sure the true unlearned model $\mathcal{U}(\mathcal{A}(D))$ is $\epsilon$-indistinguishable to $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$.
+Different from (Guo et al., 2020), in order to utilize tools of differential privacy (DP) for ML, the most straightforward manner is to forget the special dataset of erasure demands $D_{e}$ and create an unlearning mechanism $\mathcal{U}$ that solely relies on the learned algorithm $\mathcal{A}(D)$.
+In particular, the unlearning method is of the form $\mathcal{U}(D_{e},\mathcal{A}(D))=\mathcal{U}(\mathcal{A}(D))$ and makes sure the true unlearned model $\mathcal{U}(\mathcal{A}(D))$ is $\epsilon$-indistinguishable to $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$.
 Notice the difference between (Guo et al., 2020) and (Sekhari et al., 2021).
 In the definition of (Guo et al., 2020), their $\epsilon$-indistinguishable unlearning is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{A}(D\backslash D_{e})$, but here is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$.
-Such a pair of algorithms in (Sekhari et al., 2021) would be differential private for $D$, where the neighbouring datasets mean that for two datasets with an edit distance of $m$ samples.
-The guarantee of DP unlearning is stronger than the model distribution undistinguishable unlearning guarantee in (Guo et al., 2020), and thus it suffices to satisfy it.
-Based on the definition of (Sekhari et al., 2021), they pointed out that any DP algorithm automatically unlearns any $m$ data samples if it is private for datasets with edit distance $m$.
+Such a pair of algorithms in (Sekhari et al., 2021) would be differential private for $D$, where the neighboring datasets mean that for two datasets with an edit distance of $m$ samples.
+The guarantee of DP unlearning is more powerful than the model distribution undistinguishable unlearning in (Guo et al., 2020), and therefore, it suffices to satisfy it.
+Based on the definition of (Sekhari et al., 2021), they pointed out that any DP algorithm automatically unlearns any $m$ data samples if they are private for datasets with the distance $m$.
 Therefore, they derivate the bound on deletion capacity from the standard performance guarantees for DP learning.
 Furthermore, they determine that the existing unlearning algorithms can delete up to $\frac{n}{d^{1/4}}$ samples meanwhile still maintaining the performance guarantee w.r.t.
-the test loss, where $n$ is the size of the original trainset and $d$ is the dimension of trainset inputs.
-The above methods are all answering the problem of unlearning.
+the test loss, where $n$ is the size of the original trainset, and $d$ is the dimension of trainset inputs.
+The aforementioned methods are trying to address the basic unlearning problem from a certified data removal perspective.
 Neel et al.
-(Neel et al., 2021) extended the unlearning definition to the update, which includes both ’add’, $D\cup\{z\}$, and ’delete’ as $D\backslash\{z\}$, where $z\in\mathcal{Z}$ is a data point.
-They follow the definition in (Guo et al., 2020) and define the similar $(\epsilon,\delta)$-indistinguishability.
-Moreover, they extend $(\epsilon,\delta)$-indistinguishability for both ’add’ and ’delete’ updates, which can also be denoted as $(\epsilon,\delta)$-publishing.
-4.3. Bayesian-based Unlearning
-Different from certified data removal that unlearns samples by subtracting corresponding Hessian matrix estimation from trained models, Bayesian-based unlearning tries to optimize an approximate posterior of the remaining dataset only using the trained model (the posterior of the original full dataset) and the erased data.
-The exact Bayesian unlearning posterior can be derivated from the Bayesian rule as $p(\theta|D_{r})=p(\theta|D)\ p(D_{e}|D_{r})/p(D_{e}|\theta)$, where $\theta$ is the posterior (i.e., model parameters).
-$D$ is the original full trainset, and $D_{r}$ is the remaining dataset which removes the erased data $D_{e}$, denoted as $D=D_{e}\cup D_{r}$ and $D_{e}\cap D_{r}=\emptyset$.
-If the model parameters $\theta$ are discrete-valued, $p(\theta|D_{r})$ can be directly obtained from the Bayesian rule.
-The use of a conjugate prior furthermore makes unlearning relatively simple.
-However, in practice, it is not easy to achieve the exact posterior, not to mention the unlearning posterior.
+(Neel et al., 2021) extended the definition of unlearning to include updates, which encompass both ”add” ($D\cup\{z\}$) and ”delete” ($D\backslash\{z\}$), where $z=(x,y)$ and $z\in\mathcal{Z}$ is a data point.
+They follow the definition in (Guo et al., 2020) and define similar $(\epsilon,\delta)$-indistinguishability.
+Furthermore, they extend $(\epsilon,\delta)$-indistinguishability to both ”add” and ”delete” updates, which can also be denoted as $(\epsilon,\delta)$-publishing.
+4.2.3. Bayesian-based Unlearning
+Different from certified data removal that unlearns samples by subtracting corresponding Hessian matrix estimation from trained models, Bayesian-based unlearning tries to unlearn an approximate posterior as the model is trained by employing the remaining dataset.
+The exact Bayesian unlearning posterior can be derived from the Bayesian rule as $p(\theta|D_{r})=p(\theta|D)\ p(D_{e}|D_{r})/p(D_{e}|\theta)$, where $\theta$ is the posterior (i.e., model parameters).
+The erased dataset and the remaining dataset are two independent subsets of the full training dataset.
+If the model parameters $\theta$ are discrete-valued, $p(\theta|D_{r})$ can be directly obtained from the Bayesian rule (Nguyen et al., 2020).
+Additionally, employing a conjugate prior simplifies the unlearning process.
+Nevertheless, it is challenging to get the exact posterior in practice, not to mention the unlearning posterior.
 In (Nguyen et al., 2020), Nguyen et al.
 tried doing likewise at the beginning.
-They defined the loss function using the KLD between the approximate predictive distribution $q_{u}(y|D_{r})$ and the exact predictive distribution $p(y|D_{r})$ where the observation $y$ (i.e., drawn from a model with parameters $\theta$) is conditionally independent of $D_{r}$ given $\theta$.
+They defined the loss function using the KLD between the approximate predictive distribution $q_{u}(y|D_{r})$ and the exact predictive distribution $p(y|D_{r})$.
 They bounded this loss function by the KLD between posterior beliefs $q_{u}(\theta|D_{r})$ and $p(\theta|D_{r})$ and further proposed evidence upper bound (EUBO) as the loss function to unlearn the approximate unlearning posterior.
@@
 Toward this purpose, Nguyen et al.
-proposed a Markov chain Monte Cario-based machine unlearning (MCU) method deriving from the Bayesian rule.
-They experimentally proved that MCU could unlearn the erased subsets of the whole training dataset from a prepared model effectively and efficiently.
+(Nguyen et al., 2022a) proposed an MCMC-based machine unlearning method deriving from the Bayesian rule.
+They experimentally proved that MCMC-based unlearning could effectively and efficiently unlearn the erased subsets of the whole training dataset from a prepared model.
 Fu et al.
-(Fu et al., 2022) converted the Markov chain Monte Carlo unlearning problem into an explicit optimization problem.
-Then they proposed $\epsilon-$Knowledge removal, which was a little similar to certified removal methods, but they defined that the KLD between the unlearned and the retrained model must be less than $\epsilon$.
-To quantify the explicit $\epsilon-$knowledge removal, they proposed a knowledge removal estimator to evaluate the difference between the original and unlearned distributions.
+(Fu et al., 2022) converted the MCMC unlearning problem into an explicit optimization problem.
+Then they proposed $\epsilon-$knowledge removal, which was a little similar to certified removal methods, but they defined that KLD between unlearned and retrained models must be less than $\epsilon$.
+To quantify the explicit $\epsilon-$knowledge removal, they proposed a knowledge removal estimator to assess the difference between the original and unlearned distributions.
 As they indicated, though their algorithm cannot wholly remove the learned knowledge from the already trained distribution, their method can still help the unlearned model approach its local minimum.
-5. Distributed and Irregular Data Unlearning
-We have introduced popular centralized machine unlearning in the former two sections.
-Most of them are requested to delete the regularly structured data.
-Their main purpose is to reduce the computation and storage costs and the accuracy degradation caused by unlearning.
-Recently, we noticed that some researchers paid attention to realizing unlearning in other scenarios, such as in distributed and irregular data learning scenarios.
-In this section, we will introduce the literature about federated unlearning and graph unlearning as two representative directions.
-5.1. Federated Unlearning
-Federated learning (McMahan et al., 2016, 2017) is first proposed to protect the participating users’ privacy during a machine learning training process in distributed situations.
-All participants will only upload their locally trained model parameters instead of their sensitive local data to the server for the model training.
-Therefore, in a federated learning scenario, the limited access to the dataset will become a unique challenge to implement unlearning.
-Since the local data cannot be uploaded to the FL server side, most federated unlearning methods tried to unlearn a certain client’s influence from the trained model by storing and estimating the contribution of uploaded parameters.
-In this situation, they can implement federated unlearning without interacting with the client.
+Figure 7.
+(a) Regular data unlearning, which only unlearns the data sample.
+(b) Graph unlearning.
+Since the graph data contains both node features and connective edge information, we may need to unlearn edge information or both edges and nodes in graph unlearning.
+4.2.4. Graph Unlearning
+We introduce graph unlearning as a representative kind of irregular data unlearning.
+In (Chien et al., 2022a; Chen et al., 2022b; Cong and Mahdavi, 2022; Zhang et al., 2024), researchers extend regular data machine unlearning to a graph data scenario.
+Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig.
+7.
+Therefore, Chien et al.
+(Chien et al., 2022a) proposed node unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC).
+Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees.
+They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
+Chen et al.
+(Chen et al., 2022b) found that applying SISA (Bourtoule et al., 2021) unlearning methods to graph data learning will severely harm the graph-structured information, resulting in model utility degradation.
+Therefore, they proposed a method called GraphEraser to implement unlearning tailored to graph data.
+Similar to SISA, they first cut off some connecting edges to split the total graph into some sub-graphs.
+Then, they trained the constituent graph models on these sub-graphs and ensembled them for the final prediction task.
+To realize graph unlearning efficiently, they proposed two graph partition algorithms and corresponding aggregation methods based on them.
+In (Cong and Mahdavi, 2022), Cong and Mahdavi filled in the gap between regularly structured data unlearning and graph data unlearning by studying the unlearning problem on the linear-GNN.
+To remove the knowledge of a specified node, they design a projection-based unlearning approach, PROJECTOR, that projects the weights of the pre-trained model onto a subspace irrelevant to the deleted node features.
+PROJECTOR could overcome the challenges caused by node dependency and is guaranteed to unlearn the deleted node features from the pre-trained model.
+5. Federated Unlearning
+We have introduced popular centralized machine unlearning in the former section.
+Their main purpose is to reduce the calculation and storage costs and the accuracy degradation caused by unlearning.
+Recently, we noticed that some researchers paid attention to realizing unlearning in other scenarios, such as in federated learning (FL) scenarios.
+In this section, we will introduce the literature about federated unlearning.
+FL was initially introduced to protect the privacy of participating clients during the machine learning training process in distributed settings.
+All participants will only upload their locally trained model parameters instead of their sensitive local data to the FL server during model training processes.
+Therefore, in a federated learning scenario, limited access to the dataset will become a unique challenge when implementing unlearning.
+Since the local data cannot be uploaded to the federated learning (FL) server side, most federated unlearning methods try to erase a certain client’s contribution from the trained model by storing and estimating the contribution of uploaded parameters.
+In this situation, they can implement federated unlearning without interacting with the client, shown as the server-side federated unlearning in Fig.
+8 (a).
 The two representative methods are (Liu et al., 2021; Wu et al., 2022b).
 Liu et al.
-(Liu et al., 2021) proposed FedEraser to sanitize the contribution of a federated client on the global federated learning model.
-In particular, during the FL global training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates.
-Based on these operations, they reconstructed the unlearned global model instead of retraining a new model from scratch.
+(Liu et al., 2021) proposed “FedEraser” to sanitize the impact of a FL client on the global FL model.
+In particular, during FL training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates.
+Based on these operations, they reconstructed the unlearned FL model instead of retraining a new model from scratch.
 However, FedEraser can only unlearn one client’s data, which means it must unlearn all the contributions of this specific client’s data.
 It is unsuitable for a client who wants to unlearn a small piece of his data.
-Wu et al.
-(Wu et al., 2022b) tried to unlearn a client’s influence from the federated global model by removing the historical updates from the global model.
+Study (Wu et al., 2022b) tried to erase a client’s influence from the FL model by removing the historical updates from the global model.
 They implemented federated unlearning by using knowledge distillation to restore the contribution of clients’ models, which does not need to rely on clients’ participation and any data restriction.
+Figure 8. Comparison between (a) server-side federated unlearning and (b) client-side federated unlearning
 Wang et al.
-(Wang et al., 2022) explored the problem of how to selectively unlearn a category from a trained CNN classification model in FL.
-The biggest challenge is that the local data used for FL global model’s training cannot be accessed globally.
-Therefore, they explored the inner influence of each channel in deep and observed that various channels have distinct impacts on different categories.
+(Wang et al., 2022a) explored the problem of how to selectively unlearn a category from a trained CNN classification model in FL.
+The biggest challenge is that the local data utilized for the FL global model training cannot be accessed globally.
+Therefore, they explored the inner influence of each channel and observed that various channels have distinct impacts on different categories.
 They proposed a method that does not require accessing the data used for training and retraining from scratch to cleanly scrub the information of particular categories from the global FL model.
-A method called Term Frequency Inverse Document Frequency (TF-IDF) was introduced to quantize the class discrimination of channels.
+A method called TF-IDF was introduced to quantize the class discrimination of channels.
 Then, they pruned the channels with high TF-IDF scores on the erased classifications to unlearn them.
@@
 Different from unlearning a whole client’s influence and unlearning a class, Liu et al.
-(Liu et al., 2022a) investigated how to unlearn data samples in FL.
-They first defined a federated unlearning problem and proposed a fast retraining method to remove data points from the federated learning global model.
-Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods (Xu et al., 2020) and the first-order Taylor approximate technique.
-They used the practical Fisher Information Matrix (FIM) (Ly et al., 2017) to model the Hessian matrix at a low cost.
+(Liu et al., 2022a; Wang et al., 2023c) investigated how to unlearn data samples in FL, shown as the client-side federated unlearning in Fig.
+8 (b).
+In (Liu et al., 2022a), they first defined a federated unlearning problem and proposed a fast retraining method to withdraw the influence of data from the FL model.
+Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods and the first-order Taylor approximate method (Liu et al., 2022a).
+They utilized the practical Fisher Information Matrix to model the Hessian matrix at a low cost.
 When unlearning, they subtract the estimation matrix from gradients with a bound.
-5.2. Graph Unlearning
-We introduce graph unlearning as a representative kind of irregular data unlearning.
-In (Chien et al., 2022; Chen et al., 2021a; Cong and Mahdavi, 2022), researchers extend regular data machine unlearning to a graph data scenario.
-Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig.
-6.
-Therefore, Chien et al.
-(Chien et al., 2022) proposed node feature unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC).
-Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees.
-They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
-Figure 6.
-(a) Regular data unlearning, which only unlearns the data sample.
-(b) Graph unlearning.
-Since the graph data contains both node features and connective edge information, we may need to unlearn only edges information or both edges and nodes in graph unlearning.
-Chen et al.
-(Chen et al., 2021a) found that applying SISA (Bourtoule et al., 2021) unlearning methods to graph data learning will severely harm the graph-structured information, resulting in model utility degradation.
-Therefore, they proposed GraphEraser to implement unlearning tailored to graph data.
-Similar to SISA, they first cut off some connecting edges to split the total graph into some sub-graphs.
-Then they trained the constituent graph models on these sub-graphs and ensemble them for the final prediction task.
-To realize graph unlearning efficiently, they proposed two graph partition algorithms and corresponding aggregation methods based on them.
-In (Cong and Mahdavi, 2022), Cong and Mahdavi filled in the gap between regularly structured data unlearning and graph data unlearning by studying the unlearning problem on the linear-GNN.
-To unlearn the knowledge of a specific node, they propose a projection-based unlearning approach PROJECTOR that projects the weight parameters of the pre-trained model onto a subspace that is irrelevant to the deleted node features.
-PROJECTOR could overcome the challenges caused by node dependency and is guaranteed to unlearn the deleted node features from the pre-trained model.
+In (Wang et al., 2023c), they implement federated unlearning based on Bayesian inference, and they propose a parameters self-sharing method to reduce the model utility degradation.
 6. Unlearning Evaluation and Verification
-Table 2. Evaluation and Verification Metrics, Datasets
-Evaluation Metrics Formula/Description Usage Literature Popular Datasets
-Accuracy Accuracy of the unlearned model on erased datasets and remaining datasets To evaluate the predictive accuracy of the unlearned model (Bourtoule et al., 2021; Guo et al., 2020), … MNIST (LeCun et al., 1998) CIFAR SVHN LSUN Photo (Shchur et al., 2018) ImageNet (Deng et al., 2009) Adult Credit info PID (Asuncion and Newman, 2007) PWD (Liu et al., 2010) AG News (Zhang et al., 2015) GTSRB (Stallkamp et al., 2011) Covtype (Blackard and Dean, 1999) HIGGS (Baldi et al., 2014) RCV1 (Glewis et al., 2004) Market-1501 (Zheng et al., 2015) YELP2018 Movielens-1m Movielens-10m Cora (Sen et al., 2008) Citseer (Yang et al., 2016) Pubmed (Fey and Lenssen, 2019) ogbn-arxiv (Hu et al., 2020) Computers (McAuley et al., 2015) CS (Shchur et al., 2018) Physics (Shchur et al., 2018)
+Table 3. Evaluation and Verification Metrics
+Evaluation Metrics Description Usage Literature
+Accuracy Model accuracy on erased datasets and remaining datasets To evaluate the predictive accuracy of the unlearned model (Bourtoule et al., 2021; Guo et al., 2020), …
 Running time The training time of unlearning process To evaluate the unlearning efficiency (Cao and Yang, 2015; Guo et al., 2020), …
 L2-norm The parameters differences between the retrained and unlearned models $||\theta_{1}-\theta_{2}||$ To evaluate the indistinguishability between two models (Wu et al., 2020)
-KL-Divergence The KLD between the distribution of the unlearned and retrained model: $\text{KL}(A(D_{r})||U(D_{e},A(D)))$ To evaluate the indistinguishability between model parameters (Nguyen et al., 2020)
-JS-Divergence The JS divergence between the predictions of retrained and unlearned model: $JS(A,U)=0.5\cdot KL(A||Q)+0.5\cdot KL(U||Q)$ To evaluate the indistinguishability between model outputs (Chundawat et al., 2022)
-Membership inference Recall (##\##detected items /#absent#/\#/ #erased items) To verify if the erased data is unlearned by the model (Graves et al., 2021)
-Forgetting rate (FR) $FR=\frac{AF-BF}{BT}$, where AF, BF and BT are defined below To measure the rate of samples that are changed from member to non -member after unlearning (Ma et al., 2022)
+KL-Divergence The KLD between the distribution of the unlearned and retrained model: $\text{KL}(\mathcal{A}(D_{r})||\mathcal{U}(D_{e},\mathcal{A}(D)))$ To evaluate the indistinguishability between model parameters (Nguyen et al., 2020)
+JS-Divergence The distance between the predictions of retrained and unlearned model: $JS(\mathcal{A},\mathcal{U})=0.5\cdot KL(\mathcal{A}||Q)+0.5\cdot KL(\mathcal{U}||Q)$ To evaluate the indistinguishability between model outputs (Chundawat et al., 2022)
+Membership inference Recall (##\##detected objects /#absent#/\#/ #erased objects) To verify if the erased sample is unlearned by the model (Graves et al., 2021)
 […]
 EMA Ensembled Membership Auditing: Ensemble multiple membership metrics and utilizes Kolmogorov-Smirno (KS) statistical tools to obtain a final auditing score To verify if a unlearned model mem -orizes a query dataset (Liu and Tsaftaris, 2020)
-MIB Membership Inference via Backdooring: Achieve membership inference for the backdoored data by querying a certain number of black-box queries to the model To verify if the model unlearns the backdoored data (Hu et al., 2022)
-Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effect is also an important problem in unlearning.
+MIB Membership Inference via Backdooring: Achieve membership inference for the backdoored data by querying a certain number of black-box queries To verify if the model unlearns the backdoored data (Hu et al., 2022)
+Forgetting rate (FR) $FR=\frac{AF-BF}{BT}$, where AF, BF and BT are defined below To measure the rate of samples that are modified from member to non-member after unlearning (Ma et al., 2022)
+$tr(I(\theta;D))$ is the trace of $I(\theta;D)$, and $I(\theta;D)$ is the Fisher Information matrix (Becker and Liebig, 2022).
+6.1. The Unlearning Verification and Auditing Methods
+Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effect is also an important problem in machine unlearning.
 In Section 3, we have introduced the L2-norm (Wu et al., 2020), KLD (Nguyen et al., 2020), and privacy leakage as the unlearning verification metrics.
-The common metrics also include the accuracy to assess the predictive accuracy of the unlearned model and running time to evaluate the unlearning efficiency.
+The common metrics also include accuracy in assessing the performance of the unlearned model and running time in evaluating the unlearning efficiency.
 We have listed recent evaluation and verification metrics in Table 3, where the popular experimental datasets are furthermore presented.
 Since accuracy, running time, L2-norm, and KLD are fixed evaluation metrics, we do not discuss them in detail here.
-In this section, we pay attention to the new evaluation metrics that recent studies tailored to unlearning.
+In this section, we introduce some new evaluation metrics that recent studies tailored to unlearning.
 Most of them are based on privacy leakage.
@@
 They used a calibration dataset and the Kolmogorov-Smirno (KS) (Demidenko, 2004) distance to confirm if an erased dataset is removed.
-Their challenge was that the method would fail when the calibration dataset was out of high quality or when the query samples were similar to the training dataset.
-An Ensembled Membership Auditing (EMA) (Huang et al., 2021a) method was proposed to overcome these limitations when auditing data erasure from a trained model.
+Their challenge was that the method would fail when the calibration dataset was out of high quality, or the query samples were similar to the training dataset.
+The EMA (Huang et al., 2021a) method was proposed to overcome these limitations when auditing data erasure from a trained model.
 EMA is a 2-step data removal auditing method that ensembles several metrics and statistical tools.
 In the first step of EMA, the auditor uses various metrics to infer if the model remembers the samples of the query dataset.
-In the second step, EMA aggregates a final auditing score by ensembling many membership metrics and statistical tools.
+EMA aggregates a final auditing score in the second step by ensembling many membership metrics and statistical tools.
 They verified whether a trained model forgets query data by the final auditing score.
@@
 (Hu et al., 2022) proposed Membership Inference via Backdooring (MIB).
-MIB leverage the property of backdoor attacks that backdoor triggers will misadvise the trained model to predict the backdoored sample to other wrong class.
-The main idea of MIB is that the user proactively adds the trigger to her data when publishing them online so that she can implement the backdoor attacks to determine if a model has been trained based on her data.
-MIB evaluate the membership inference for the triggered data by calculating the results of a certain number of black-box queries to the targetted model.
-Then, they observed the results between the targeted model and clean models to infer if the model was trained using the backdoored data.
+MIB leverages the property of backdoor attacks that backdoor triggers will misadvise the trained model to predict the backdoored sample to other wrong classes.
+The main idea of MIB is that the user proactively adds the trigger to her data when publishing them online so that she can implement the backdoor attacks to determine if the model has been trained using her dataset.
+MIB evaluates the membership inference for the triggered data by calculating the results of a certain number of black-box queries to the targetted model.
+Sequentially, they observed the results between the targeted model and clean models to infer whether the model was trained based on the backdoored samples.
 A similar membership-inference-based method was proposed in (Ma et al., 2022).
 Ma et al.
-verified the effectiveness of an unlearning method by their proposed uniform metric called forgetting rate (FR).
-Their methods were inspired by membership inference attacks and defined the evaluation metrics based on the observation of membership inference.
-Suppose the dataset $D_{e}$ is the erased dataset; the FR of an unlearning method is denoted as $FR=\frac{AF-BF}{BT}$.
-In their FR definition, $BF$ and $AF$ mean the number of samples in $D_{e}$ that are predicted to be false by a membership inference attack before and after unlearning.
-$BT$ is the number of samples in $D_{e}$ predicted to be true by a membership inference attack before machine unlearning.
-According to this definition, $FR$ presents an instinctive evaluation of how many data points are changed from member to non-member after unlearning.
+(Ma et al., 2022) verified the effect of unlearning methods by the proposed forgetting rate (FR) metric.
+The evaluation metrics were defined using the observation of membership inference.
+Suppose $D_{e}$ is the erased dataset; the FR of an unlearning method is denoted as $FR=\frac{AF-BF}{BT}$.
+In their FR definition, $BF$ and $AF$ are the samples in $D_{e}$ that are predicted as false by a membership inference attack before and after machine unlearning operations.
+$BT$ is the sample size in $D_{e}$, which is indicated to be correct by a membership inference attack before machine unlearning.
+According to the definition, $FR$ presents an instinctive evaluation of how many data points are altered from member to non-member by unlearning.
 If an unlearning method achieves that $AF>BF$ on the condition that $BT>0$, it means this method is effective.
@@
 They also proposed an unlearning method called Forsaken and evaluated Forsaken’s performance on eight standard benchmark datasets.
-The main idea of Forsaken is to continuously generate masked gradients, apply them to the model’s neurons, and model them to forget the contribution of the specific samples.
+Forsaken’s main idea is to generate masked gradients continuously, apply them to the model’s neurons, and model them to forget the contribution of the specific samples.
 To implement this process, they introduced a mask gradient generator.
-The proposed mask gradient generator could adjust the volume of the mask gradients to avoid unlearning catastrophes based on the observation of the updated model.
+The proposed mask gradient generator could adjust the volume to prevent unlearning catastrophes by observing the updated model.
 Sommer et al.
-(Sommer et al., 2022) introduced Athena, which leverages the property of backdoors, to verify the effectiveness of unlearning.
-Athena effectively certifies whether the data is deleted from an unlearning method with high confidence.
+(Sommer et al., 2022) introduced “Athena”, which leverages the property of backdooring techniques to verify the effectiveness of unlearning.
+Athena effectively and confidently certifies whether the data is deleted from an unlearning method.
 Thus, it provides a basis for quantitatively inferring unlearning.
-In their backdoor-based verification scheme, they first backdoor users’ data and then test the backdoor success probability to infer if the data is deleted.
-7. Privacy Threats on Machine Unlearning
-Although machine unlearning was first proposed to protect users’ privacy, many researchers have noticed that unlearning brings new privacy threats simultaneously.
+In their backdoor-based verification scheme, they first backdoor users’ data and then test the backdoor success probability to infer if the data is unlearned.
+Table 4. The Employed Datasets in Machine Unlearning
+Data Type Name of Datasets Feature Dimension #. Smaples Task Type Employed by
+Image MNIST (LeCun et al., 1998) $28\times 28\times 1$ $70,000$ Classification (Gupta et al., 2021; Wang et al., 2023a), …
+CIFAR10 $32\times 32\times 3$ $60,000$ Classification (Wang et al., 2023a; Tarun et al., 2023), …
+CIFAR100 $32\times 32\times 3$ $60,000$ Classification (Tarun et al., 2023; Cha et al., 2024), …
+SVHN $32\times 32\times 3$ $99,289$ Classification (Bourtoule et al., 2021; Chundawat et al., 2023), …
+ImageNet (Deng et al., 2009) $224\times 224\times 3$ $1,281,167$ Classification (Cha et al., 2024; Tarun et al., 2023), …
+GTSRB (Stallkamp et al., 2011) $32\times 32\times 3$ $51,839$ Classification (Wu et al., 2022b; Wei et al., 2023), …
+Market-1501 (Zheng et al., 2015) $128\times 64\times 3$ $32,668$ Person re-identification (Mehta et al., 2022)
+Tabular Adult $14$ $48,842$ Classification (Chen et al., 2021; Warnecke et al., 2024), …
+Credit info $30$ $284,807$ Classification and anomaly detection (Warnecke et al., 2024; Du et al., 2019), …
+Covtype $54$ $581,012$ Classification (Mahadevan and Mathioudakis, 2021; Wu et al., 2023c), …
+HIGGS (Baldi et al., 2014) $28$ $11,000,000$ Binary classification (Mahadevan and Mathioudakis, 2021, 2022), …
+YELP2018 $5$ $1,561,406$ Recommendation (Chen et al., 2022a)
+Movielens-1m $4$ $1,000,209$ Recommendation (Chen et al., 2022a)
+Movielens-10m $4$ $10,000,054$ Recommendation (Chen et al., 2022a)
+Text AG News (Zhang et al., 2015) $3$ $127,600$ Text classification (Wu et al., 2020)
+RCV1 (Glewis et al., 2004) $3$ $804,414$ Text classification (Wu et al., 2020)
+Graph Amazon Photo (Shchur et al., 2018) Features per Node: $745$ Nodes: $7,650$ Edges: $119,081$ Node classification Link prediction (Chien et al., 2022a, b), …
+Cora (Sen et al., 2008) Features per Node: $1,433$ Nodes: $2,708$ Edges: $5,429$ Node classification Link prediction (Chien et al., 2022b; Chen et al., 2022b), …
+Citseer (Yang et al., 2016) Features per Node: $3,703$ Nodes: $3,327$ Edges: $4,732$ Node classification Link prediction (Chen et al., 2022b)
+Pubmed (Fey and Lenssen, 2019) Features per Node: $500$ Nodes: $19,717$ Edges: $44,338$ Node classification Link prediction (Chen et al., 2022b; Chien et al., 2022a), …
+ogbn-arxiv (Hu et al., 2020) Features per Node: $128$ Nodes: $169,343$ Edges: $1,166,243$ Node classification Link prediction (Chien et al., 2022b, a), …
+Computers (McAuley et al., 2015) Features per Node: $767$ Nodes: $13,752$ Edges: $245,861$ Node classification Link prediction (Chen et al., 2022b; Chien et al., 2022a), …
+CS (Shchur et al., 2018) Features per Node: $767$ Nodes: $18,333$ Edges: $327,476$ Node classification Link prediction (Chen et al., 2022b)
+Physics (Shchur et al., 2018) Features per Node: Text Nodes: $27,770$ Edges: $352,807$ Node classification Link prediction (Chen et al., 2022b)
+6.2. The Employed Datasets
+We collect the commonly employed datasets in machine unlearning studies and present the detail introduction of them in Table 4.
+There are four main types of data: Image, Tabular, Text and Graph.
+Most of the unlearning studies use image datasets and train classification models based on these image datasets.
+For tabular datasets, most of them are used in recommendation systems.
+The unlearning studies that investigate how to unlearn a recommendation model will use these tabular datasets.
+Graph data is employed for node classification and link prediction tasks, which is usually used in graph unlearning studies.
+For convenience to find the related studies, we link the corresponding unlearning studies at the last column in Table 4.
+7. Privacy and Security Threats on Machine Unlearning
+Although unlearning methods were initially proposed to safeguard users’ privacy, many researchers have noticed that unlearning brings new privacy threats simultaneously.
 Similar to machine learning scenarios, there are two main kinds of privacy threats in unlearning.
 They are membership inference attacks and private information reconstruction attacks.
-In this section, we introduce the two prevalent unlearning privacy threats detailedly.
-7.1. Membership Inference Attacks in Unlearning
+In this section, we detail the two prevalent unlearning privacy threats.
+7.1. Privacy Threats on Machine Unlearning
+7.1.1. Membership Inference Attacks in Unlearning
 Chen et al.
-(Chen et al., 2021b) first pointed out that when a model is unlearned, the difference in the outputs from the model before and after unlearning leaks the privacy of the erased data.
-Then they proposed the corresponding membership inference attack pipeline in unlearning, which includes three phases: posteriors generation, feature construction, and membership inference.
+(Chen et al., 2021) first pointed out that when a model is unlearned, the discrepancy in the outputs from the model before and after unlearning leaks privacy of erased data.
+Then, they proposed the corresponding membership inference attack pipeline in unlearning, which includes three phases: posterior generation, feature construction, and membership inference.
   1. (1)
     Posteriors Generation.
-Suppose that the attacker has the ability to access two versions of the ML model, the model $\theta_{\mathcal{A}}$ before unlearning and the unlearned model $\theta_{\mathcal{U}}$.
-Assume a target sample $e$, the attacker queries $\theta_{\mathcal{A}}$ and $\theta_{\mathcal{U}}$, and obtains the corresponding posteriors, $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, which also called as confidence values in (Shokri et al., 2017).
+Suppose that the attacker can access two versions of the trained model, the model $\theta_{\mathcal{A}}$ before unlearning and the model $\theta_{\mathcal{U}}$ after unlearning.
+Assume a target data point $e$, the attacker queries $\theta_{\mathcal{A}}$ and $\theta_{\mathcal{U}}$, and has the corresponding posteriors, $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, which also called as confidence values in (Shokri et al., 2017).
   2. (2)
@@
 After achieving the two posteriors $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, the attacker sums them to make the inference feature vector $F$.
-Common methods exist to construct the feature vector shown in (Chen et al., 2021b).
+Common methods exist to construct the feature vector shown in (Chen et al., 2021).
   3. (3)
     Inference.
-After the attacker finishes training the attack model based on the created features $F$, he inputs the collected feature to the inference model to predict if the specific sample $e$ is in the erased dataset of unlearning models.
-In (Chen et al., 2021b), they assumed that the attacker has admission to two versions of ML models before and after unlearning, but it is sometimes impractical, especially in black-box learning scenarios.
+After the attacker finishes training the attack model based on the created features $F$, he inputs the collected feature to the inference model to predict if the specific sample $e$ occurs in the erased dataset of unlearning models.
+In (Chen et al., 2021), they assumed that the attacker has admission to two ML models before and after unlearning, but it is sometimes impractical, especially in black-box learning scenarios.
 Lu et al.
-(Lu et al., 2022) further proposed a label-only membership inference attack method to infer if a data point is unlearned by the model, eliminating the dependence on accessing posteriors.
+(Lu et al., 2022) further proposed a label-only membership inference method to imply if a sample is unlearned, eliminating the dependence on accessing posteriors.
 Their basic idea is that the same noise injection on candidate data points will show different results for the sample in or not in the training dataset.
-Thus, they made the adversary continuously query the original and unlearned models and add noise to change their outputing labels.
-By observing the disturbance amplitude, they can determine whether an item is deleted.
+Thus, they made the adversary continuously query the original and unlearned models and add noise to modify their outputs.
+Observing the disturbance amplitude lets them determine whether an item is deleted.
 Golatkar et al.
 (Golatkar et al., 2020b) derived an upper bound to confirm the maximizing knowledge that can be extracted from a black-box model.
-They queried the model with an image and obtained the related output.
-They then used the entropy of the result probabilities to construct an effective black-box membership inference (Shokri and Shmatikov, 2015) attack in machine unlearning.
-7.2. Privacy Reconstruction Attacks in Unlearning
-Privacy reconstruction is another popular attacking kind in machine unlearning.
+They queried the model with a picture and obtained the related output.
+They used the entropy of the result probabilities to construct an effective black-box membership inference (Shokri and Shmatikov, 2015) attack in machine unlearning.
+7.1.2. Privacy Reconstruction Attacks in Unlearning
+Privacy reconstruction is another popular attack in machine unlearning.
 In an unlearning scenario, Gao et al.
 (Gao et al., 2022) proposed the deleted reconstruction attacks to recover the removed data from the outputs of the original and unlearning models.
-In their work, they formalized erasure inference and erasure reconstruction attacks, in which the attacker aims to either infer which record is deleted or reconstruct the erased sample.
-In particular, for the deletion inference, they formalize the goal of an erasure inference to distinguish a data example $e$ was in or not in the erased dataset, $e\in D_{e}$ or $e^{\prime}\notin D_{e}$.
+In their work, they formalized erasure inference and erasure recovering attacks.
+The attacker seeks to infer which sample is removed or recovers the erased sample.
+In particular, for the deletion inference, they formulate the objective of an erasure inference to decide if a data instance $e$ was in or not in the erased dataset, $e\in D_{e}$ or $e^{\prime}\notin D_{e}$.
 For the deletion reconstruction, they focused on reconstructing the erased example $e$.
-In all their reconstruction attacks, the attacker is not given any particular examples, and its goal is to extract knowledge about the features of the erased instance.
+In all their reconstruction attacks, the attacker does not have any particular samples, and the purpose is to extract the features knowledge of the erased example.
 Specifically, the deletion reconstructions include the deleted instance reconstruction and deleted label reconstruction.
-As named, the deleted instance reconstruction is to extract all of the features of the erased instance, and the erased label reconstruction is to infer the label of the erased point in the classification problem.
+As named, the deleted instance reconstruction is to extract all of the information of the erased example, and the erased label reconstruction is to infer the label of the erased point in the classification problem.
 Zanella-Béguelin et al.
 (Zanella-Béguelin et al., 2020) indicated that the releasing snapshot of overlapped language models would leak the privacy of the training dataset.
-They verified that the model updates pose a significant risk to the private information added to or removed from the training dataset by a huge amount of experimental results.
+They verified that the model updates significantly threaten the private information added to or deleted from the training dataset by many experimental results.
 Zanella-Béguelin et al.
@@
 First, an attacker can extract particular sentences used or not in the training dataset by comparing two models.
-Second, analysing more model snapshots shows more information about the updated data than considering fewer model snapshots.
+Second, analyzing more model snapshots shows more information about the updated data than considering fewer model snapshots.
 Third, adding or deleting other non-private data during model updates can not mitigate privacy leakage.
-Fourth, differential privacy can reduce privacy leakage risks, but it also decreases the accuracy of trained models.
+Fourth, differential privacy can reduce privacy leakage risks and decrease trained models’ accuracy.
 Fifth, to mitigate the privacy leakage risks while keeping the model utility, the server can limit the model parameters access or only output a subset of the results.
-Many studies furthermore utilised these privacy threats to evaluate the unlearning effect.
+Many studies further utilized these privacy threats to evaluate the unlearning effect.
 Huang et al.
 (Huang et al., 2021a) proposed Ensembled Membership Auditing (EMA) for auditing data erasure.
-They use the membership inference to evaluate the removing effect of unlearning.
+They use the membership inference to assess the removing effectiveness of unlearning.
 Graves et al.
-(Graves et al., 2021) indicated that if an attacker can infer the sensitive information that was wanted to be erased, then it means that the server has not guarded the user’s rights to be forgotten.
+(Graves et al., 2021) indicated that if an attacker can infer the sensitive information that was wanted to be erased, then it means that the server has not guarded the rights to be forgotten.
 Baumhauer et al.
-(Baumhauer et al., 2022) developed linear filtration to sanitise classification models with logits prediction after class-wide deletion requests.
-They verified their methods by testing how well it defends against privacy attacks.
-Both (Chen et al., 2021b) and (Zanella-Béguelin et al., 2020) pointed out that differential privacy guarantees that a model does not reveal too much information about any particular training data point.
-A differentially private model can furthermore guarantee the group’s privacy by bounding the impacts of a group of training samples.
-If using DP to protect the privacy of a group of $|D\backslash D_{e}|$training samples against snapshot attacks on $\theta_{D},\theta_{D\backslash D_{e}}$, it means that $\theta_{D\backslash D_{e}}$ cannot be more useful than $\theta_{D}$.
+(Baumhauer et al., 2022) developed linear filtration to sanitize classification models with logits prediction after class-wide deletion requests.
+They verified their methods by testing how well the method defends against privacy attacks.
+Both (Chen et al., 2021) and (Zanella-Béguelin et al., 2020) pointed out that differential privacy guarantees that a model does not reveal too much knowledge about any training sample.
+A DP-protected model can further guarantee the group’s privacy by binding the impacts of a bunch of training samples.
+If using DP to protect the privacy of a bunch of $|D\backslash D_{e}|$training examples against snapshot attacks on $\theta_{D},\theta_{D\backslash D_{e}}$, it means that $\theta_{D\backslash D_{e}}$ cannot be more useful than $\theta_{D}$.
+7.2. Security Threats on Machine Unlearning
+There are several papers that address security threats targeting machine unlearning.
+For instance, Marchant et al.
+(Marchant et al., 2022) examine an attacker’s strategy to increase the computational cost of performing machine unlearning.
+Studies by Zhao et al.
+(Zhao et al., 2024) and Hu et al.
+(Hu et al., ) investigate security vulnerabilities within unlearning systems, specifically by uploading customized malicious data update requests to negatively influence the model utility.
+Additionally, some studies propose backdooring or poisoning methods that exploit unlearning requests to achieve backdoor insertion, as discussed by (Di et al., 2022) and (Liu et al., 2024).
+These works collectively highlight the need for robust security measures in machine unlearning to mitigate these emerging threats.
 8. Machine Unlearning Applications
 Table 5. Machine Unlearning Application
-Literature Application Scenarios Algorithms Realization Methods Evaluation Metric Year
-(Huang et al., 2021b) Data unlearnable EMP Add noise ACC 2021
+Literature Application Scenarios Methods Realization Methods Evaluation Metric Year
+(Huang et al., 2021b) Data “unlearnable” EMP Add noise ACC 2021
 (Liu et al., 2022b) Mitigate backdoor BAERASER Gradient ascent method ASR and ACC 2022
 (Wang et al., 2019) Mitigate backdoor - Median Absolute Deviation L1-Norm, ##\##FP, ACC, ASR 2019
-(Du et al., 2019) Anomaly detection Unlearning Maintain memory set ##\##FP and ##\##FN 2019
+(Du et al., 2019) Anomaly detection - Maintain memory set ##\##FP and ##\##FN 2019
 (Cao et al., 2018) Repair pollution KARMA Cluster and unlearn DA 2018
@@
   • DA:Detection Accuracy, EMP: Error-minimizing Perturbations
-  • ##\##FP: number of false positive, ##\##FN: number of false negative
-Besides the inherent demands in machine unlearning that draw much research attention, many researchers also find that machine unlearning can be applied in many other scenarios to solve related problems.
+  • ##\##FN: number of false negative, ##\##FP: number of false positive
+Besides the inherent demands of machine unlearning that draw much research attention, many researchers also find that machine unlearning can be applied in many other scenarios to solve related problems.
 We list the recent unlearning applications in Table 5.
-The most popular application of machine unlearning is to mitigate the anomaly, including the backdoor triggers and pollution, from a trained model.
-Wang et al.
-(Wang et al., 2019) applied machine unlearning in detecting and mitigating backdoor attacks in DNNs.
-They designed two approaches to eliminate backdoor-related neurons/weights from the infected model and repair the infected model to be robust against malicious images.
-Then, Liu et al.
-(Liu et al., 2022b) proposed that a backdoor model learned the poisoned decision boundary.
+The most popular application of machine unlearning is to mitigate the anomalies, including backdoor triggers and pollution from an already-trained model.
+Study (Wang et al., 2019) applied machine unlearning in detecting and mitigating backdoor attacks in DNNs.
+They designed two approaches to eliminate backdoored neurons from the backdoored model and repair the model to be strong against malicious pictures.
+Then, in (Liu et al., 2022b), they offered that a backdoor model learned the poisoned decision boundary.
 Data points with triggers are all classified into the target class.
 They reversed the backdoor injection process to defend against it in machine unlearning, which is simple but effective.
-Their method contains two primary steps: first, they use a max-entropy staircase approximator to complete trigger reconstruction; second, they remove the added backdoor triggers using unlearning.
+Their method contains two primary steps: first, they utilize a max-entropy staircase approximator to complete trigger reconstruction; second, they remove the added triggers using unlearning.
 They named these two key steps of BAERASER as trigger pattern recovery and trigger pattern unlearning.
@@
 Cao et al.
-(Cao et al., 2018) proposed KARMA to search different subsets of training datasets and return the subset that with the highest misclassifications.
-KARMA includes three parts.
-First, KARMA searches for potential causality that leads to the wrong classification of a ML model.
-It clusters the misclassified data points into different parts and extracts the centres of clusters.
-Then KARMA prioritizes the search for similar data samples in the training datasets based on these extracted centres.
-Second, KARMA grows the causality found in the first stage by finding more training samples and forming a cluster.
+(Cao et al., 2018) proposed “KARMA” to search various subsets of original datasets and return the subset with the highest misclassifications.
+First, KARMA searches for possible reasons that lead to the wrong ML model classification.
+It clusters the misclassified samples into various domains and extracts the middle of clusters.
+KARMA prioritizes the search for matching examples in the original datasets using these extracted centers.
+Second, KARMA grows the reason discovered in the first step by discovering more training samples and creating a cluster.
 Third, KARMA determines if a causality cluster is polluted and calculates how many samples the cluster contains.
@@
 In their presentation, they explain their techniques, emphasizing handling false negatives.
-When meeting such case $x_{t}$, their goal is, therefore, to modify the model to reduce the probability of $p(x_{t})$, or to increase the loss $\mathcal{L}(x_{t})$.
+When meeting such case $x_{t}$, their purpose is, therefore, to modify the model to reduce the possibility of $p(x_{t})$, or to increase the loss $\mathcal{L}(x_{t})$.
 Huang et al.
-(Huang et al., 2021b) presented a method that can make training examples unlearnable by injecting a type of error-minimizing noise.
-This kind of noise is intentionally generated to reduce the error of the training examples close to zero, which can mislead the model into considering there is ”nothing” to learn from these examples.
-They first tried the error-maximizing (Goodfellow et al., 2014) noise but found that this noise cannot prevent DNN learning when applied in a sample-wise manner to the training data points.
+(Huang et al., 2021b) presented a method that can make samples unlearnable by injecting error-minimizing noise.
+This noise is intentionally synthesized to diminish the error of the samples close to zero, which can mislead the model into considering there is ”nothing” to learn.
+They first tried the error-maximizing noise but found that it could not prevent DNN learning when used sample-wise to the training data points.
 Therefore, they then begin to study the opposite direction of error-maximizing noise.
@@
 For example, in exact unlearning, they designed split algorithms that divide the final model into several sub-models and avoid the stochasticity and incrementality of one sub-model to influence other sub-models.
-In approximate unlearning, they bounded the removed estimation to aviod the accuracy degradation.
+In approximate unlearning, they bounded the removed estimation to avoid accuracy degradation.
 However, it is still not easy to analyze them clearly and solve them totally.
-Moreover, as the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met new challenges.
-In this section, we will discuss the differences between various unlearning scenarios and their corresponding challenges.
+Moreover, as the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met corresponding new challenges.
+Here, we discuss the differences between various unlearning scenarios and their corresponding challenges.
 Based on these challenges, we list some potential research directions on which we can focus.
@@
 Although they proposed many methods, they just mitigated the influence of the main challenges and still have not solved them.
-Existing unlearning methods were at the beginning step that tried to implement effective unlearning.
+Existing unlearning methods were at the beginning phase, which is trying to implement effective unlearning.
 They proposed mechanisms, but most of them cannot guarantee the final results, so they also bound the unlearned item (Guo et al., 2020) or set an unlearning threshold (Nguyen et al., 2020).
-Although they also introduced many metrics to verify unlearning effect, the unlearning and verification processes are split.
-It means that the unlearning result is uncertain during the unlearning period before they finish the verification.
+Although they also introduced many metrics to verify the unlearning effect, the unlearning and verification processes are split.
+This means that the unlearning result is uncertain during the unlearning period before the verification is finished.
 Therefore, we find and list some open questions in centralized unlearning.
   1. (1)
-    Optimizing the original three challenges.
-The first two challenges are hard to solve, so recent researches focus on reducing the utility degradation after unlearning.
-Even though many studies exist, there is still a long way to eliminate the unlearning degradation.
+    Optimizing the existing challenges.
+The first two challenges are hard to solve exactly, so recent research focuses on reducing utility degradation after unlearning.
+Even though many studies exist, there is still a long way to go to eliminate the degradation of unlearning.
   2. (2)
-    How to unlearn with a certain or exact goal? In other words, can we unlearn as learning? We know the purpose, to what extent we have unlearned and when we can finish or stop unlearning process.
-Machine unlearning in distributed scenarios has many differences to which in centralized scenarios.
+    How do we unlearn with a certain or exact goal?
+In other words, can we unlearn as learning?
+If we know the purpose, to what extent we have unlearned, and when we can finish or stop the unlearning process.
+Machine unlearning in distributed scenarios has many differences from centralized scenarios.
 We take federated unlearning as a representative example of distributed unlearning.
-The first difference is that federated unlearning can only be implemented locally on the client’s side if they want to unlearn some specific samples because clients’ do not upload their data to the FL server in a federated scenario.
+The first difference is that federated unlearning can only be implemented locally on the client’s side if they want to unlearn some specific samples because clients do not upload their data to the FL server in a federated scenario.
 To avoid interacting with clients during unlearning, researchers (Liu et al., 2021; Wu et al., 2022b) proposed to unlearn the contribution of a whole client while not some samples of the client.
@@
 Therefore, federated unlearning can only be implemented using approximate unlearning methods.
-However, as we know, approximate unlearning is easy to cause catastrophic unlearning, and federated learning is more vulnerable to degradation, so how to control the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
+However, as we know, approximate unlearning easily causes catastrophic unlearning, and federated learning is more vulnerable to degradation, so controlling the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
   2. (2)
-    Privacy in federated learning is more sensitive than in other situations because FL was designed to protect participants’ privacy at the beginning.
-May unlearning bring other privacy demands in federated learning, such as secret or unnoticeable unlearning?
+    Existing federated unlearning methods require the activation of all users, including those without unlearning requests, to assist in the unlearning process.
+These approaches are impractical and inefficient, particularly when unlearning requests are frequent.
+Therefore, it is crucial to study how to balance the unlearning effect and efficiency in federated unlearning.
 Besides exploring machine unlearning based on regularly structured data, researchers tried to implement unlearning in graph data.
 Exact unlearning may be suitable for centralized graph unlearning if graph data is sparse.
-However, the challenges of approximate unlearning may be more difficult than structured-data-based unlearning because, in graph unlearning, the relationship and influence between data samples are more complex than structured data (Chen et al., 2021a; Chien et al., 2022).
+However, the challenges of approximate unlearning may be more difficult than structured-data-based unlearning because, in graph unlearning, the relationship and influence between data samples are more complex than structured data (Chen et al., 2022b; Chien et al., 2022a).
 In particular, graph data includes not only the node feature value but also the connecting edge information.
@@
 The original problems in regular data unlearning will be more challenging in graph unlearning.
-Besides these problems, graph unlearning also face unique problems that is related to edge structure information.
-In graph unlearning, how to unlearn some edges or a sub-graph is a big question.
+Besides these problems, graph unlearning also faces unique problems that are related to edge structure information.
+In graph unlearning, unlearning some edges or sub-graphs is a big question.
+After designing machine unlearning algorithms, effective verification and auditing methods are necessary (Thudi et al., 2022b).
+Most existing unlearning verification methods rely on backdooring techniques.
+However, these methods inherently degrade model utility because they require mixing backdoored samples into the model training process.
+Investigating ways to preserve model utility in backdoor-based unlearning verification methods is an under-explored area.
+Additionally, backdoor-based verification methods must include backdoored samples in the unlearning requests to verify the effectiveness of these requests.
+This requirement restricts these methods from supporting verification for single-sample unlearning requests.
+Exploring new strategies that can effectively verify unlearning requests without compromising model utility or being suitable for both single-sample and multi-sample unlearning scenarios remains a significant challenge in this field.
 Another important part is the privacy and security issues in machine unlearning.
-Although machine unlearning was proposed to protect users’ privacy when they want to leave, it brings new threats in that adversaries have a chance to infer the information about the removed data.
-Literature (Chen et al., 2021b; Zanella-Béguelin et al., 2020) have pointed out that updates of unlearning will leak privacy information, and they proposed corresponding attacks to infer this private information.
-However, most recent unlearning privacy leakage attacks are similar to those in a learning situation.
+Machine unlearning was first proposed to protect users’ privacy, but it brings new threats in that adversaries have a chance to infer the information about the removed data.
+Literature (Chen et al., 2021; Zanella-Béguelin et al., 2020) have pointed out that updates of unlearning will leak privacy information, and they proposed corresponding attacks to infer this private information.
+However, the most recent unlearning privacy leakage attacks have been similar to those in a learning situation.
 An attack that is tailored to unlearning mechanisms is expected.
@@
 Although researchers proposed to unlearn a backdoor trigger (Liu et al., 2022b) or pollution (Cao et al., 2018), they only used a few unlearning techniques and paid more effort to detect those anomalies.
-One important reason is that unlearning mechanism is not mature enough.
-We are at the beginning of machine unlearning, and there are still many problems in unlearning itself that need to be solved.
+One important reason is that the unlearning mechanism is not mature enough.
+We are at the beginning of machine unlearning investigation, and there are still many under-explored problems in unlearning itself.
 Finding machine unlearning applying situation and tailoring unlearning techniques to this situation is the direction of unlearning application.
 10. Summary
-The survey aims to offer a comprehensive and up-to-date overview of machine unlearning techniques in a systematic way.
+The survey aims to offer a comprehensive and systematic overview of machine unlearning techniques.
 We organize the main challenges, research advancements, corresponding techniques, and privacy and security issues in machine unlearning.
@@
 We first illustrate the complete unlearning framework, including the learning, request, unlearning, and verification.
-Then, we catalogue recent studies briefly into exact unlearning and approximate unlearning, and we introduce the technical details of these unlearning methods.
+Then, we briefly categorize recent studies into exact and approximate unlearning and introduce the technical details of corresponding unlearning methods.
 Moreover, we noticed some new unlearning scenarios, such as federated unlearning and graph unlearning, which are also introduced in the survey.
-Besides, we consider privacy and security issues in machine unlearning as an important consisting part.
+Besides, we consider privacy and security issues in machine unlearning to be an important part of the studies.
 We collect and summarize the related literature about unlearning privacy threats and applications.
-Ultimately, the survey provides clear summaries and comparisons between various unlearning scenarios and corresponding methods, giving a comprehensive view of current work and listing the challenges and open problems of different scenarios of machine unlearning.
+Ultimately, the survey provides clear summaries and comparisons between various unlearning scenarios and corresponding methods, giving a comprehensive picture of existing work and listing the challenges and open problems of different scenarios of machine unlearning.
 We hope our survey can help classify future unlearning studies, achieve a more in-depth understanding of unlearning methods, and address complex challenges.
 We believe the open problems listed in Section 9 will still be challenging in the following years, and we will try to optimize some of them.
-Last but not least, we expect this survey can help researchers in the study of machine unlearning, regardless of the unlearning strategies or unlearning privacy and security threats or applications of unlearning.
+Last but not least, we expect the survey can help researchers in the study of machine unlearning, regardless of the unlearning strategies or unlearning privacy and security threats or applications of unlearning.

### v3

@@
 This survey aims to systematically classify a wide range of machine unlearning studies, discussing their differences, connections, and open problems.
-We categorize current unlearning methods into four key areas: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in unlearning.
-Given that centralized unlearning is the primary focus, we further divide its discussion into two parts: first, at a higher level, we classify centralized unlearning into exact unlearning and approximate unlearning; second, we provide a detailed introduction to the techniques used in these methods.
-Next, we introduce federated unlearning, another emerging area that has garnered significant attention.
-Following the discussion on unlearning methods, we review studies on unlearning verification and audit, which assess the effectiveness of existing unlearning algorithms.
-Additionally, we consider the privacy and security issues essential in machine unlearning and compile the related literature.
+We categorize current unlearning studies into four key areas: traditional unlearning methods, unlearning verification, domain-centric machine unlearning, and privacy and security issues in unlearning.
+In the traditional unlearning part, we first classify traditional unlearning into exact unlearning and approximate unlearning at a higher level; then, we provide detailed introductions to the techniques used in these studies.
+Next, we review studies on unlearning verification to introduce how to assess the effectiveness of unlearning operations.
+Following the discussion of traditional unlearning and verification methods, we introduce domain-centric machine unlearning, including graph unlearning, federated unlearning, diffusion model unlearning, and large language model unlearning.
+Additionally, we consider privacy and security issues essential in machine unlearning and compile related literature.
 Finally, we discuss the challenges of various unlearning scenarios and highlight potential research directions.
-Federated Unlearning; Machine Unlearning; Unlearning Verification
-ccs: Computing methodolgies Machine learningccs: Computing methodolgies Aritificial intelligenceccs: Security and privacyccs: Theory of computation
+Machine Unlearning; Unlearning Verification; Federated Unlearning; Graph Unlearning; Diffusion Model Unlearning; Large Language Model Unlearning
+$c$s: Computing methodologies Machine learningccs: Computing methodologies Aritificial intelligenceccs: Security and privacyccs: Theory of computation
 1. Introduction
@@
 The deletion is not only erasing their data from a database, but it also needs to delete the influence of the specified samples from trained models.
-The process of data removal from models was first conceptualized as machine unlearning (Cao and Yang, 2015).
+The process of data removal from models was conceptualized as machine unlearning (Cao and Yang, 2015).
 Specifically, suppose a user (Alice) wants to exercise her right (Mantelero, 2013) when quitting a ML application, then the trained model of such application must ”unlearn” her data.
@@
 (Nguyen et al., 2020) indicated that an unlearned model typically has worse model utility than the model retrained from scratch.
-The degradation would be exponential, especially when a method tries to delete a huge amount of data samples.
+The degradation would be severe, especially when a method tries to delete a huge amount of data samples.
 They referred to such sharp degradation as catastrophic unlearning (Nguyen et al., 2020).
@@
 Recently, many studies have put efforts into solving these three main challenges and proposed many novel mechanisms that promote the progress of machine unlearning.
-[…]
 ARC
-Machine UnlearningCentralized UnlearningUnlearning CategoriesExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a; Schelter et al., 2023; Koch and Soll, 2023; Hu et al., 2024a)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021; Tarun et al., 2023; Wang et al., 2023a; Chundawat et al., 2023; Warnecke et al., 2024; Liu et al., 2023a; Lin et al., 2023; Cha et al., 2024) Unlearning TechniquesSplit Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a)Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Graph Unlearning: (Chien et al., 2022a; Chen et al., 2022b; Li et al., 2024; Cong and Mahdavi, 2022; Wang et al., 2023b; Wu et al., 2023a; Zhang et al., 2024; Wu et al., 2023b)Federated (Distributed) Unlearning: (Wang et al., 2022a, 2024; Fraboni et al., 2024; Wang et al., 2023c; Liu et al., 2021, 2022a; Zhang et al., 2023a; Gupta et al., 2021; Wu et al., 2022b)Unlearning Verification: (Hu et al., 2022; Guo et al., 2023; Thudi et al., 2022a; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a; Thudi et al., 2022b) Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b; Zhang et al., 2023b)Security Threats in Unlearning : (Marchant et al., 2022; Zhao et al., 2024; Di et al., 2022; Hu et al., ; Liu et al., 2024)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Wei et al., 2023; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
+Machine UnlearningTraditional Unlearning MethodsExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a; Schelter et al., 2023; Koch and Soll, 2023; Hu et al., 2024a)Split Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021; Tarun et al., 2023; Wang et al., 2023a; Chundawat et al., 2023; Warnecke et al., 2024; Liu et al., 2023a; Lin et al., 2023; Cha et al., 2024) Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Unlearning Verification: (Hu et al., 2022; Guo et al., 2023; Thudi et al., 2022a; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a; Thudi et al., 2022b) Domain-Centric Machine UnlearningGraph Unlearning: (Chien et al., 2022a; Chen et al., 2022b; Li et al., 2024a; Cong and Mahdavi, 2022; Wang et al., 2023b; Wu et al., 2023a; Zhang et al., 2024a; Wu et al., 2023b)Federated (Distributed) Unlearning Methods: (Wang et al., 2022a, 2024a; Fraboni et al., 2024; Wang et al., 2023c; Liu et al., 2021, 2022a; Zhang et al., 2023a; Gupta et al., 2021; Wu et al., 2022b)Diffusion Model Unlearning: (Alberti et al., 2025; Chen et al., 2024; Wu et al., 2025a; Li et al., 2024b; Zhang et al., 2024b; Wu et al., 2025b; Cywiński and Deja, 2025; Zhang et al., 2024c; Ko et al., 2024; George et al., 2025)Large Language Models Unlearning: (Yao et al., 2024a, b; Liu et al., 2025a, 2024a; Huang et al., 2024a; Wang et al., 2025a; Liu et al., 2024b; Jin et al., 2024; Li et al., 2024c)Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b; Zhang et al., 2023b)Security Threats in Unlearning : (Marchant et al., 2022; Zhao et al., 2024; Di et al., 2022; Hu et al., ; Liu et al., 2024c)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Wei et al., 2023; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
 Figure 1.
@@
 The introduction order will also follow this figure.
-We classify the current unlearning literature into four main scenarios: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in machine unlearning.
-This survey aims to classify and systematize machine unlearning methods based on the research problems and objectives in the unlearning process and to review their differences and connections, as well as their advantages and disadvantages.
-We first categorize unlearning approaches into four main areas: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in machine unlearning, as shown in Fig.
+We classify the current unlearning literature into four main scenarios: traditional unlearning methods, unlearning verification, domain-centric machine unlearning, and privacy and security issues in machine unlearning.
+This survey aims to classify and systematize machine unlearning methods based on the research problems and objectives in the unlearning process and to review their differences, connections, as well as their advantages and disadvantages.
+Our survey includes four main categories: traditional unlearning methods, unlearning verification, domain-centric unlearning methods, and privacy and security issues in machine unlearning, as shown in Fig.
 1.
-Since many studies focus on centralized unlearning, we dedicate two sections to introducing corresponding studies in this area.
-We classify these unlearning methods into: Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022a; Nguyen et al., 2020), where exact unlearning extends the idea of naive unlearning to design efficient retraining scheme with less computation cost.
-Approximate unlearning tries to unlearn a model approximate to the model retrained using the remaining dataset.
-We also provide a detailed introduction to the popular techniques used to implement these two types of unlearning.
-In addition to centralized unlearning methods, we introduce federated (distributed) unlearning studies (Wang et al., 2022a; Liu et al., 2021), which are gaining increasing importance.
-After discussing unlearning implementation methods, verifying the unlearning effect is equally crucial and has garnered significant research attention (Thudi et al., 2022b).
+In the traditional unlearning category, we will introduce the representative Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022a; Nguyen et al., 2020) methods, including basic concepts and detailed technical implementation.
+After discussing traditional unlearning methods, verifying the unlearning effectiveness is equally crucial and has garnered significant research attention (Thudi et al., 2022b).
 Therefore, we subsequently review related research on unlearning verification.
+Next, we will introduce domain-centric unlearning methods, including graph unlearning (Chien et al., 2022a; Chen et al., 2022b), federated unlearning (Wang et al., 2022a; Liu et al., 2021), diffusion model unlearning (Alberti et al., 2025; Chen et al., 2024), and large language model (LLM) unlearning (Yao et al., 2024a, b).
 Finally, we consider privacy and security issues in machine unlearning essential, organizing and reviewing the related publications on privacy and security threats, defenses, and unlearning applications.
+Table 1. Comparison of this survey with representative existing surveys on machine unlearning.
+Survey Year Main focus Difference from this survey
+Mahadevan and Mathioudakis, Certifiable Machine Unlearning for Linear Models (Mahadevan and Mathioudakis, 2021) 2021 Experimental review of approximate/certifiable unlearning methods for linear models, with emphasis on efficiency, effectiveness, and certifiability trade-offs.
+We provide a broader taxonomy spanning traditional unlearning, verification, domain-centric unlearning, such as federated unlearning and LLM unlearning, and privacy/security issues.
+Nguyen et al., A Survey of Machine Unlearning (Nguyen et al., 2022b) 2022 Broad overview of formulations, design criteria, removal requests, algorithms, scenarios, and applications.
+Offers a strong general foundation, but places more emphasis on concepts and formulations than on systematically comparing scenario-specific challenges.
+Wang et al., Federated Unlearning and Its Privacy Threats (Wang et al., 2023d) 2023 Federated unlearning, especially unlearner roles, privacy leakage, membership inference risks, and defenses.
+Focused on federated settings only, particularly privacy threats in federated unlearning, rather than the full machine unlearning landscape.
+Liu et al., A Survey on Federated Unlearning: Challenges, Methods, and Future Directions (Liu et al., 2023b) 2024 Dedicated taxonomy of federated unlearning methods, workflows, challenges, applications, and future directions.
+Comprehensive for federated unlearning, but not designed to compare federated unlearning with centralized unlearning, verification, and broader privacy/security issues within one unified framework.
+Liu et al., Threats, Attacks, and Defenses in Machine Unlearning: A Survey (Liu et al., 2025b) 2025 Security-oriented review of vulnerabilities, attacks, and defenses in machine unlearning systems.
+Centers on threat taxonomy and defensive mechanisms, rather than on the full algorithmic and scenario-based landscape of machine unlearning.
+This survey Unified, SLR-based survey covering traditional centralized unlearning, unlearning verification, domain-centric unlearning, privacy/security issues, applications, open questions, and their interconnections.
+Distinctive in integrating previously separate strands into one taxonomy and explicitly comparing their differences, connections, advantages, limitations, and research gaps.
 There are only a few surveys on machine unlearning because it is a relatively new research domain, and most of them are preprints on Arxiv, focusing on certain unlearning scenarios.
@@
 (Nguyen et al., 2022b) summarized the general unlearning framework and added the unlearning verification part to it, they focused primarily on introducing problem formulations and technical definitions.
-However, none of these surveys reviewed the privacy and security issues in machine unlearning, nor did they comprehensively discuss the differences and connections between various unlearning scenarios.
+For an security-oriented review of vulnerabilities in machine unlearning systems, see (Liu et al., 2025b).
+We provide a comparison difference between existing surveys in Table 1.
 Compared with existing surveys on machine unlearning, we performed a systematic literature review (SLR) following the guidelines presented by (Keele et al., 2007) and as exemplified in (Kubrak et al., 2022).
-The contributions of this paper include:
-  • We systematically catalog machine unlearning studies according to their scenarios, which can be briefly classified into centralized unlearning, distributed unlearning, unlearning verification, and privacy and security issues in machine unlearning.
-  • For the centralized unlearning scenario, which draws the most attention, we divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
+The main contributions are as follows.
+  • We systematically catalog machine unlearning studies into traditional unlearning methods, unlearning verification, domain-centric unlearning, and privacy and security issues in machine unlearning.
+  • For the traditional centralized unlearning scenario, which draws the most attention, we divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
   • We further explore the privacy and security threats that target machine unlearning and discuss the applications of unlearning in defending against traditional security and privacy issues.
-  • We discuss different challenges between various machine unlearning areas, from centralized unlearning to federated unlearning, unlearning verification, and privacy and security issues in unlearning.
+  • We discuss different challenges in various machine unlearning areas, from traditional centralized unlearning to federated unlearning and generative model unlearning, unlearning verification, and privacy and security issues in unlearning.
 We list the related open questions of each scenario and present potential research directions to solve them.
@@
 1.
-The primary content of centralized unlearning is presented in Section 4, which includes an introduction from two main unlearning categories and corresponding in-depth techniques.
-We collect and introduce the federated unlearning scenarios in Section 5.
+The primary content of traditional unlearning is presented in Section 4, which includes two main unlearning categories and corresponding in-depth techniques.
 Section 5 introduces the unlearning evaluation and verification methods.
+We collect and introduce the unlearning methods for specific domains such as federated unlearning, graph unlearning, diffusion model unlearning, and LLMs unlearning, in Section 6.1.
 The privacy and security issues in machine unlearning are divided into two sections.
@@
 Section 8 discusses machine unlearning applications, which are mainly applied in dealing with security issues.
-In Section 9, we discuss the challenges of current unlearning methods in a general way.
-After that, we enumerate the unaddressed issues for further research.
+In Section 9, we discuss the challenges of current unlearning methods in a general way and enumerate the potential directions for further research.
 At last, in Section 10, we provide a summary of the survey.
@@
 2, and we summarize the notations in Table 2.
-Let $\mathcal{Z}$ denote as a space of data items; the particular (full) training dataset is $D\in\mathcal{Z}$.
+Let $\mathcal{Z}$ denote a space of data items; the particular (full) training dataset is $D\in\mathcal{Z}$.
 A learning process can be demonstrated as Step 0 in Fig.
@@
 The requested unlearning data $D_{e}$ can be data samples (Bourtoule et al., 2021), classes (Warnecke et al., 2023), or graph nodes (Chen et al., 2022b).
-Then, in Step 2, the server removes the contribution of $D_{e}$ using a machine unlearning algorithm $\mathcal{U}$.
+Then, in Step 2, the server removes the contribution of $D_{e}$ using a designed machine unlearning algorithm $\mathcal{U}$.
 The unlearned model can be described as $M_{D\setminus D_{e}}=\mathcal{U}(M,D,D_{e})$.
-When dealing with an unlearning request, we hope that the unlearned model $\mathcal{U}(M,D,D_{e})$ is the same as the retrained model $\mathcal{A}(D\setminus D_{e})$ (i.e., $\mathcal{U}(M,D,D_{e})\simeq\mathcal{A}(D\setminus D_{e})$).
+The standard aim of unlearning is to ensure the unlearned model $\mathcal{U}(M,D,D_{e})$ is the same as the retrained model $\mathcal{A}(D\setminus D_{e})$ (i.e., $\mathcal{U}(M,D,D_{e})\simeq\mathcal{A}(D\setminus D_{e})$).
 Most unlearning studies ended at this step.
-However, we do need an effective evaluation metric to assess the unlearning effect, as in Step 3 in Fig.
+However, we do need an effective evaluation metric to assess the unlearning effectiveness, as in Step 3 in Fig.
 2.
 Therefore, we add the verification step as the last step in the unlearning process, and we will introduce the related literature later.
-Figure 2. Machine Unlearning Process
 Table 2. Notations in Machine Unlearning
@@
 $D_{\text{probe}}$ the probing data used in attacking $M_{D\backslash D_{e}}$ the unlearned model
+Figure 2. Overview of the machine unlearning pipeline, from model training and unlearning request types to unlearning execution and verification methods.
 3.2. Machine Unlearning Evaluation Metrics
@@
 In Bayes-based or Markov chain Monte Carlo-based unlearning methods (Nguyen et al., 2020), researchers utilize KLD (Joyce, 2011) to optimize approximate models, employing it to measure the distance between two probability distributions.
-Recent unlearning studies have also used KLD to estimate the unlearning effect by comparing the distributions of retrained and unlearned models (Nguyen et al., 2020).
+Recent unlearning studies have also used KLD to estimate the unlearning effectiveness by comparing the distributions of retrained and unlearned models (Nguyen et al., 2020).
 Evaluation Metric based on Privacy Leakage.
@@
 Conversely, if the backdoor trigger cannot attack the unlearned model, it proves that the unlearning algorithm is effective.
-Similar methods were also used in (Thudi et al., 2022a; Ma et al., 2022) to evaluate the unlearning effect.
-3.3. Majority of Tools Used in Unlearning
+Similar methods were also used in (Thudi et al., 2022a; Ma et al., 2022) to evaluate the unlearning effectiveness.
+3.3. Tools Used in Unlearning
 Differential Privacy (DP).
@@
 3.
-4. Centralized Machine Unlearning
-In this section, we classify existing centralized unlearning methods by their inherent mechanism and designed purposes and present the corresponding detailed techniques.
+4. Traditional Centralized Machine Unlearning
+In this section, we classify existing traditional centralized unlearning methods by their inherent mechanism and designed purposes and present the corresponding detailed techniques.
 4.1. Unlearning Solution Categories
@@
 However, it is inefficient because it requires storing the entire original dataset and retraining the model from scratch, which consumes significant storage and computational resources, especially in deep learning scenarios.
-Therefore, researchers tried to design effective and efficient unlearning mechanisms.
-Exact unlearning was proposed to reduce the computation cost by splitting training sub-models based on pre-divided data sub-sets (Bourtoule et al., 2021).
-It also divides the stochasticity and the incrementality into sub-models.
-It can unlearn an exact model by ensembling the consisting submodels, but it still needs to store all the split data subsets for fast retraining.
-Another approach is approximate unlearning (Sekhari et al., 2021).
-It can reduce both storage and computation costs because it unlearns only based on the erased data.
-However, it is difficult for approximate unlearning methods to control the accuracy degradation due to the challenges in estimating stochasticity and incrementality during the training process.
-Most of them bounded their estimation to avoid dramatic accuracy reduction.
+Therefore, researchers tried to design effective and efficient unlearning mechanisms, and two representative solutions are exact unlearning and approximate unlearning.
 4.1.1. Exact Unlearning
-Exact unlearning could also be called fast retraining, whose basic idea is derived from naive retraining from scratch.
-Following the background introduction of unlearning, we know the learning and unlearning algorithm, $\mathcal{A}(D)$ and $\mathcal{U}$, based on the trainset $D$ and erased dataset $D_{e}\subseteq D$, respectively.
+Exact unlearning is also called fast retraining, whose basic idea is derived from naive retraining from scratch.
+Following the background of unlearning, we know the learning and unlearning algorithm, $\mathcal{A}(D)$ and $\mathcal{U}$, based on the trainset $D$ and erased dataset $D_{e}\subseteq D$, respectively.
 If $\mathcal{U}(\cdot)$ is implemented as naive retraining, the equality between $\mathcal{A}(D\setminus D_{e})\in\mathcal{H}$ and $\mathcal{U}(M,D,D_{e})\in\mathcal{H}$ is absolutely guaranteed.
@@
 They then ensemble the retrained sub-model and other sub-models as the unlearned model.
-As we introduced above, exact unlearning aims to mitigate the computation cost when retraining a new model by transforming the original learning algorithms into an ensembling form.
+Exact unlearning aims to mitigate the computation cost when retraining a new model by transforming the original learning algorithms into an ensembling form.
 It divides the stochasticity and incrementality into several sub-models to reduce their influence.
@@
 The popular solutions are certified-removal (Guo et al., 2020) and Bayes-based mechanisms (Nguyen et al., 2020), which are introduced in technical detail in Section 4.2.
-Here, we briefly summarize current studies about approximate unlearning.
-A certified-removal mechanism (Guo et al., 2020) was an approximate unlearning method similar to differential privacy.
-Guo et al.
-(Guo et al., 2020) defined the $\epsilon-$approximate unlearning, which makes sure that the model after and before unlearning must be $\epsilon$-indistinguishable as the definition in DP.
-The difference between $\epsilon-$approximate unlearning and $\epsilon$-DP is that the mechanism $\mathcal{A}$ on differential privacy is needed never to memorize the data in the first place, which is impossible in machine unlearning.
-The machine learning model does not learn anything from the training dataset if $\mathcal{A}$ is differentially private (Bourtoule et al., 2021).
-A similar solution to (Guo et al., 2020) is introduced in (Sekhari et al., 2021), which erases the influence of the specified samples from the gradients during unlearning.
-To mitigate accuracy degradation, model parameter updates are constrained by a differential unlearning definition.
-Different from indistinguishable unlearning, other methods used the Bayes theorem to unlearn an approximate posterior.
-In (Nguyen et al., 2020), Nguyen et al.
-employed Bayesian inference to unlearn an approximate model using the erased data.
-Moreover, Fu et al.
-(Fu et al., 2022) developed a similar unlearning method, which uses Markov chain Monte Carlo (MCMC) methods.
-In (Nguyen et al., 2022a), the authors also explained the effectiveness of the approximate unlearning method from the perspective of MCMC.
-Nevertheless, since those techniques are approximately unlearning the contribution of all the erasing data, including the inputs and labels, they inevitably decrease the model accuracy to some extent after unlearning.
+Although those techniques are approximately unlearning the contribution of all the erasing data, including the inputs and labels, they inevitably decrease the model accuracy to some extent after unlearning.
 Main Challenges of Approximate Unlearning.
-As we discussed at the beginning, there are three main challenges in machine unlearning.
-In centralized scenarios, these studies mainly aim at solving the basic machine unlearning problem, which unavoidably faces three challenges: stochasticity of training, incrementality of training, and catastrophe of unlearning.
+In centralized scenarios, researchers aiming at solving the basic machine unlearning problem will unavoidably face three challenges: stochasticity of training, incrementality of training, and catastrophe of unlearning.
 The exact unlearning methods extend the retraining idea, which avoids facing these challenges but consumes lots of storage costs.
@@
   • To overcome the stochasticity and incrementality challenges when estimating the unlearning influence, one popular strategy is based on the first-order and second-order influence function (Basu et al., 2020), which is calculated based on the perturbation theory (Avrachenkov et al., 2013).
+At the same time, since classical influence functions are best justified under smooth and strongly convex objectives, whereas deep neural networks usually involve highly non-convex loss landscapes, the local Taylor approximation and inverse-Hessian estimation can become unstable (Koh and Liang, 2017; Basu et al., 2020).
   • The unlearning catastrophe appears commonly in approximate unlearning, and many studies try to propose some methods to solve this problem.
@@
 solves this problem by adding a model utility compensation task during unlearning optimization and finding the optimal balance based on multi-objective training methods.
-4.2. Detailed Techniques of Centralized Unlearning
-Table 2. Popular Unlearning Techniques
+4.2. Detailed Techniques of Traditional Centralized Unlearning
+Table 3. Traditional Unlearning Techniques
 Unlearning Literature Taxonomy Requests Type Techniques Realization Method Year
@@
 Gif (Wu et al., 2023b) Approximate unlearning Graph nodes Certified Data Removal Influence function 2023
-SUMMIT (Zhang et al., 2024) Approximate unlearning Graph nodes Certified Data Removal Multi-Objective Optimization 2024
+SUMMIT (Zhang et al., 2024a) Approximate unlearning Graph nodes Certified Data Removal Multi-Objective Optimization 2024
 EUBO, rKL (Nguyen et al., 2020) Approximate unlearning Samples Bayesian Unlearning VBI 2020
@@
 Two primary techniques used in approximate unlearning are certified data removal and Bayesian-based unlearning.
-The introduction of these techniques will primarily follow the sequence of the unlearning techniques presented in Fig.
-1.
 4.2.1. Split Unlearning
@@
 4.2.2. Certified Data Removal
+Certified data removal unlearning methods usually define their unlearning algorithms as $\epsilon$-indistinguishable unlearning, which is similar to the differential privacy definition (Dwork et al., 2014).
+An example is presented in Figure 6.
+Most of them use the Hessian matrix (Bishop, 1992) to evaluate the contribution of erased data samples for unlearning subtraction.
+After estimating the impact of the erased data samples, they unlearn by subtracting these impacts with an updating bound from the unlearning model.
 Figure 6.
@@
 The unlearned model is expected to approach the retrained model.
-Certified data removal unlearning methods usually define their unlearning algorithms as $\epsilon$-indistinguishable unlearning, which is similar to the differential privacy definition (Dwork et al., 2014).
-An example is presented in Figure 6.
-Most of them use the Hessian matrix (Bishop, 1992) to evaluate the contribution of erased data samples for unlearning subtraction.
-After estimating the impact of the erased data samples, they unlearn by subtracting these impacts with an updating bound from the unlearning model.
 In (Guo et al., 2020), Guo et al.
@@
 the test loss, where $n$ is the size of the original trainset, and $d$ is the dimension of trainset inputs.
-The aforementioned methods are trying to address the basic unlearning problem from a certified data removal perspective.
-Neel et al.
-(Neel et al., 2021) extended the definition of unlearning to include updates, which encompass both ”add” ($D\cup\{z\}$) and ”delete” ($D\backslash\{z\}$), where $z=(x,y)$ and $z\in\mathcal{Z}$ is a data point.
-They follow the definition in (Guo et al., 2020) and define similar $(\epsilon,\delta)$-indistinguishability.
-Furthermore, they extend $(\epsilon,\delta)$-indistinguishability to both ”add” and ”delete” updates, which can also be denoted as $(\epsilon,\delta)$-publishing.
+[…]
 4.2.3. Bayesian-based Unlearning
@@
 As they indicated, though their algorithm cannot wholly remove the learned knowledge from the already trained distribution, their method can still help the unlearned model approach its local minimum.
-Figure 7.
-(a) Regular data unlearning, which only unlearns the data sample.
-(b) Graph unlearning.
-Since the graph data contains both node features and connective edge information, we may need to unlearn edge information or both edges and nodes in graph unlearning.
-4.2.4. Graph Unlearning
-We introduce graph unlearning as a representative kind of irregular data unlearning.
-In (Chien et al., 2022a; Chen et al., 2022b; Cong and Mahdavi, 2022; Zhang et al., 2024), researchers extend regular data machine unlearning to a graph data scenario.
-Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig.
-7.
-Therefore, Chien et al.
-(Chien et al., 2022a) proposed node unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC).
-Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees.
-They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
-Chen et al.
-(Chen et al., 2022b) found that applying SISA (Bourtoule et al., 2021) unlearning methods to graph data learning will severely harm the graph-structured information, resulting in model utility degradation.
-Therefore, they proposed a method called GraphEraser to implement unlearning tailored to graph data.
-Similar to SISA, they first cut off some connecting edges to split the total graph into some sub-graphs.
-Then, they trained the constituent graph models on these sub-graphs and ensembled them for the final prediction task.
-To realize graph unlearning efficiently, they proposed two graph partition algorithms and corresponding aggregation methods based on them.
-In (Cong and Mahdavi, 2022), Cong and Mahdavi filled in the gap between regularly structured data unlearning and graph data unlearning by studying the unlearning problem on the linear-GNN.
-To remove the knowledge of a specified node, they design a projection-based unlearning approach, PROJECTOR, that projects the weights of the pre-trained model onto a subspace irrelevant to the deleted node features.
-PROJECTOR could overcome the challenges caused by node dependency and is guaranteed to unlearn the deleted node features from the pre-trained model.
-5. Federated Unlearning
-We have introduced popular centralized machine unlearning in the former section.
-Their main purpose is to reduce the calculation and storage costs and the accuracy degradation caused by unlearning.
-Recently, we noticed that some researchers paid attention to realizing unlearning in other scenarios, such as in federated learning (FL) scenarios.
-In this section, we will introduce the literature about federated unlearning.
-FL was initially introduced to protect the privacy of participating clients during the machine learning training process in distributed settings.
-All participants will only upload their locally trained model parameters instead of their sensitive local data to the FL server during model training processes.
-Therefore, in a federated learning scenario, limited access to the dataset will become a unique challenge when implementing unlearning.
-Since the local data cannot be uploaded to the federated learning (FL) server side, most federated unlearning methods try to erase a certain client’s contribution from the trained model by storing and estimating the contribution of uploaded parameters.
-In this situation, they can implement federated unlearning without interacting with the client, shown as the server-side federated unlearning in Fig.
-8 (a).
-The two representative methods are (Liu et al., 2021; Wu et al., 2022b).
-Liu et al.
-(Liu et al., 2021) proposed “FedEraser” to sanitize the impact of a FL client on the global FL model.
-In particular, during FL training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates.
-Based on these operations, they reconstructed the unlearned FL model instead of retraining a new model from scratch.
-However, FedEraser can only unlearn one client’s data, which means it must unlearn all the contributions of this specific client’s data.
-It is unsuitable for a client who wants to unlearn a small piece of his data.
-Study (Wu et al., 2022b) tried to erase a client’s influence from the FL model by removing the historical updates from the global model.
-They implemented federated unlearning by using knowledge distillation to restore the contribution of clients’ models, which does not need to rely on clients’ participation and any data restriction.
-Figure 8. Comparison between (a) server-side federated unlearning and (b) client-side federated unlearning
-Wang et al.
-(Wang et al., 2022a) explored the problem of how to selectively unlearn a category from a trained CNN classification model in FL.
-The biggest challenge is that the local data utilized for the FL global model training cannot be accessed globally.
-Therefore, they explored the inner influence of each channel and observed that various channels have distinct impacts on different categories.
-They proposed a method that does not require accessing the data used for training and retraining from scratch to cleanly scrub the information of particular categories from the global FL model.
-A method called TF-IDF was introduced to quantize the class discrimination of channels.
-Then, they pruned the channels with high TF-IDF scores on the erased classifications to unlearn them.
-Evaluations of CIFAR10 and CIFAR100 demonstrate the unlearning effect of their method.
-Different from unlearning a whole client’s influence and unlearning a class, Liu et al.
-(Liu et al., 2022a; Wang et al., 2023c) investigated how to unlearn data samples in FL, shown as the client-side federated unlearning in Fig.
-8 (b).
-In (Liu et al., 2022a), they first defined a federated unlearning problem and proposed a fast retraining method to withdraw the influence of data from the FL model.
-Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods and the first-order Taylor approximate method (Liu et al., 2022a).
-They utilized the practical Fisher Information Matrix to model the Hessian matrix at a low cost.
-When unlearning, they subtract the estimation matrix from gradients with a bound.
-In (Wang et al., 2023c), they implement federated unlearning based on Bayesian inference, and they propose a parameters self-sharing method to reduce the model utility degradation.
 5. Unlearning Evaluation and Verification
+Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effectiveness is also an important problem in machine unlearning.
+In this section, we will introduce some common and basic unlearning verification methods and evaluation datasets.
 Table 4. Evaluation and Verification Metrics
@@
 Running time The training time of unlearning process To evaluate the unlearning efficiency (Cao and Yang, 2015; Guo et al., 2020), …
-L2-norm The parameters differences between the retrained and unlearned models $||\theta_{1}-\theta_{2}||$ To evaluate the indistinguishability between two models (Wu et al., 2020)
+L2-norm The parameters differences between the retrained and unlearned models $\theta_{1}-\theta_{2}||$ To evaluate the indistinguishability between two models (Wu et al., 2020)
 KL-Divergence The KLD between the distribution of the unlearned and retrained model: $\text{KL}(\mathcal{A}(D_{r})||\mathcal{U}(D_{e},\mathcal{A}(D)))$ To evaluate the indistinguishability between model parameters (Nguyen et al., 2020)
 JS-Divergence The distance between the predictions of retrained and unlearned model: $JS(\mathcal{A},\mathcal{U})=0.5\cdot KL(\mathcal{A}||Q)+0.5\cdot KL(\mathcal{U}||Q)$ To evaluate the indistinguishability between model outputs (Chundawat et al., 2022)
-Membership inference Recall (##\##detected objects /#absent#/\#/ #erased objects) To verify if the erased sample is unlearned by the model (Graves et al., 2021)
+Membership inference Recall (#\#detected objects /#/\#erased objects) To verify if the erased sample is unlearned by the model (Graves et al., 2021)
 […]
-EMA Ensembled Membership Auditing: Ensemble multiple membership metrics and utilizes Kolmogorov-Smirno (KS) statistical tools to obtain a final auditing score To verify if a unlearned model mem -orizes a query dataset (Liu and Tsaftaris, 2020)
+EMA Ensembled Membership Auditing: Ensemble multiple membership metrics and utilizes Kolmogorov-Smirnov (KS) statistical tools to obtain a final auditing score To verify if an unlearned model memorizes a query dataset (Liu and Tsaftaris, 2020)
 MIB Membership Inference via Backdooring: Achieve membership inference for the backdoored data by querying a certain number of black-box queries To verify if the model unlearns the backdoored data (Hu et al., 2022)
@@
 $tr(I(\theta;D))$ is the trace of $I(\theta;D)$, and $I(\theta;D)$ is the Fisher Information matrix (Becker and Liebig, 2022).
-6.1. The Unlearning Verification and Auditing Methods
-Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effect is also an important problem in machine unlearning.
+5.1. The Unlearning Verification Methods
 In Section 3, we have introduced the L2-norm (Wu et al., 2020), KLD (Nguyen et al., 2020), and privacy leakage as the unlearning verification metrics.
 The common metrics also include accuracy in assessing the performance of the unlearned model and running time in evaluating the unlearning efficiency.
-We have listed recent evaluation and verification metrics in Table 3, where the popular experimental datasets are furthermore presented.
-Since accuracy, running time, L2-norm, and KLD are fixed evaluation metrics, we do not discuss them in detail here.
-In this section, we introduce some new evaluation metrics that recent studies tailored to unlearning.
-Most of them are based on privacy leakage.
-In (Liu and Tsaftaris, 2020), Liu and Tsaftaris indicated that the data auditing problem in machine unlearning is becoming popular.
-They used a calibration dataset and the Kolmogorov-Smirno (KS) (Demidenko, 2004) distance to confirm if an erased dataset is removed.
-Their challenge was that the method would fail when the calibration dataset was out of high quality, or the query samples were similar to the training dataset.
-The EMA (Huang et al., 2021a) method was proposed to overcome these limitations when auditing data erasure from a trained model.
-EMA is a 2-step data removal auditing method that ensembles several metrics and statistical tools.
-In the first step of EMA, the auditor uses various metrics to infer if the model remembers the samples of the query dataset.
-EMA aggregates a final auditing score in the second step by ensembling many membership metrics and statistical tools.
-They verified whether a trained model forgets query data by the final auditing score.
-They conducted the experiments using standard datasets, including MNIST, SVHN, and the Chest X-ray dataset.
+We have listed these evaluation metrics in Table 4.
+Here, we will introduce some new evaluation metrics that recent studies tailored to unlearning.
+Attack-based verification.
 Inspired by backdoor attacks in ML, Hu et al.
@@
 MIB evaluates the membership inference for the triggered data by calculating the results of a certain number of black-box queries to the targetted model.
-Sequentially, they observed the results between the targeted model and clean models to infer whether the model was trained based on the backdoored samples.
+Although MIB is effective to verify the unlearning of backdoored samples, it is hard to directly verify the unlearning of benign samples.
 A similar membership-inference-based method was proposed in (Ma et al., 2022).
 Ma et al.
-(Ma et al., 2022) verified the effect of unlearning methods by the proposed forgetting rate (FR) metric.
+(Ma et al., 2022) verified the effectiveness of unlearning methods by the proposed forgetting rate (FR) metric.
 The evaluation metrics were defined using the observation of membership inference.
@@
 By contrast, the unlearning will be meaningless.
-They also proposed an unlearning method called Forsaken and evaluated Forsaken’s performance on eight standard benchmark datasets.
-Forsaken’s main idea is to generate masked gradients continuously, apply them to the model’s neurons, and model them to forget the contribution of the specific samples.
-To implement this process, they introduced a mask gradient generator.
-The proposed mask gradient generator could adjust the volume to prevent unlearning catastrophes by observing the updated model.
+The membership inference is an important verification tool, but failing an MIA is a necessary but not sufficient condition for successful unlearning.
 Sommer et al.
@@
 In their backdoor-based verification scheme, they first backdoor users’ data and then test the backdoor success probability to infer if the data is unlearned.
+Like the MIB method, Athena is also hard to directly verify the unlearning of benign data, and the backdooring technique should be processed before the original model training.
+Model-centric (influence or model-difference) audits.
+Another line of research views the difference between the model before and after unlearning as a meaningful source of evidence.
+EMU trains reconstruction models over simulated model differences to infer attributes of the deleted data, thereby measuring residual privacy leakage without depending on explicit backdoor design (Wang et al., 2025b).
+Building on this perspective, TAPE focuses on tailored posterior differences and enhances auditing with data perturbation and influence-based partitioning, enabling a more fine-grained assessment of how much private information about forgotten samples can still be recovered across different unlearning settings (Wang et al., 2025c).
+Similarly, TruVRF adopts a non-invasive, model-centric perspective by leveraging model sensitivity to verify unlearning at class-, volume-, and sample-level granularities, extending model-change-based auditing toward finer-grained white-box verification without relying on explicit backdoor design (Zhou et al., 2025).
+These approaches move auditing beyond direct output probing and toward more principled evidence grounded in influence patterns and model-change signatures.
 Table 5. The Employed Datasets in Machine Unlearning
-Data Type Name of Datasets Feature Dimension #. Smaples Task Type Employed by
+Data Type Name of Datasets Feature Dimension #. Samples Task Type Employed by
 Image MNIST (LeCun et al., 1998) $28\times 28\times 1$ $70,000$ Classification (Gupta et al., 2021; Wang et al., 2023a), …
@@
 HIGGS (Baldi et al., 2014) $28$ $11,000,000$ Binary classification (Mahadevan and Mathioudakis, 2021, 2022), …
-YELP2018 $5$ $1,561,406$ Recommendation (Chen et al., 2022a)
-Movielens-1m $4$ $1,000,209$ Recommendation (Chen et al., 2022a)
-Movielens-10m $4$ $10,000,054$ Recommendation (Chen et al., 2022a)
-Text AG News (Zhang et al., 2015) $3$ $127,600$ Text classification (Wu et al., 2020)
-RCV1 (Glewis et al., 2004) $3$ $804,414$ Text classification (Wu et al., 2020)
+YELP2018 55 $1,561,406$ Recommendation (Chen et al., 2022a)
+Movielens-1m 44 $1,000,209$ Recommendation (Chen et al., 2022a)
+Movielens-10m 44 $10,000,054$ Recommendation (Chen et al., 2022a)
+Text AG News (Zhang et al., 2015) 33 $127,600$ Text classification (Wu et al., 2020)
+RCV1 (Glewis et al., 2004) 33 $804,414$ Text classification (Wu et al., 2020)
 Graph Amazon Photo (Shchur et al., 2018) Features per Node: $745$ Nodes: $7,650$ Edges: $119,081$ Node classification Link prediction (Chien et al., 2022a, b), …
@@
 For convenience to find the related studies, we link the corresponding unlearning studies at the last column in Table 5.
-7. Privacy and Security Threats on Machine Unlearning
+6. Domain-Centric Machine Unlearning
+After introducing the classic centralized machine unlearning techniques and corresponding evaluation methods, in this section, we will introduce the studies of domain-centric unlearning, such as federated unlearning, graph unlearning, diffusion model unlearning, and large language model unlearning.
+6.1. Federated Unlearning
+Figure 7. Comparison between (a) server-side federated unlearning and (b) client-side federated unlearning
+FL was initially introduced to protect the privacy of participating clients during the machine learning training process in distributed settings.
+All participants will only upload their locally trained model parameters instead of their sensitive local data to the FL server during model training processes (McMahan et al., 2016).
+Therefore, in a federated learning scenario, limited access to the dataset will become a unique challenge when implementing unlearning.
+According to the unlearning target of a whole client’s contribution or samples’ contribution, we can roughly divide existing unlearning studies into two categories: client-level and sample-level federated unlearning.
+Since the client-level unlearning is usually operated in the server side and the sample-level unlearning usually needs the clients’ participation, we can also call them server-side and client-side federated unlearning, as shown in Figure 7.
+Client-level (Server-side) Federated Unlearning.
+Since the local data cannot be uploaded to the federated learning (FL) server side, most federated unlearning methods try to erase a certain client’s contribution from the trained model by storing and estimating the contribution of uploaded parameters.
+In this situation, they can implement federated unlearning without interacting with the client, shown as the server-side federated unlearning in Fig.
+7 (a).
+The two representative methods are (Liu et al., 2021; Wu et al., 2022b).
+Liu et al.
+(Liu et al., 2021) proposed “FedEraser” to sanitize the impact of a FL client on the global FL model.
+In particular, during FL training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates.
+Based on these operations, they reconstructed the unlearned FL model instead of retraining a new model from scratch.
+However, FedEraser can only unlearn one client’s data, which means it must unlearn all the contributions of this specific client’s data.
+It is unsuitable for a client who wants to unlearn a small piece of his data.
+Study (Wu et al., 2022b) tried to erase a client’s influence from the FL model by removing the historical updates from the global model.
+They implemented federated unlearning by using knowledge distillation to restore the contribution of clients’ models, which does not need to rely on clients’ participation and any data restriction.
+Smaple-level (Client-side) Federated Unlearning.
+Different from unlearning a whole client’s influence and unlearning a class, Liu et al.
+(Liu et al., 2022a; Wang et al., 2023c) investigated how to unlearn data samples in FL, shown as the client-side federated unlearning in Fig.
+7 (b).
+In (Liu et al., 2022a), they first defined a federated unlearning problem and proposed a fast retraining method to withdraw the influence of data from the FL model.
+Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods and the first-order Taylor approximate method (Liu et al., 2022a).
+They utilized the practical Fisher Information Matrix to model the Hessian matrix at a low cost.
+Another similar work based on influence function to implement federated unlearning was introduced in (Wang et al., 2024b).
+Moreover, in (Wang et al., 2023c), the authors implement federated unlearning based on Bayesian inference, and they propose a parameters self-sharing method to reduce the model utility degradation.
+Although these methods effectively implement federated unlearning, they still need some benign clients to participate in the unlearning training, which limits the realistic deployment.
+Figure 8.
+(a) Regular data unlearning, which only unlearns the data sample.
+(b) Graph unlearning.
+Since the graph data contains both node features and connective edge information, we may need to unlearn edge information or both edges and nodes in graph unlearning.
+6.1.1. Graph Unlearning
+We introduce graph unlearning as a representative kind of irregular data unlearning.
+In (Chien et al., 2022a; Chen et al., 2022b; Cong and Mahdavi, 2022; Zhang et al., 2024a), researchers extend regular data machine unlearning to a graph data scenario.
+Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig.
+8.
+Therefore, Chien et al.
+(Chien et al., 2022a) proposed node unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC).
+Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees.
+They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
+Chen et al.
+(Chen et al., 2022b) found that applying SISA (Bourtoule et al., 2021) unlearning methods to graph data learning will severely harm the graph-structured information, resulting in model utility degradation.
+Therefore, they proposed a method called GraphEraser to implement unlearning tailored to graph data.
+Similar to SISA, they first cut off some connecting edges to split the total graph into some sub-graphs.
+Then, they trained the constituent graph models on these sub-graphs and ensembled them for the final prediction task.
+To realize graph unlearning efficiently, they proposed two graph partition algorithms and corresponding aggregation methods based on them.
+In (Cong and Mahdavi, 2022), Cong and Mahdavi filled in the gap between regularly structured data unlearning and graph data unlearning by studying the unlearning problem on the linear-GNN.
+To remove the knowledge of a specified node, they design a projection-based unlearning approach, PROJECTOR, that projects the weights of the pre-trained model onto a subspace irrelevant to the deleted node features.
+PROJECTOR could overcome the challenges caused by node dependency and is guaranteed to unlearn the deleted node features from the pre-trained model.
+6.2. Unlearning Diffusion Models
+Exploring unlearning solutions for diffusion models and large language models (LLMs) is popular in recent years.
+We will first introduce the literature about unlearning methods for diffusion models and then introduce the LLMs unlearning in the next subsection.
+Unlike the classic classification unlearning setting, unlearning for generative models must consider about open-ended conditional output distributions.
+The unlearned model should cease to reproduce specific training examples and protected artistic styles.
+Commonly, diffusion model unlearning methods can be organized by what is removed.
+Firstly, data-centric deletion targets the influence of designated training points and aims to approximate retraining on a pruned dataset (Li et al., 2024b; Alberti et al., 2025; Chen et al., 2024).
+Secondly, concept-centric deletion removes higher-level semantic notions (e.g., nudity or protected styles) by intervening in the model’s internal representations and denoising dynamics (Wu et al., 2025a; Zhang et al., 2024b; Wu et al., 2025b; Cywiński and Deja, 2025).
+Data-centric deletion.
+A line of the diffusion model unlearning treats unlearning as an approximate data removal problem.
+The goal of these studies is to match the behaviour of a model retrained from scratch based only on the remaining dataset.
+Li et al.
+(Li et al., 2024b) studied machine unlearning for image-to-image generative models.
+They contributed a general unlearning framework for structured translation models such as diffusion models, VQ-GAN, and MAE.
+They showed on ImageNet-1k and Places-365 that their method can effectively forget target samples with only a small degradation on retrained ones.
+Alberti et al.
+(Alberti et al., 2025) focused on data unlearning in diffusion models.
+They defined retraining without the target data as the gold standard, then proposed SISS (Subtracted Importance Sampled Scores), a method that combines quality preservation and forgetting strength through importance sampling.
+Experiments on CelebA-HQ, MNIST with T-shirt, and Stable Diffusion show that SISS achieves a strong trade-off between forgetting and generation quality.
+Chen et al.
+(Chen et al., 2024) proposed Score Forgetting Distillation (SFD), a data-free unlearning method for diffusion models.
+Instead of using original training data, SFD adds an unlearning objective into score distillation, encouraging the model to align the scores of unsafe classes.
+Concept-centric deletion.
+A second line of diffusion unlearning studies targets semantic removal rather than individual data points.
+They usually aim to erase concepts or styles that may be harmful.
+EraseDiff (Wu et al., 2025a) formulates concept removal as a constrained optimization problem: the model should move away from denoising trajectories associated with the forget set, while staying close to behaviors needed for retained content.
+AdvUnlearn (Zhang et al., 2024b) shifts the focus to robustness, integrating adversarial training into the unlearning process so that erased concepts cannot be easily recovered through adversarial prompts.
+Importantly, it shows that editing the text encoder can yield stronger and more transferable robustness than editing the UNet alone.
+DoCo (Wu et al., 2025b) further targets generalization and utility preservation by aligning the output domains of sensitive and anchor concepts and using gradient surgery to reduce interference with related benign concepts.
+In contrast to these parameter-level approaches, SAeUron (Cywiński and Deja, 2025) offers a more interpretable feature-level intervention by training sparse autoencoders on internal activations and blocking the features associated with unwanted concepts.
+Thereby, it enables transparent and multi-concept unlearning with competitive robustness and generation quality.
+Cross-cutting evaluation.
+For diffusion models, unlearning evaluation should go beyond checking whether a target concept disappears in a few post-edit samples.
+A convincing evaluation framework must assess at least four properties: forgetting efficacy, utility preservation, robustness, and durability.
+Zhang et al.
+(Zhang et al., 2024c) make this point explicit by introducing UnlearnCanvas, a high-resolution benchmark for object and style unlearning that provides a standardised and automated evaluation pipeline with seven quantitative metrics.
+They also examine harder settings such as adversarial prompts, finer-grained concept removal, and sequential unlearning, showing that immediate deletion performance alone is insufficient for judging success.
+Ko et al.
+(Ko et al., 2024) further show that diffusion unlearning can substantially impair text–image alignment, and therefore argue that evaluation must jointly measure forgetting and alignment rather than treating utility as generic sample quality alone.
+Their post-unlearning alignment framework is motivated precisely by this tension.
+George et al.
+(George et al., 2025) extend evaluation to the temporal dimension by showing that supposedly erased concepts can re-emerge after benign fine-tuning.
+These studies suggest that diffusion unlearning should be evaluated not only by how much is forgotten immediately after editing, but also by whether the model retains non-target capabilities, resists adversarial recovery, and remains stably unlearned after subsequent updates.
+6.3. Unlearning Large Language Models
+Unlearning in large language models (LLMs) seeks to remove the concepts and behaviours, such as copyrighted text, personal information, and unsafe responses.
+Recent studies conceptualize this problem as eliminating undesirable training influence and the capabilities associated with it under two key requirements: locality and utility, meaning that changes should be confined to the targeted knowledge and general performance on unrelated tasks should remain largely intact (Yao et al., 2024a, b; Liu et al., 2025a).
+From this perspective, existing approaches can be broadly divided into two categories according to where forgetting is implemented.
+The first is parametric post-hoc unlearning, which directly updates model parameters so that forgetting becomes embedded within the model itself (Yao et al., 2024b, a; Liu et al., 2024a; Huang et al., 2024a; Wang et al., 2025a).
+The second is non-parametric or suppression-based unlearning, which leaves the underlying model unchanged and instead prevents undesirable outputs through inference-time intervention or control (Liu et al., 2024b).
+Parametric post-hoc unlearning.
+Most LLMs unlearning studies are about post-hoc unlearning directly on pretrained LLMs by updating weights to suppress targeted information.
+Large Language Model Unlearning (Yao et al., 2024a) presents unlearning primarily as an alignment mechanism, arguing that targeted forgetting can suppress harmful outputs, remove copyrighted content, and reduce hallucinations using only negative examples.
+Machine Unlearning of Pre-trained Large Language Models (Yao et al., 2024b), in contrast, takes a more systematic and benchmark-driven view, evaluating multiple unlearning strategies on pre-trained models and showing that gradient-based post-hoc unlearning can be highly efficient relative to retraining, especially when forgetting updates are combined with utility-preserving regularisation on retained data.
+Building on early gradient-based approaches, subsequent work develops more targeted unlearning mechanisms to better satisfy the locality requirement.
+Liu et al.
+(Liu et al., 2024a) propose Selective Knowledge-negation Unlearning (SKU), a two-stage framework that first identifies harmful knowledge and then explicitly negates it, aiming to suppress harmful responses while preserving normal utility.
+Wang et al.
+(Wang et al., 2025a) further refine the granularity of forgetting with SeUL, which performs unlearning at the level of sensitive spans instead of whole sequences and introduces dedicated metrics to quantify sensitive-information deletion.
+These methods reflect a broader shift from coarse unlearning toward more selective and locality-aware interventions, although they differ in whether forgetting is enforced through parameter negation, output-space offsets, or span-level objective design.
+Non-parametric unlearning (inference-time control).
+Unlike parametric unlearning methods that edit model weights, Liu et al.
+(Liu et al., 2024b) shift forgetting to the inference stage.
+Their Embedding-Corrupted (ECO) framework treats unlearning as a conditional control problem: the base model remains intact, while a detector identifies prompts associated with forgotten content and an embedding-level intervention steers generation away from that content.
+In this sense, ECO does not make forgetting intrinsic to the model; rather, it imposes an “unlearned state” on demand for selected inputs.
+This design is attractive operationally because it is modular, scalable, and easy to deploy across different LLM backbones, especially in settings where direct access to model parameters is unavailable or repeated fine-tuning would be too costly.
+Evaluation infrastructure for LLMs unlearning.
+Across recent work, evaluation in LLM unlearning is becoming both broader and more demanding.
+Jin et al.
+(Jin et al., 2024) contribute a dedicated text-only benchmark, RWKU, which evaluates forgetting of real-world entity knowledge on LLaMA3-Instruct (8B) and Phi-3 Mini-4K-Instruct (3.8B) using membership-inference attacks, adversarial probes, and retain tasks drawn from MMLU, BBH, TruthfulQA, TriviaQA, and AlpacaEval.
+Liu et al.
+(Liu et al., 2025a) provide a more general evaluation framework, arguing that unlearning should be assessed not only by immediate deletion success but also by scope precision, privacy leakage, robustness, consistency, and effects on causally unrelated capabilities.
+Extending this agenda beyond text-only models, Li et al.
+(Li et al., 2024c) propose SIU for multimodal LLMs and introduce MMUBench, built from concepts sampled from MIKE, image collections from web search, and evaluation on LLaVA 7B/13B together with standard multimodal benchmarks such as GQA, VQA-v2, VizWiz, ScienceQA-IMG, TextVQA, POPE, MMBench, and MM-Vet.
+Taken together, these studies show that unlearning evaluation is moving from coarse deletion scores toward more rigorous, attack-aware, locality-sensitive, and increasingly multimodal assessment.
+7. Privacy and Security Issues on Machine Unlearning
 Although unlearning methods were initially proposed to safeguard users’ privacy, many researchers have noticed that unlearning brings new privacy threats simultaneously.
-Similar to machine learning scenarios, there are two main kinds of privacy threats in unlearning.
-They are membership inference attacks and private information reconstruction attacks.
-In this section, we detail the two prevalent unlearning privacy threats.
+In this section, we will discuss these privacy and security studies about machine unlearning.
 7.1. Privacy Threats on Machine Unlearning
@@
 Zanella-Béguelin et al.
-found five phenomenons.
+found five phenomena.
 First, an attacker can extract particular sentences used or not in the training dataset by comparing two models.
@@
 Fifth, to mitigate the privacy leakage risks while keeping the model utility, the server can limit the model parameters access or only output a subset of the results.
-Many studies further utilized these privacy threats to evaluate the unlearning effect.
+Many studies further utilized these privacy threats to evaluate the unlearning effectiveness.
 Huang et al.
@@
 They verified their methods by testing how well the method defends against privacy attacks.
-Both (Chen et al., 2021) and (Zanella-Béguelin et al., 2020) pointed out that differential privacy guarantees that a model does not reveal too much knowledge about any training sample.
-A DP-protected model can further guarantee the group’s privacy by binding the impacts of a bunch of training samples.
-If using DP to protect the privacy of a bunch of $|D\backslash D_{e}|$training examples against snapshot attacks on $\theta_{D},\theta_{D\backslash D_{e}}$, it means that $\theta_{D\backslash D_{e}}$ cannot be more useful than $\theta_{D}$.
+[…]
 7.2. Security Threats on Machine Unlearning
-There are several papers that address security threats targeting machine unlearning.
+There are several papers that introduced security threats targeting machine unlearning.
 For instance, Marchant et al.
@@
 (Hu et al., ) investigate security vulnerabilities within unlearning systems, specifically by uploading customized malicious data update requests to negatively influence the model utility.
-Additionally, some studies propose backdooring or poisoning methods that exploit unlearning requests to achieve backdoor insertion, as discussed by (Di et al., 2022) and (Liu et al., 2024).
+Additionally, some studies propose backdooring or poisoning methods that exploit unlearning requests to achieve backdoor insertion, as discussed by (Di et al., 2022) and (Liu et al., 2024c).
+Huang et al.
+(Huang et al., 2024b) introduced an unlearning-activated backdoor attack that uses influence-based analysis to select camouflage to trigger samples, so that the backdoor remains stealthy before unlearning but becomes effective after specific deletions have taken place.
 These works collectively highlight the need for robust security measures in machine unlearning to mitigate these emerging threats.
+7.3. Defending Methods
+From a defensive standpoint, only a limited number of studies have begun to design privacy-preserving and robust unlearning methods to defend against these emerging threats.
+Representative examples are BlindU (Wang et al., 2026) and Compressive Representation Forgetting (CRFU) (Wang et al., 2025d).
+CRFU mitigates reconstruction attacks and related forms of information leakage attacks by imposing an information-bottleneck-inspired compression objective on the representations of data designated for removal (Wang et al., 2025d).
+In doing so, CRFU aims to suppress the mutual information between latent representations and the erased inputs, while preserving model utility through remembering objectives and a controllable unlearning rate (Wang et al., 2025d).
+BlindU provides a stricter privacy-preserving unlearning framework compared with CRFU.
+Privacy preservation of BlindU is achieved by uploading compressed and differential privacy-enabled data as the unlearning requests, and an unlearning method in BlindU for compressed data is designed correspondingly (Wang et al., 2026).
+More broadly, this line of work signals a move away from viewing unlearning as a simple post hoc algorithmic adjustment, and toward framing it as a joint optimisation problem that simultaneously addresses forgetting, privacy protection, and robustness.
 8. Machine Unlearning Applications
@@
 Literature Application Scenarios Methods Realization Methods Evaluation Metric Year
+(Liu et al., 2022b) Mitigate backdoor BAERASER Gradient ascent method ASR and ACC 2022
+(Wang et al., 2019) Mitigate backdoor - Median Absolute Deviation L1-Norm, #\#FP, ACC, ASR 2019
+(Du et al., 2019) Anomaly detection - Maintain memory set #\#FP and #\#FN 2019
+(Cao et al., 2018) Repair pollution KARMA Cluster and unlearn DA 2018
 (Huang et al., 2021b) Data “unlearnable” EMP Add noise ACC 2021
-(Liu et al., 2022b) Mitigate backdoor BAERASER Gradient ascent method ASR and ACC 2022
-(Wang et al., 2019) Mitigate backdoor - Median Absolute Deviation L1-Norm, ##\##FP, ACC, ASR 2019
-(Du et al., 2019) Anomaly detection - Maintain memory set ##\##FP and ##\##FN 2019
-(Cao et al., 2018) Repair pollution KARMA Cluster and unlearn DA 2018
   • ASR:Attack Success Rate, ACC: Accuracy
   • DA:Detection Accuracy, EMP: Error-minimizing Perturbations
-  • ##\##FN: number of false negative, ##\##FP: number of false positive
+  • #\#FN: number of false negative, #\#FP: number of false positive
 Besides the inherent demands of machine unlearning that draw much research attention, many researchers also find that machine unlearning can be applied in many other scenarios to solve related problems.
@@
 The most popular application of machine unlearning is to mitigate the anomalies, including backdoor triggers and pollution from an already-trained model.
+Mitigating Anomaly.
 Study (Wang et al., 2019) applied machine unlearning in detecting and mitigating backdoor attacks in DNNs.
@@
 Third, KARMA determines if a causality cluster is polluted and calculates how many samples the cluster contains.
-Machine unlearning was also applied to lifelong anomaly detection problems.
 Du et al.
-(Du et al., 2019) proposed new approaches to solve the challenges of exploding loss and sharp accuracy degradation caused by unlearning.
-In their presentation, they explain their techniques, emphasizing handling false negatives.
-When meeting such case $x_{t}$, their purpose is, therefore, to modify the model to reduce the possibility of $p(x_{t})$, or to increase the loss $\mathcal{L}(x_{t})$.
+(Du et al., 2019) also explored unlearning in lifelong anomaly detection and tried to mitigating exploding loss and sharp accuracy degradation caused by unlearning.
+Data Unlearnable.
 Huang et al.
@@
 9. Lessons Learnt and Discussions
-In the research domain of machine unlearning, researchers mainly face three difficult challenges, the stochasticity of training, incrementality of training, and catastrophe of unlearning, when they try to design efficient and effective unlearning mechanisms replacing naive retraining.
-They tried many mechanisms to mitigate the influence of these challenges.
-For example, in exact unlearning, they designed split algorithms that divide the final model into several sub-models and avoid the stochasticity and incrementality of one sub-model to influence other sub-models.
-In approximate unlearning, they bounded the removed estimation to avoid accuracy degradation.
+In the research domain of machine unlearning, researchers mainly face three difficult challenges, the stochasticity of training, incrementality of training, and catastrophe of unlearning.
+Researchers tried many mechanisms to mitigate the influence of these challenges.
+For example, in exact unlearning, they designed split algorithms that divide the final model into several sub-models and avoid the stochasticity and incrementality of one sub-model to influence other sub-models (Bourtoule et al., 2021).
+In approximate unlearning, they bounded the removed estimation to avoid accuracy degradation (Guo et al., 2020).
 However, it is still not easy to analyze them clearly and solve them totally.
-Moreover, as the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met corresponding new challenges.
+As the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met corresponding new challenges.
 Here, we discuss the differences between various unlearning scenarios and their corresponding challenges.
@@
 Existing methods try to design new unlearning mechanisms to reduce the cost from the two aspects.
-Although they proposed many methods, they just mitigated the influence of the main challenges and still have not solved them.
-Existing unlearning methods were at the beginning phase, which is trying to implement effective unlearning.
-They proposed mechanisms, but most of them cannot guarantee the final results, so they also bound the unlearned item (Guo et al., 2020) or set an unlearning threshold (Nguyen et al., 2020).
-Although they also introduced many metrics to verify the unlearning effect, the unlearning and verification processes are split.
-This means that the unlearning result is uncertain during the unlearning period before the verification is finished.
-Therefore, we find and list some open questions in centralized unlearning.
+Although researchers proposed many methods, they just solved the challenges partially and cannot guarantee the final unlearning effect.
+Here, we find and list some open questions in traditional centralized unlearning.
   1. (1)
-    Optimizing the existing challenges.
-The first two challenges are hard to solve exactly, so recent research focuses on reducing utility degradation after unlearning.
-Even though many studies exist, there is still a long way to go to eliminate the degradation of unlearning.
+    Challenge: The stochasticity of training and incrementality of training challenges are hard to solve exactly, so recent research mainly focuses on reducing utility degradation after unlearning.
+Potential Direction: Exploring the unlearning problem with a theoretical guarantee, such as from a differential geometry perspective, will provide the theoretical contribution and explanation to unlearning.
   2. (2)
-    How do we unlearn with a certain or exact goal?
-In other words, can we unlearn as learning?
-If we know the purpose, to what extent we have unlearned, and when we can finish or stop the unlearning process.
+    Challenge: What unlearning mechanisms should be to forget a certain or exact goal in different models, such as in diffusion model or LLMs?
+Potential Direction: Exploring to what extent we have to unlearn for different unlearning purposes remains under-explored, and how to utilize the learning mechanism to implement unlearning purposes in also an interesting direction.
 Machine unlearning in distributed scenarios has many differences from centralized scenarios.
@@
   1. (1)
-    Federated unlearning cannot use the fast retraining methods because data is out of reach for the server.
-Therefore, federated unlearning can only be implemented using approximate unlearning methods.
-However, as we know, approximate unlearning easily causes catastrophic unlearning, and federated learning is more vulnerable to degradation, so controlling the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
+    Challenge: Federated unlearning is hard to be implemented by using the fast retraining methods because data is out of reach for the server.
+Most federated unlearning methods are implemented by using approximate unlearning methods.
+However, as we know, approximate unlearning easily causes catastrophic unlearning, and federated learning is more vulnerable to degradation.
+Potential Direction: Controlling the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
   2. (2)
-    Existing federated unlearning methods require the activation of all users, including those without unlearning requests, to assist in the unlearning process.
+    Challenge: Existing federated unlearning methods require the activation of all users, including those without unlearning requests, to assist in the unlearning process.
 These approaches are impractical and inefficient, particularly when unlearning requests are frequent.
-Therefore, it is crucial to study how to balance the unlearning effect and efficiency in federated unlearning.
+Potential Direction: It is crucial to study how to balance the unlearning effectiveness and efficiency in federated unlearning.
 Besides exploring machine unlearning based on regularly structured data, researchers tried to implement unlearning in graph data.
@@
 In graph unlearning, unlearning some edges or sub-graphs is a big question.
+Unlearning in LLMs and diffusion models introduces additional challenges beyond those in conventional discriminative models.
+In these generative models, the target to be removed is often distributed across parameters and entangled with related knowledge, so forgetting a specific sample, entity, or concept does not necessarily prevent the model from reproducing semantically related variants (Liu et al., 2025a; Chen et al., 2024).
+Existing evaluation is also still immature: LLM unlearning is heavily concentrated on a small number of fixed forget-retain benchmarks (Jin et al., 2024), while diffusion unlearning studies have shown that erased concepts may reappear under nearby or adversarial prompts (George et al., 2025).
+Therefore, future research should focus on defining unlearning targets and scopes more precisely, such as at the entity, concept, and hierarchical knowledge levels, developing more realistic and robust benchmarks, improving continual and black-box unlearning for real deployment settings, and enhancing robustness and interpretability.
 After designing machine unlearning algorithms, effective verification and auditing methods are necessary (Thudi et al., 2022b).
@@
 We first illustrate the complete unlearning framework, including the learning, request, unlearning, and verification.
-Then, we briefly categorize recent studies into exact and approximate unlearning and introduce the technical details of corresponding unlearning methods.
-Moreover, we noticed some new unlearning scenarios, such as federated unlearning and graph unlearning, which are also introduced in the survey.
+Then, we briefly categorize recent studies into exact and approximate unlearning and introduce the technical details of the corresponding unlearning methods.
+Moreover, we noticed some new unlearning scenarios, such as graph unlearning, federated unlearning, diffusion model unlearning, and LLMs unlearning, which are also introduced in the survey.
 Besides, we consider privacy and security issues in machine unlearning to be an important part of the studies.